A Complete Bibliography of Publications in R News and the R Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

21 May 2021 Version 1.02

Title word cross-reference

3 [MSC14]. 4 [MSC14]. g [Pra20]. H [HNL12, Pra20]. K [DDC07, Pra20, WS11]. O(1) [O’N15]. R × C [LMK07]. t [AHvD09, AH10, Hof13, HBG01, Liu14, MUX20, Suc13].

-and- [Pra20]. -EGARCH [Suc13]. -likelihood [HNL12]. -means [WS11]. -Statistics [Liu14]. -table [DDC07].

2.0.0 [Rip04]. 2.1.0 [Rip05a, Rip05b]. 2013 [Use13, Ulr13]. 2015 [Tve15]. 2016 [BAB^{+}17, Ric16]. 2017 [BDH17, Ver17]. 2018 [AdSQ18, BKJ^{+}18, Dar18, Ric18]. 2019 [BPC^{+}20, CvDMvE19]. 22K [Dem16].

4 [Ray01]. 4.0 [KMH20].

5 [SFMR16].

6th [LG13].

Aalen [ABS08, WBS06]. abctools [NP15]. Academic [BSZWW14]. Accelerated [KF16, SP15]. Acceleration [BK19]. Accelerometer [VP14]. Access [Car06, KRJ16, Wal16]. Accessing [Lig06]. Accountable [GGHB18]. Achievement [LDM03]. acids [WV06]. across [COGRB16]. ACSNMineR [DBB^{+}16]. Actual [OOFdlFS^{+}16]. actuar [GP08]. Adaptive [AHvD09, Yoo18]. Add [OS18]. Add-on [OS18]. added [BGK18]. Addendum [GE14]. addhaz [dLN^{+}18].

Additive [BTH^{+}17, SS18]. ade4 [CT06, CDT04, DDC07]. adegraphics [SJLD^{+}17]. Adjusted [BFK^{+}20]. Adjusting [GK17]. AdMit [AHvD09]. Advanced [Bür18, OS18]. AFM [BBTH17]. afmToolkit [BBTH17]. Again [LDB13]. Against [AG20]. Age [Nie15]. Aggregated [KBBH15]. Agnostic [GB19]. Agriculture [OC07]. Aid [KRJ16]. Air [RC12]. Ake [WSK16]. algebras [Han06b]. Algorithm [ALB^{+}19, BACMÁK17, KLK20, MMH09, Yoo18]. Algorithms [CS16, DY16, Kap14, MNT18, Pal18, BC18]. Alignments [DSN17]. alineR [DSN17]. Allele [MRS17]. Allow [HvWDM17]. Alternatives [AG20]. ALTopt [SP15]. amap [LJ06]. Amazon [Lee16]. AMMI [OC07]. Among [BKvB^{+}17, IAA16]. Amplification [RBB^{+}15]. Analyse [DSTT06]. Analyses [NP15, WFRM15]. Analysing [BBD12, DGF08, HHB18, KE06, Lum03]. Analysis [AHLL19, BT15, Bea16, BBTH17, BRS18, BFHM11, CCFVM17, CKL17, Chr18, Dal07, Dam09, Dja19, DYB02, ERK16, EZ20, FKW^{+}18, FR06a, FKP19, FLO20, GGHB18, God13, GKZK16, Gra14, Hol06, IPRT19, IW17, IMA16, JA18, JM17, JM16, KM08, KG06, KZDZ18, KBBH13, KBBH15, LHKB17, LCT20, LB16, Luq07, MD18, MSN^{+}16, NDWL20, Nie15, OC07, OSS^{+}17a, PMZ20, PFMCFC18, PBCV06, PS16, RD01, RBB^{+}15, RBS13, RC12, SHW18, Sch01, Sch07, SH06, Smi03, SFdC18, SB19, Sto14, TDAP18, TD13, TVMZ10, Wie17, Mar02, Mur05b]. Analyze [GHM16, SJB20, Wri16]. AnalyzeFMRI [Mar02]. Analyzing [Bec16, CGG13, HWW12, KLN11, KCK^{+}20, KB14, LH06, LZN^{+}20, MKL20, OK07, PT07, PLH^{+}10, Sha13, XY19]. anchoredDistr [SHR17]. animation [XC08]. Animations [XC08]. Annealing [XGSH13]. Annotation [Lig02]. anomalies [WAD^{+}19]. Anomaly [GBB^{+}17]. anomalyDetection [GBB^{+}17]. Anonymization [LOM20]. Antimicrobial [ORVT15]. apc [Nie15]. apComplex [Sch06]. API [GdC11, Win17]. Apple [AZJH12]. Application [BND11, BS12, CKL17, JM16, MHR13, NW07, TVMZ10]. Applications [FR06b, NEB^{+}20]. Applied [Jar07, MQ06]. Applying [ARA^{+}14]. Approach [BRS18, CT06, PGCE17, VB15]. Approaches [OSS^{+}17a]. Approximate [NP15, van14]. Approximated [NS17]. Approximating [LQ18]. AQuadtree [LOM20]. Aquatic [CGG13]. arc [HJKK19]. ARCH [Ott19]. Archive [VFB16]. Archiving [Lee14]. ArCo [FMMV18]. Area [BTH^{+}17, MM15, Mor18, Wie17]. Areal [Hug14]. Argument [Lan17]. ari [KLM20]. Arrays [RH06]. Artificial [FMMV18]. arulesCBA [HJKK19]. arulesViz [Hah17]. Aspects [Fox09]. Assembling [GSBT^{+}20]. Assess [DD18]. Assessing [BWM14, FLSNG19, KGZ14]. Assessment [ZC16]. Associated [WSK16]. Association [Hah17]. Associative [HJKK19]. Assumptions [KCK^{+}20]. Asymmetric [Zag13]. asympTest [CDLR09]. atable [Str19]. Attribution [LK13]. auditor [GB19]. Augmented [GBB^{+}17]. Augmenting [Rob07]. Authenticating [BS12]. Authors [HMZ12]. autoimage [Fre17]. Autologistic [Hug14]. Automated [BBTH17, SFdC18, SB19, Jur12, KLM20]. Automatic [Oh14]. Automation [Lig02]. Autonomous [Gra14]. Autoregressive [ARS15, CLLG08, HWW12, JP18b, WC17]. Average [SY19]. Averaging [GDF18, RPV05]. Avoid [LF08].

B [GF19]. Back [LDB13]. Backtests

[CEGK07]. Bacterial [TVMZ10]. Balance [OOFdlFS^{+}16]. Balances [BKvB^{+}17]. Balloon [JW06]. Base [Mur03, Sch03]. Based [BRS18, BS13, BFHM11, BACMÁK17, BW15, CFF20, FR06b, FKP19, IMA16, KH19, Liu14, PR17, SJLD^{+}17, Sto14, Wai16, FLSNG19, FR06a, KM12, Smi03, SD20, WBS06, WCAR16]. Basis [PBMSN20, Yoo18]. Basis-Adaptive [Yoo18]. Basket [KCK^{+}20]. Bass [Dug17]. BayesBD [SL17]. BayesBinMix [PR17]. bayesdfa [WAD^{+}19]. Bayesian [AB17, AH10, Bür18, CLLG08, CG06, Fer10, GDF18, Jar07, KG06, LC15, LK15, MQ06, MBL18, NP15, Pap18, PBMSN20, RPV05, SHR17, SCLV20, SL17]. BayesMallows [SCLV20]. BayesVarSel [GDF18]. beadarray [DSTT06]. beadarrayFilter [FVK^{+}13]. BeadArrays [DSTT06]. Beads [FVK^{+}13]. Behaviour [Luq07]. Beowulf [CMM03]. Best [LGE^{+}18]. Beta [LK15, Suc13]. Beta-Skew- [Suc13]. betategarch [Suc13]. BETS [SFdC18]. between [PMMEGM19]. BFGS [GF19]. Bias [MMN17]. Biased [LZN^{+}20]. Biclustering [RPO^{+}19]. biclustermd [RPO^{+}19]. Big [Qia14, Wri16, ZB20]. biglasso [ZB20]. Binary [GMBL18, Koe06, LRMLLCS17, Mur07, PR17, SY19, Tou16]. BINCOR [PMMEGM19]. binGroup [BZST10]. Binomial [LQ18]. bioassays [Ran06]. Bioconductor [GC02, ACD^{+}16, Car06, LH06, LC06, Smi03, The14a, The15a, The15b, Tea16, The16a, Tea17a, Tea17b, Tea18a, Tea18b, Tea19, Tea20]. Biological [DBB^{+}16, Dur06, Wri16]. biomaRt [Dur06]. bipartite [DGF08]. Birnbaum [LHR06]. Bivariate [Liu14, SD20]. Blind [God13, GE14]. BMA [RPV05]. bnclassify [MBL18]. BNSP [Pap18]. Bond [Dja19]. BondValuation [Dja19]. Book [KLN11, Lia07, Hol06, Lei07]. Boolean [Dus18, LVV11]. Boosting [MHL^{+}20]. boot [Can02]. Bootstrap [CGLN18]. Boundaries [SL17]. Boxland [SWCH16]. Bradley [Han17]. Brain [Chr18]. brainR [MSC14]. breakDown [SB18]. brinton [MMO20]. brms [Bür18]. Brobdingnag [Han07a]. Broken [Mug08]. Broken-Line [Mug08]. BSGS [LC15]. bshazard [RSR14]. Budget [JA18]. BUGS [LDB13, Tho06, TOLS06, Yan04]. Builder [PS16]. Building [YR03]. Bundle [Bra05]. Burden [dLN^{+}18]. Bureau [Wal16].

C [Lan01a, Plu11]. C/C [Lan01a]. CA [KBBH15]. CA-GALT [KBBH15]. Calculating [DBB^{+}16, HTM15]. Calculations [Bat04, DEVK19, GS16, MUX20]. Can [LF08]. Capable [KF16]. car [FFW13]. carboxylic [WV06]. Carlo [Jon18b]. carx [WC17]. Case [AE17, Jon18a, Mor20, SS16]. Case-Cohort [Jon18a]. Case-Control [SS16]. Categorical [HR07, KZDZ18, KB14, NR12]. Causal [BT15, CCFVL20]. Causality [Hma20]. cchs [Jon18a]. cDNA [Saw02]. Censored [GSHB14, MMZ16, SD20, WC17, AB17]. Censoring [LGSG18]. Census [Wal16]. Centered [Hug14]. Central [Liu14, Dal01a]. Centralized [SN18]. Chain [Pfe17]. Chained [SS18]. Chains [Sar13, Spe17]. Change [EZ20, Zei01, Poc06, WBS06]. changeLOS [WBS06]. Changes [Dal02, HZ13a, HZ13b, HZ14a, HZ14b, HZ15a, HZ15b, HZ16a, HZ16b, HLZ17a, HLZ17b, HLZ18a, HLZ18b, HLZ19a, HLZ19b, HLZ20a, HLZ20b, KMH20, Mur05a, Mur13, The13a, The13b, The14b, The14c, The15d, The15c, The16b, The16c, The17a, The17b, The18a, The18b]. Channel [KY15]. charting [Scr04]. Charts [GK17]. Checking [ZH02]. checkmate [Lan17]. Checks [Lan17]. Chemistry [FR06b]. chi [Dal01a]. chi-squared [Dal01a]. Chinese [LG13]. Chronic [dLN^{+}18]. ciuupi [MK19]. Ckmeans.1d.dp [WS11]. Class [BB16, Lum04b, LMT14]. Classes [GP04, PB05, RH01, RKSC06]. Classical [CDLR09]. Classification [CJB^{+}13, HJKK19, Jur12, LW02, LRMLLCS17, PAC^{+}17, RdC18, SFMR16].

Classifiers [MBL18]. cleanNLP [Arn17]. clere [YJBC16]. Climate [AEB17, Poc06]. Clinical [Str19, GB06, PZK06]. Clone [BG20]. Cloning [Sól10]. Cloud [WCAR16]. Cloud-based [WCAR16]. Club [SP19]. Cluster [AHLL19, CMM03, HP19, LR01, PWD17, YLGT19]. Cluster-correlated [AHLL19]. Clustered [CA19, YLGT19]. Clustering [BW15, FGS19, FR06b, HQ16, JGB14, LJ06, PR17, SE19, SFMR16, SP19, Sze18, Szk16, VCQ15, WS11, YJBC16]. clustering.sc.dp [Szk16]. clustMixType [Sze18]. ClustVarLV [VCQ15]. CMatching [CA19]. CMD [LM05]. CODA [PBCV06]. Code [BG20, MKL20, RdV15]. Coding [Hol06, NtGP17, Mur05b]. Coefficient [LGSG18, dCSPM19]. Cognitive [Chr18]. Cohort [Jon18a, Nie15]. Cohorts [KM08]. Coincidence [BT15]. Cointegrating [JP18b]. Collaborative [KMKT20, TZ09, WCAR16]. Collection [Ger16]. Collections [Bar18]. Collinear [IAA17]. Collinearity [IAA16]. Color [VS15]. Column [BCD^{+}17, GDM^{+}17]. Combinations [WRR18]. comf [Sch16]. Comfort [Sch16]. Commander [Fox07]. common [WAD^{+}19]. Community [RC12]. Comparative [TD13]. Compare [BBACK18]. Comparing [wCTKL20, COGRB16, Hoc19, Mur08]. Comparison [CS16, FRL^{+}16]. Comparisons [BHW02, VS15]. Compatible [ALH18]. Compendium [Lun11]. Competing [KSLS20, AHLL19]. Complex [BT15, GHM16, Han15a, Sch06]. CompLognormal [NB13]. Composed [MRS17]. Composite [NB13, SH20]. Comprehensive [IAA17, UAA18]. Computation [Fer10, KF16, KY15, NP15, ZB20]. Computation-Efficient [ZB20]. Computational [CMM03, ERK16, SH20]. Computations [VDK15]. compute [BECDN17]. Computer [ORC^{+}18].

Computers [AZJH12]. Computing

[BF14, KPBS09, KH19, LR01, MK19, Roo16, WS13, Yu02]. Concept [Lun11]. Conditional [ARS15, BILL16, LP14, MMS16]. Condor [Xie05]. condSURV [MMS16]. ConectaR [CvDMvE19]. Conference [AdSQ18, BDH17, BSZWW14, BAB^{+}17, BKJ^{+}18, BPC^{+}20, Use13, CvDMvE19, Dar18, GT14, LG13, LAL18, Ric16, Ric18, RBK18, SD13, Tve15, Ulr13, Ver17, FVS^{+}20, LG13]. Confidence [CDLR09, Fay10, MK19, WS13]. Configurational [BT15]. Connecting [KLC19, MSLC15, WCAR16]. Connections [Rip01a]. Consistency [Dus18]. Consistent [Wic10]. Consortium [Plu15]. Constrained [ZZZ^{+}19, YLGT19]. Constraint [Szk16]. Constructing [Gre17]. Containerisation [NEB^{+}20]. Containers [BE17]. Contamination [ANRG18]. Content [BFHM11]. Content-Based [BFHM11]. Contingency [Dem16, KBBH13]. Continuous [FT19, HR07, LS20, LRMLLCS17, NR12, Pfe17, Sar13, WSK16]. Continuous-Time [Pfe17]. Contrasts [Arm13]. Contribution [dLN^{+}18]. Control [Bet05, GK17, Saw02, SS16, Scr04]. Controlling [OL09]. Conventions [Båå12]. Convergence [ALB^{+}19, LL09, Mar11, PBCV06, SP19]. ConvergenceClubs [SP19]. ConvergenceConcepts [LL09]. Conversion [Oh14]. Converting [Bat03]. Convex [KLK20, HY18]. Coordinate [Bro16, Sto14]. Coordinate-Based [Sto14]. Coordinates [Fre17]. copas [CRS09]. Copula [SD20]. CopulaCenR [SD20]. coRanking [KRM18]. Correction [MMN17]. Correlated [FT19, Tou16, AHLL19]. Correlation [LGSG18, PF16, PMMEGM19]. Correspondence [KBBH15, LB16, Mur05b, Hol06]. Cosmological [FLO20]. Costa [Mor20]. Costs [EKN^{+}08]. Count [FT19]. Counterfactual [CCFVM17, CCFVM17]. Counterfactuals [FMMV18]. Counting [NTE14]. Country [WG10]. Coupon [Dja19]. Covariance [PF16]. Covariance/Correlation [PF16]. Covariate [KM08, XY19, YLGT19]. Covariate-constrained [YLGT19]. Covariates [BHM^{+}17, IW17, WC17]. Cover [EZ20]. Cox [KH19, LZN^{+}20, NS17, OSS^{+}17b]. coxed [KH19]. CoxPhLb [LZN^{+}20]. coxphMIC [NS17]. CRAN [HZ13a, HZ13b, HZ14a, HZ14b, HZ15a, HZ15b, HZ16a, HZ16b, HLZ17a, HLZ17b, HLZ18a, HLZ18b, HLZ19a, HLZ19b, HLZ20a, HLZ20b, Zei05]. crch [MMZ16]. Create [NW07, Str19]. Creating [BND11, Law12, Rou05, SC20]. Criteria [NS17]. crossdes [Sai05]. Crossing [wCTKL20]. crossover [Sai05]. Crowdsourced [Lee16]. crs [NR12]. CRTgeeDR [PWD17]. Cryptographic [DB16, LVV11]. CryptRndTest [DB16]. ctc [LJ06]. ctmcd [Pfe17]. Cubes [Dus18]. CUDA [Fer10, KF16]. cudaBayesreg [Fer10]. Curve [GKZK16, PFMCFC18, RBS13]. Curves [BBTH17, Ran06]. Customizable [Her11]. Cut [LRMLLCS17]. Cut-off [LRMLLCS17]. cvcrand [YLGT19]. Cytometry [LH06].

D [MSC14]. D3 [KLC19]. DAG [Xie05]. Daily [AEB17]. Darwin [dL01]. Darwin/X11 [dL01]. Data [AHLL19, AB17, AE17, AEB17, Arn17, AdlF18, BGBV20, BBD12, BT15, Bec16, BTH^{+}17, Biv01, Bra02, Bro16, CI17, CA19, Cha20, Dam09, DBB^{+}16, DSU18, DYB02, Dur06, EO03, Fay10, FT19, FKP19, FLO20, GGHB18, Gen02, GJP09, GSHB14, Han17, HR07, HGF14, Hol06, HWW12, Hug14, IPRT19, IW17, KSLS20, KG06, KZDZ18, LOM20, LMK07, LH06, LZN^{+}20, Lee16, LGSG18, LDC20, LP14, Lum03, MMS16, MH17, MSC14, NtP12, NtGP17, OK07, ORVT15, OSS^{+}17a, PR17, Peb18, Pen06, Pfe17, Poi11, PT07, PWD17, PKH17, Qia14, RC12, SB10, SHR17, SJB20, SJLD^{+}17, Sie14, Sól10, Sou11, SC20, SHG14, SB19, SS16, SD20, Sze18, TVMZ10, VAD20, VP14, Vil17, VFB16, Wal16, Wil09, WCAR16, Wri16, XY19, ZB20, KLC19, Kom06, Mur05b, PB05, WAD^{+}19]. Data [LHKB17]. Database [BB18, Roo16]. Databases [MRS17, Rip01c, WW06]. Dataframes [Oh14]. Datasets [CC15, CT06, GSBT^{+}20, MW13, Mar02, HUC13]. Dataverse [Lee14]. Date [GP04, RH01]. Day [MZB^{+}18]. dclone [Sól10]. dCovTS [PF16]. DDF [HM20]. Dealing [Gen02]. Death [AdlF18]. debate [Poc06]. Debugging [Bra03, ML12]. DECIPHER [Wri16]. Decision [Sch04]. Deck [ICK18]. Decomposition [KO09]. Defensive [Lan17]. Delayed [Bra02]. Density [EB18, RH06, SFMR16, Dal01a]. DEoptim [ABC^{+}11]. Dependence [GW17, GSL19, Gre17]. Dependencies [BT15, Fal06]. Dependency [Oom13]. Depletion [DBB^{+}16]. Deploying [BND11]. Design [ACD^{+}16, SP15, VDK15, Sai05]. Designing [PZK06]. Designs [CCT15, DDK18, Fey16, FKP19, PZK06]. Desk [GP04, Lig02, Lig03a, Lig03b, LM05, Lig06, LF08, Sch03]. deSolve [Dug17]. Detecting [NtP12, SB10, YKT18]. Detection [ANRG18, BG20, GBB^{+}17, HM20, Hma20, IAA16]. Determination [MRS17]. Deuxièmes [SD13]. Developers [Len15]. Developing [Kap14]. Development [KMKT20, TZ09]. Deviation [HvWDM17]. dGAselID [MH17]. Diagnosis [PBCV06]. Diagnostic [LRMLLCS17, ZH02]. Diagnostics [GB19, GMBL18, Saw02]. Diagrams [Mur09, Zha06]. Did [HMZ12]. DIF [HM20]. Difference [WS13]. Differential [ABC^{+}11, DHS19, MSN^{+}16, SPS10]. Differentiation [DD18]. Diffusion [Dug17, WV14]. difNLR [HM20]. digitize [Poi11]. Dimension [Hin17, LMK07, WS11, ZZZ^{+}19]. Dimensional [ANRG18, CHS16, EO03, IPRT19, IW17, MH17, MHL^{+}20, YJBC16, HHB18, SC20]. Dimensionality [KRM18]. Dimensions [HB19]. dimRed [KRM18]. Directions [Oom13]. Disability [dLN^{+}18]. Disaggregation [SS13]. Discontinuity [CCT15]. Discovery [GCG^{+}16, OL09]. Discrete [AE11, BBC^{+}18, Fay10, HBH15, LS20, Pfe17, SKZS17, Spe17, WSK16]. Discrete-Continuous [LS20]. Discrete-Time [Pfe17, BBC^{+}18]. discreteRV [HBH15]. Discriminant [IMA16, PMZ20]. Disease [GSL19, Mor18, VAD20]. Diseases [dLN^{+}18]. Dispersion [Law12]. Dissimilarities [BW15]. Distance [BBTH17, HTM15, IMA16, MML16, PF16]. Distances [DSN17]. Distributed [AZJH12, JP18b, Cc20]. Distribution [BF14, GS16, Han15a, Hof13, IBG18, LHR06, MsHCV18, Min03, Mor20, MHPO15, WV14, WM10, Dal01a]. Distribution-free [MsHCV18]. Distributions [Abe15, AHvD09, AE11, FT19, GP08, LDB13, NB13, Pra20, RKSC06, SH20, Ste02]. diverse [GHM16]. Diversity [GHM16]. Diving [Luq07]. division [Han06b]. DNA [DYB02, RBB^{+}15, TVMZ10]. Do [Fox05]. doBy [Høj06]. Docker [BE17]. DOCSTRIP [Lun11]. Document [Kuh06]. Documents [TKM15]. doex [Cc20]. Dominance [BRS18]. Don’t [Mur12a]. Dose [OSS^{+}17a, PLH^{+}10, Ran06]. Dose-Response [OSS^{+}17a, PLH^{+}10, Ran06]. Dot [MZ18]. Dot-Pipe [MZ18]. Doubly [PWD17]. Downside [ABC18]. DPpackage [Jar07]. dr [Yoo18]. dr-package [Yoo18]. dr4pl [ALB^{+}19]. drat [AE17]. Draw [Mur12a]. Drawing [Her11, Mur09, Zha06]. dtwclust [SE19]. Dual [AMRdMR15]. Duration [BHM^{+}17, KH19]. Duration-Based [KH19]. Dynamic [Boe13, JP18b, OHC19, PGCE17, Szk16, WS11, dVdMF19, KLC19]. Dynamics [MO17]. dynr [OHC19].

e-Rum2020 [FVS^{+}20]. Early [Saw02]. Easier [KPBS09, Ray01]. Easy [ORC^{+}18]. easyROC [GKZK16]. Ecological [COGRB16, DGF08, LMK07, Pet03]. Economic [dVdMF19]. Ecosystem [CGG13]. EcoTroph [CGG13]. EDA [MMO20]. Editorial [Biv17a, Biv17b, Grü15a, Grü15b, Kan19, Kan20a, Kan20b, Law16a, Law16b, Mat19, Sar14a, Sar14b, Ver18a, Ver18b, Wic13a, Wic13b]. Educational [MD18]. Effect [BGBV20, BWM14, EHD07, NtGP17, SY19]. Effects [BS13, CCFVL20, FO16, GW17, NtP12, PGCE17]. Efficient [ZB20, Dus18]. EGARCH [Suc13]. eiCompare [COGRB16]. eiPack [LMK07]. Electrical [Han06a]. Electronic [KLN11]. Embedded [BS12]. EMD [KO09]. Emphasis [VP14]. Empirical [KO09, dCSPM19]. EMSaov [CKL17]. Emulation [GPB19, ORC^{+}18]. Encourage [NDWL20]. Energy [OOFdlFS^{+}16]. Enforce [MD18]. Enforcing [MVT19]. Engineer [DSU18]. Engineering [SORS06]. Enrichment [DBB^{+}16]. Entry [JCWL19]. Environment [DEVK19, GKZK16, RBB^{+}15, WFRM15]. Environmental [Bec16]. EpiILM [VAD20]. Epistemic [BC18]. Equality [AG20, Cc20]. Equation [JP18b, MSN^{+}16, VZ19]. Equations [DHS19, PWD17, SS18, SPS10, Yan02]. Equity [LK13, KE06]. Equivalence [Liu14]. ercv [dCSPM19]. Error [BTH^{+}17, GK17, HvWDM17, MUX20, Min03, XY19]. Errors [SB10, UPA18]. eRum [Dar18]. Escape [SWCH16]. Essential [FLSNG19].

Estimability [Len15]. Estimate [Adä19, FMMV18, LRMLLCS17, OOFdlFS^{+}16, WC17]. Estimates [BGBV20, COGRB16]. Estimating [LDM03, Pfe17, PGCE17, PMMEGM19, PWD17, WdM13, Yan02, ÇT18]. Estimation [AMRdMR15, AH10, BBC^{+}18, BTH^{+}17, BS13, CLLG08, GK17, GS14, Hin17, IAA17, KLK20, KM12, LGSG18, LS20, MMS16, MM15, Mor18, NS17, OL09, Sch06, SFMR16, SY19, Suc13, War06, RFGC20]. Estimations [PWD17, WSK16]. Estimator [ABS08, WBS06]. Estuaries [Bec16]. European [BAB^{+}17]. Eurostat [LHKB17]. eUtils [Win17]. Evaluation [ABC18, Han15b, MUX20, SG16, VB15]. Evapotranspiration [OOFdlFS^{+}16]. evd [Ste02]. Even [Ray01]. Event [OSS^{+}17b, FVS^{+}20]. everyday [GB06]. Evolution [ABC^{+}11, Scr17, VD13]. Evolutionary [GCG^{+}16]. Exact [Fay10, Hot01, Liu14, WS13, Dus18]. ExactCIdiff [WS13]. Example [LM05]. Excel [BND11]. Exchangeability [KCK^{+}20]. Existing [RdV15]. Exogenous [WC17]. Expected [CKL17, MUX20]. Experimental [SP15]. Experiments [OL09, PLH^{+}10, RBB^{+}15, WFRM15, War06]. Expert [GJP09, GJP09]. Explanations [SB18]. exploration [Mar02]. Exploratory [SB19]. Explore [Zab14, KLC19]. Exploring [FO16, ZG04]. Exposure [AE17]. expressions [Hoc19]. Extending [Fox07]. Extensible [MZ18, WFRM15]. Extensions [Scr17]. External [FLSNG19]. Externally [BF14]. Extracting [Poi11]. Extreme [LS20, Ste02]. extremes [WAD^{+}19].

FA [FLSNG19]. FA-based [FLSNG19]. Facets [Cha09]. Facilities [Lig03a]. FactoMineR [KBBH13, KBBH15]. Factor [BFK^{+}20, FO16, KBBH13, PS16, dVdMF19, HHB18]. Factor-Adjusted [BFK^{+}20]. Factorial [Fey16]. Factorization [KF16]. factorplot [Arm13]. Factors [Gau13, Gro03]. Failure [AWB17, MMS16]. Fake [LP14]. False [OL09]. Families [MRS17]. fanplot [Abe15]. Farewell [HGF14]. FarmTest [BFK^{+}20]. Fast [KSLS20, Lan17, MHR13, Dus18]. fastcmprsk [KSLS20]. Faster [LF08]. FastICA [MNT18]. fclust [FGS19]. Feature [DSN17, PMM^{+}19]. Feature-Weighted [DSN17]. Features [MH17, Peb18, Rip05a]. FHDI [ICK18]. fICA [MNT18]. Fields [BILL16, Sar13, Sch01]. FieldSim [BILL16]. filehash [Pen06]. Files [Mur02a, Mur07]. Filter [FVK^{+}13, VFB16]. Finance [Ulr13]. Financial [Alp07, Zag13]. Finding [LGE^{+}18]. Fingerprint [TVMZ10]. Fingerprinting [SMBTS19]. Finite [GL07, SFMR16, VDK15]. Fionn [Hol06]. First [Wai16]. Fit [AE11, DHS19, LLM16, Mug08]. Fitting [ARS15, Bat05, BGK18, CRS09, EHD07, Gau13, HNL12, Hug14, Mar11, Pap18, Ran06, RSA10, ZB20, dCSPM19, HB19]. fitzRoy [NDWL20]. Fixed [BK19, Dja19, HvWDM17, PGCE17]. flan [MDDY17]. Flexibility [BKvB^{+}17]. Flexible [Han07b, QGC19, RSR14, VP14]. FlexMix [GL07]. FLIM [HGF14]. Flow [LH06, VFB16]. fMRI [Mar02, Sto14, PT07]. Fonts [Mur04, MR06, Qiu15]. Force [BBTH17]. Force-Distance [BBTH17]. Forecast [WRR18]. ForecastComb [WRR18]. Forecasting [BACMÁK17, DY16, FRG^{+}11, MFCR19, Mey02, Suc13, WAD^{+}19]. Foreign [Adl12, Mur02a]. Forests [GPTM15, IK07]. Forge [TZ09]. Format [Kuh06]. FORTRAN [Lan01a]. Forwards [BCD^{+}17, Tur20]. Foundation [Hor13a, Hor13b, Hor14, Hot16, Hot17a, Hot17b, Hot18a, Hot18b, Hot19a, Hot19b, Hot20a, Hot20b, MH14, MH15, Plu15]. Fourier [BECDN17]. fourierin [BECDN17]. Fractional [BILL16, ICK18, MO20, Ben05]. Frailty [HNL12]. frailtyHL [HNL12]. Frame [AMRdMR15]. Frames2 [AMRdMR15]. Framework [BWM14, SH20]. free [MsHCV18]. Friendly [NW07, Sze18]. Frontiers [Roo16]. FSL [MSLC15]. fslr [MSLC15]. ftsa [Sha13]. Function [Han15b, MMS16, RSR14, CM06]. Functional [Sha11, Sha13]. Functions [Adä19, Boe13, Bos17, wCTKL20, CE18, Dal07, Han17, Jon18b, LVV11, LLM16, OOFdlFS^{+}16, PBMSN20, VP14, WV14, Han06c]. Furniture [BB17a]. Fusing [Yan04]. futureheatwaves [AEB17]. Fuzzy [FGS19, GCG^{+}16]. FWDselect [SVMMRP16].

GA [Scr17]. Gains [CMM03]. GALT [KBBH15]. Game [BC18]. Gamma [CI17]. Gamma-Poisson [CI17]. GAMs [Woo01]. GARCH [AH10, Ott19]. GAS [ABC18]. Gauss [Han15b, HBG01]. Gaussian [BILL16, GS16, GPB19, Han15a, SFMR16]. GEE [MHR13]. geepack [Yan02]. Gender [Wai16]. genderizeR [Wai16]. GeneNet [SORS06]. General [Min03, Suc20, WFRM15]. General-Purpose [WFRM15]. General-to-Specific [Suc20]. Generalised [CR02, KBBH15, Pra20]. Generalized [BKvB^{+}17, BZ10, Han17, HM20, Hug14, LDC20, LRMLLCS17, Mar11, MHPO15, PWD17, RSA10, SS18, TF07, WB10, Woo01, WR18, XGSH13, Yan02]. Generating [Guh06, TKM15]. Generation [Dem16, LDC20]. Genes [SKI06]. Genetic [BWM14, SORS06]. genetics [War03]. Genomic [Gen02]. GenSA [XGSH13]. Geocomputing [MSB17]. Geodata [Kol19]. Geographic [Bro16, Wal16]. Geometric [DEVK19]. geoR [RD01]. geoRglm [CR02]. Geospatial [EB18].

Geostatistical [RD01, SHR17]. Get [Wic11]. Getting [Lig03a]. ggfortify [THL16]. ggmap [KW13]. ggplot2 [ALH18, KW13, MMO20, TBH17]. Giving [Hyn11]. gk [Pra20]. gldrm [WR18]. glm2 [Mar11]. glmmBUGS [BZ10]. glmmTMB [BKvB^{+}17]. glmperm [WB10]. GLMs [Arm13]. Global [Sou11, XGSH13]. GMDH [DY16]. GMDH-Type [DY16]. gnm [TF07]. Goodness [AE11]. Goodness-of-Fit [AE11]. Google [GdC11]. GPU [KF16]. GPU-Accelerated [KF16]. Gradient [MHL^{+}20, Nas20]. GrapheR [Her11]. Graphical [BVB13, MMO20, SM12, TKM15]. Graphics [HUC13, Mur02b, Mur03, Mur04, Mur05a, Mur11, ML12, Qiu15, Sch03, KLC19, MR06]. Graphs [Fal06, Her11, Law12, LC06, SM12, BCB06]. graphviz [Zha06]. grid [Mur03, Mur05a, Mur02b, ML12, Mur13]. gridGraphics [Mur15]. gridSVG [MP14]. Groovy [GD18]. Group [BZST10, LC15, Cc20, PZK06]. Groups [AZJH12, DDK18, LLM16, TDAP18]. GroupSeq [PZK06]. GrpString [TDAP18]. gset [Liu14]. gsl [Han06c]. gstat [GPH16]. GsymPoint [LRMLLCS17]. GUI [Her11, Wil09]. GUIProfiler [RdV15]. gWidgets [Ver07].

Handling [KRJ16]. Haplotype [MRS17]. Hardware [KF16]. Hazard [RSR14]. Hazards [wCTKL20, KH19, MHL^{+}20, NS17]. HCmodelSets [HB19]. hdm [CHS16]. Headers [CE18]. HEALPix [FLO20]. Health [Dam09]. Heat [Fre17]. Help [GDRF09, GP04, Lig02, Lig03a, Lig03b, LM05, Lig06, LF08, MU09, Sch03]. Helping [PKH17]. Hermite [MHPO15]. Heterogeneous [AZJH12]. Heteroscedastic [MMZ16, Vil17]. hexView [Mur07]. hglm [ARS15, RSA10]. HHG [BHH18]. hiddenf [FO16]. Hierarchical [RSA10]. High [AKU17, CHS16, HHB18, HB19, IPRT19, IW17, MH17, MHL^{+}20, MSC14, RH06, YJBC16]. High-Density [RH06]. High-Dimensional [CHS16, IW17, MHL^{+}20, YJBC16, HHB18]. High-Performance [AKU17]. Higher [Bra05, LMK07]. Higher-Dimension [LMK07]. Hilbert [KO09]. Hint [WW07]. hoa [Bra05]. hospital [WBS06]. Hosting [AE17]. Hot [ICK18]. HRM [HHB18]. Huge [LJ06]. Human [Zab14]. Hurricane [AE17]. hwriter [PH09]. Hybrid [PBMSN20, Scr17]. hyper2 [Han17]. hypergeo [Han15b]. Hypergeometric [Han15b]. Hypothesis [FFW13, Liu14, Vil17].

I/O [AKU17]. ICSOutlier [ANRG18]. Identical [LQ18, Mur08]. Identification [HP19]. Identifying [BT15, SKI06, VZ19, WAD^{+}19]. ider [Hin17]. idmTPreg [AdlF18]. IDSpatialStats [GSL19]. II [DDC07, Lei03]. III [LAL18]. Illness [AdlF18]. Illumina [DSTT06]. Image [FR06a, SH06, SL17]. Images [Mur11, MSC14]. Imbalance [LMT14]. Implementation [CI17, GBB^{+}17, KSLS20, MHR13]. Implementing [Dug17, Lun11, O’N15]. Importance [GB20]. Improved [MNT18, Oom13, PHL02]. Improvements [TLP07]. Improving [Arm13]. Impulse [Adä19]. Imputation [BBACK18, HW20, ICK18, MRS17, MBB17, QGC19, SS18]. ImputeRobust [SS18]. imputeTestbench [BBACK18]. imputeTS [MBB17]. Including [FT19]. Increments [HGF14]. Independence [BHH18, MsHCV18]. Independent [AHLL19, DDK18, FRL^{+}16, LQ18, Cc20]. Indicator [Suc20]. Individual [BGBV20, VAD20]. Individual-Level [BGBV20, VAD20]. Indoor [SMBTS19]. Inequality [BRS18]. Infectious [GSL19, VAD20]. Inference [Bra05, CCT15, CGLN18, COGRB16, Ger16, Jar07, LK15, MQ06, MDDY17, SL17, WB10]. Inferences [DSH07, LMK07]. Inflated [LK15, BKvB^{+}17]. Inflation [Gro03]. influence.ME [NtP12]. Influential [NtP12]. Information [FLSNG19, Guh06, KM08, MK19, MUX20, NS17, Zag13]. informed [ÇT18]. InfoTrad [ÇT18]. Innovations [AH10]. Installing [Rip01b]. Instance [CCY^{+}17]. Instructor [KLM20]. Insurance [BDH17, GT14]. integrals [BECDN17]. Integrated [BVB13]. Integrating [CT06, Dur06, MSB17, Mur03]. Integration [CC19, GD18]. Intel [YR03]. Interacting [Pen06]. Interaction [Fey16, FO16]. Interactions [LGE^{+}18]. Interactive [GKZK16, Hah17, JM16, MSC14]. Interactively [ZG04]. Interdependent [VZ19]. Interesting [SKI06]. Interface [Adl12, CE18, THL16]. Interfaces [NW07]. Internationalization [Rip05a]. Interpolation [GPH16]. Interval [AB17, GSHB14, LGSG18, You16]. Interval-Censored [GSHB14, AB17]. Intervals [CDLR09, Fay10, MK19, WS13]. Intrinsic [Hin17]. Introducing [DGF08, Han06b, Han07a, Han05, Han06c]. Introduction [BE17, BACMÁK17, CGG13, Fei08, GMBL18, GB20, LK06, RdV15, SG16, Sch03, Ver07, WLLS08]. Introductory [Lia07, Ver05]. Inventory [SN18]. Inventorymodel [SN18]. Inverse [GS16, GS14, SS16]. Inversion [SHR17]. Investigate [LL09]. Investigate-Various-Modes-of-Convergence [LL09]. investr [GS14]. iotools [AKU17]. ipft [SMBTS19]. ipred [PHL02]. ipwErrorY [SY19]. IRT [BB16, BKPW20]. Islands [Scr17]. Isochrones [VD13]. IsoGene [PLH^{+}10].

IsoGeneGUI [OSS^{+}17a]. Item [PS16, KLWD18].

Java [Hol06, GD18, Mur05b]. JavaScript [GD18]. jobs [Xie05]. Johansen [WBS06]. John [Lia07]. Joint [QGC19]. jomo [QGC19]. Journalism [Alp07]. JRuby [GD18]. jsr223 [GD18]. Jython [GD18].

k-sample [BHH18]. Kernel [HR07, SJO^{+}20, WSK16]. Key [AL16]. keyplayer [AL16]. KNN [MFCR19]. Kotlin [GD18]. KSPM [SJO^{+}20]. Kuhn [LS20].

L [GF19]. L-BFGS-B [GF19]. Label [MLG^{+}17, RdC18]. Lag [JP18b, Sar13]. Land [EZ20, OOFdlFS^{+}16]. Landscape [SB19]. Language [Arn17, GKZK16, Tho06, YKT18]. Languages [GD18]. Large [BHH18, CDLR09, DBB^{+}16, Han07a, Dal01a]. Large-scale [DBB^{+}16]. Lasso [CCY^{+}17, ZB20]. Latent [BB16, JA18, VCQ15]. LatinR [AdSQ18]. Lattice [Sar02, SJLD^{+}17]. Lattice-Based [SJLD^{+}17]. Laying [GCGG04]. Lazy [Rip04]. lba [JA18]. LeArEst [BTH^{+}17]. Learning [CJB^{+}13, LMT14, MBL18, PBMSN20, Qia14]. Least [Bat04, CFF20]. Left [LGSG18]. Length [BTH^{+}17, LZN^{+}20, WBS06]. Let [HUC13]. Level [BGBV20, VAD20, QGC19]. Levels [Gau13]. Lexical [KBBH15]. lfe [Gau13]. Library [Adl12]. Life [SP15]. Likelihood [Ago01, Bra05, Han17, LGSG18, MMN17, HNL12]. Limit [KLN11]. Line [Mug08]. Linear [Bat05, BKvB^{+}17, BZ10, CR02, FFW13, GDF18, Gau13, HGF14, Hug14, IAA17, JP18a, KM12, Kos08, Mar11, MO20, MKRPC19, OHC19, PMZ20, Rah18, RSA10, WB10, WR18, YJBC16, Zei01, WV06]. Lines [Mur04]. Linguistic [DSN17]. Linkage [MVT19]. Linking [BB18]. Links [Koe06, VZ19]. Linux [YR03]. Lists [BFHM11]. Literate [Rou05]. Liu [IAA17]. liureg [IAA17]. live [SB18]. lmeSplines [Bal03]. lmridge [UAA18]. Loading [Rip04]. Local [Adä19]. Locating [AL16]. Location [AG20]. Locations [Wie17]. Log [GBB^{+}17]. Logistic [ALB^{+}19, CCY^{+}17, HM20, Jur12]. Lognormal [NB13]. Logs [KRJ16]. Long [GW17, Xie05]. Long-Term [GW17]. Longitudinal [IW17, LDM03, XY19]. Loop [LF08]. Loss [GP08, KLWD18]. Low [ANRG18, Jur12, SC20]. Low-Dimensional [ANRG18, SC20]. Low-memory [Jur12]. LP [Pal18]. lpirfs [Adä19]. lspartition [CFF20]. LyX [Gor08].

Mac [dL01]. Machine [Qia14]. Machines [FKW^{+}18, Mey01]. Macintosh [Iac01]. Macros [Lum01]. made4 [CT06]. magic [Han05]. Mailing [BFHM11]. Maintaining [Rou05]. Maintenance [Gil04]. Make [LM05, LF08]. Making [Ray01, TOLS06, BCB06]. Mallows [SCLV20]. Manage [VD13]. Management [LMK07, Lig03b, Rip05b, Tie01, Tie03]. Managing [GDGJC20]. Manipulation [HBH15]. Manly [ZM17]. ManlyMix [ZM17]. MANOVA.RM [FKP19]. Manual [SC20]. Manuals [Lig03a]. Many [Bra03, Gau13]. MapGAM [BGBV20]. mapmisc [Bro16]. Mapping [BGBV20, EO03, RH06, Sou11, WW06, WG10]. Maps [Bro16, DBB^{+}16, Fre17]. Marginal [Tou16]. Marked [HTM15]. Market [Wie17]. Markets [Zag13]. Markov [BBC^{+}18, Pfe17, SM12, Sar13, Spe17]. MARSS [HWW12]. Matches [MVT19]. Matching [CA19, Fay10, Han07b, JCWL19, van14]. Mathematical [Lig02]. mathematics [Han05]. Matrix [BR20, KF16, MUX20, MHR13]. maxent [Jur12]. Maximally [HL02]. Maximum [MMN17]. MCI [Wie17]. mclust [SFMR16]. MCMC [BZ10, PBCV06]. MCMCpack [MQ06]. MCSIMEX [LK06]. mctest [IAA16]. MDFS [PMM^{+}19]. MDplot [MO17]. Mean [CKL17]. Means [DSH07, Cc20, WS11]. Measured [BTH^{+}17]. Measurement [PMH16, UPA18, XY19]. Measures [Chr18, FKP19, HHB18, MML16]. Measuring [WG10, Zag13]. Mechanical [Lee16]. Medicine [RBK18]. Meeting [BAB^{+}17, BSZWW14]. Melting [RBS13]. Membership [Sch06]. Memory [Tie01, ZB20, Jur12]. Memory- [ZB20]. MessagePack [CE18]. Meta [Bea16, Sch07, Sto14, Sch07]. Meta-Analysis [Bea16, Sch07, Sto14]. Meta-Regression [Bea16]. Metadata [Car06]. metaplus [Bea16]. Metapopulation [Dug17]. Method [GF19, KSLS20, VB15, Dus18]. Methodology [Zab14]. Methods [BBACK18, Can02, CGLN18, CDT04, Chr18, DDC07, HR07, Jon18b, Mey02, RdV15, Suc20, Wai16, PB05]. Metrics [CHS16, HTM15]. mfp [Ben05]. mgcv [Woo01]. MGLM [KZDZ18]. micompr [FRL^{+}16]. Microarray [BWM14, DYB02, FR06a, OL09, OSS^{+}17a, PLH^{+}10, Smi03, War06]. Microarrays [Saw02]. Microscopy [SH06]. Microsoft [YR03]. Midrange [BF14]. milr [CCY^{+}17]. MINimal [OGP17]. minimization [Dus18]. Mining [BVB13, CI17, Fei08, ORVT15, Wil09]. minval [OGP17]. Misclassified [SY19]. Missing [HGF14, KM08, LGSG18, MBB17, PWD17, RPO^{+}19]. Missingness [XY19]. mistr [SH20]. miWQS [HW20]. Mixed [Bat05, BKvB^{+}17, BZ10, DHS19, Hug14, NtP12, SM12, Sze18, Vil17]. Mixed-Type [Sze18]. mixedsde [DHS19]. Mixing [Lei02]. mixtox [ZC16]. Mixture [FT19, GL07, SH20, SFMR16, ZC16, ZM17].

Mixtures [AHvD09, PBMSN20]. mldr [CC15]. mle.tools [MMN17]. mlr [PAC^{+}17]. mmpf [Jon18b]. mmpp [HTM15]. Mode [KO09]. Model [ALB^{+}19, AEB17, AH10, Arn17, AdlF18, BGK18, CI17, CRS09, Dug17, FR06a, FR06b, GDF18, GB19, GSHB14, HGF14, IAA17, IBG18, KH19, LZN^{+}20, LS20, MHL^{+}20, ORC^{+}18, PR17, PGCE17, PS16, RPV05, SS18, SCLV20, SFdC18, SB18, Tou16, WV14, ZB20]. Model-Agnostic [GB19]. Model-Based [FR06b, FR06a]. Modeling [BKvB^{+}17, Bür18, Fal06, GP08, GJP09, Mor20, OHC19, WAD^{+}19, ZM17]. Modelling [BKPW20, GL07, MHPO15, Pet03, QGC19, Rib07, Sar13, VAD20]. Models [ARS15, ABS08, AWB17, AB17, BB16, BBC^{+}18, Bat05, BS13, BB17b, BZ10, CLLG08, CR02, EHD07, FFW13, GW17, GDF18, Gau13, Ger16, GMBL18, GZLW09, HNL12, Han17, HM20, HB19, HWW12, Hug14, JM17, JP18b, KM12, Kos08, LDM03, Mar11, MDDY17, MKRPC19, Mug08, NS17, NtP12, OOFdlFS^{+}16, OSS^{+}17b, Pap18, RSA10, SM12, SJO^{+}20, SFMR16, SVMMRP16, Suc13, SD20, TF07, VZ19, Vil17, WB10, WdM13, WR18, dVdMF19, Ott19]. Modern [MsHCV18, Wic10]. Modes [LL09]. Module [DBB^{+}16]. Molecular [DBB^{+}16, Guh06, MO17]. Monte [Jon18b]. Monte-Carlo [Jon18b]. mopa [IBG18]. mosaic [PKH17]. MoTBFs [PBMSN20]. Motion [MO20]. move [BCB06]. mpoly [Kah13]. MRCV [KB14]. MRI [Mar02]. mtk [WFRM15]. mudfold [BKPW20]. Multi [Bos17, HHB18, RdC18, SHR17]. Multi-factor [HHB18]. Multi-label [RdC18]. Multi-Objective [Bos17]. Multi-scale [SHR17]. Multi-type [SHR17]. Multidimensional [ORC^{+}18, PMM^{+}19]. Multilabel [CC15, PAC^{+}17]. Multilevel [Bür18]. Multinomial [Jur12, Tou16]. Multiplatform [Her11]. Multiple [BFK^{+}20, BHW02, CS16, CCdUÁ14, wCTKL20, CCY^{+}17, FT19, Fre17, Gau13, HW20, KBBH13, KB14, LLM16, LS20, OSS^{+}17a, QGC19, SS18, TLP07, VS15]. Multiple-Instance [CCY^{+}17]. multipleNCC [SS16]. Multisource [KCK^{+}20]. Multistate [ABS08]. Multivariable [Ben05]. Multivariate [BGK18, CT06, Dal07, FRL^{+}16, FFW13, FKP19, Han08, Han15a, Hof13, HWW12, HBG01, JGB14, Kah13, KZDZ18, KGZ14, LDC20, MMS16, MMH09, PR17, SKZS17, SJLD^{+}17, SC20, TVMZ10, Vil17, WM10, KLC19, WAD^{+}19]. multtest [TLP07]. Murtagh [Hol06]. Mutation [MDDY17]. MVN [KGZ14]. mvna [ABS08]. mvtnorm [MMH09].

Naive [Mey02]. Name [Mur12b, Tie03]. namedCapture [Hoc19]. Names [Wai16]. Naming [Båå12]. National [VFB16]. Natural [Arn17, YKT18]. Navigating [SNG18]. NCBI [Win17]. ncpen [KLK20]. Need [WW07]. Negative [KF16]. Nelson [ABS08]. Nested [BS13, SS16]. NetCDF [MW13]. Network [BFHM11, CGLN18, Chr18, CGG13, DY16, GBB^{+}17, JM16, MBL18, TBH17]. Networks [AL16, DGF08, GF10, JM17, LC06, PBMSN20, SORS06, CC19, Han06a]. NetworkToolbox [Chr18]. Neural [DY16, GF10]. neuralnet [GF10]. Neuroimage [MSC14]. News [Hor13a, Hor13b, Hor14, Hot16, Hot17a, Hot17b, Hot18a, Hot18b, Hot19a, Hot19b, Hot20a, Hot20b, MH14, MH15, Bio13a, Bio13b, Bio14, The14a, The15a, The15b, Tea16, The16a, Tea17a, Tea17b, Tea18a, Tea18b, Tea19, The19a, The19b, Tea20, Tur20]. ngspatial [Hug14]. NHANES [VP14]. Niche [Car08, Fox05, Han08, Lum01, Lum04b, Ven01, Ven02]. NlinTS [Hma20]. nlme [HvWDM17]. NMF [KF16]. nmfgpu4R [KF16]. NMMAPSdata [PW04]. Noise [MLG^{+}17]. NoiseFiltersR [MLG^{+}17]. Noisy [SKZS17]. Non [Jar07, KLK20, KF16, Liu14, LQ18, MR06, Mur08, PFMCFC18, Vil17, WV06, Dal01a, RFGC20]. Non- [Jar07]. Non-Central [Liu14, Dal01a]. Non-Convex [KLK20]. Non-Identical [LQ18, Mur08]. Non-linear [WV06]. Non-Negative [KF16]. Non-normal [Vil17]. non-probability [RFGC20]. Non-Standard [MR06, PFMCFC18]. noncentrality [Dal01a]. Nonlinear [LCT20, OHC19, TF07]. Nonparametric [AG20, AE11, BKPW20, BHH18, Fey16, LBCFV16, NR12, RSR14, SKZS17]. NonProbEst [RFGC20]. Normal [BF14, DSH07, MMH09, WM10, You16, Cc20, Vil17]. Normality [KGZ14]. Normed [Han06b]. Nowcasting [dVdMF19]. np [HR07]. npordtests [AG20]. nsROC [PFMCFC18]. NTS [LCT20]. Nuclear [MRS17]. Null [AE11]. Number [Fox05, MH17]. Numbers [Han07a, LL05]. Numerical [Han15b, MMH09, Poi11, WFRM15].

O [AKU17]. Object [CL01a].

Object-Oriented [CL01a]. Objective [Bos17]. Objectives [Kos08]. Objects [Mur08]. Observational [Han07b, NtGP17]. Observations [FRL^{+}16]. Occurrence [Mor20]. ODE [Gra14]. odfWeave [Kuh06]. off [LRMLLCS17]. oligoMask [BWM14]. oLongForm [Oh14]. Omegahat [CL01b]. ONCOlogy [ACD^{+}16]. One [CDT04, DDK18, LK15, MVT19, WS11]. One-table [CDT04]. One-to-One [MVT19]. One-Way [DDK18]. onewaytests [DDK18]. onion [Han06b]. Online [TVMZ10, WCAR16]. Open [GZLW09, KY15, LHKB17, TOLS06, Kuh06]. Open-Channel [KY15]. openair [RC12]. openEBGM [CI17]. OpenLand [EZ20]. OpenMx [PS16]. OpenStreetMap [ES13].

Opinion [GJP09]. Optimal [CPS15, Han07b, LRMLLCS17, SP15, Szk16, WS11, HY18]. Optimisation [Scr17]. optimising [WV06]. Optimization [Bos17, GF19, MVT19, Pal18, XGSH13, ZZZ^{+}19]. Optimizers [Nas20]. Optimizing [DSN17]. optimParallel [GF19]. Options [KB14]. Optmatch [Han07b]. Order [Bra05, KLN11]. Ordered [AG20, MMS16]. Ordinal [GMBL18, LP14, SJB20]. ordinalClust [SJB20]. Organization [Fox09]. Organizing [Bec16]. Oriented [CL01a]. orthoDr [ZZZ^{+}19]. Orthogonal [PGCE17]. Orthogonality [ZZZ^{+}19]. OrthoPanels [PGCE17]. osmar [ES13]. other [Hoc19]. Outcome [SY19]. Outlier [ANRG18]. outliers [Kom06]. Output [AEB17, Mur03, ORC^{+}18, PBCV06].

Package [Abe15, Adä19, Ago01, ABS08, AWB17, AG20, AHLL19, AL16, AEB17, ACD^{+}16, AMRdMR15, ABC18, BKPW20, BBC^{+}18, BF14, Bea16, Bec16, BBACK18, BHM^{+}17, BBTH17, BZST10, BS13, BR20, BACMÁK17, BFK^{+}20, Bra05, BHH18, BGK18, BILL16, Bür18, BW15, CA19, Can02, CRS09, CCdUÁ14, CCT15, CMW11, CPS15, CC15, CCFVM17, CDT04, CKL17, CR02, CDLR09, CJB^{+}13, DY16, DDK18, Dal01b, DB16, Dem16, DHS19, Dja19, DEVK19, DGF08, DSN17, DDC07, Dug17, DSTT06, ERK16, EHD07, FRL^{+}16, Fal06, FLSNG19, FGS19, FVK^{+}13, FFW13, FO16, FKP19, FLO20, GK17, GCG^{+}16, GPTM15, GF19, Gil04, GSL19, GDGJC20, GSBT^{+}20, GB19, Gra14, GS14, Gre17, GMBL18, GB20, GL07, GHM16, HNL12, Han07a, Han15b, Han17, HHB18, HR07, HvWDM17, HTM15, Hma20, HB19, HGF14, HP19]. Package [HMZ12, Hug14, IPRT19, ICK18, IAA16, IAA17, IW17, IMA16, IBG18, JGB14, JA18, JM17, Jon18a, JCWL19, Jur12, KSLS20, KO09, KZDZ18, KLK20, KGZ14, Kos08, KBBH13, KBBH15, KB14, KH19, Kuh06, LL09, LOM20, LHKB17, LGE^{+}18, Lan01b, Law12, Lec03, LZN^{+}20, Lei03, LHR06, Len15, LDC20, Lig03b, LLM16, LBCFV16, Liu14, LK15, LCT20, LS20, LP14, LRMLLCS17, LE18, Lum04a, LMT14, MFCR19, MDDY17, MMN17, MO20, MHR13, MKRPC19, MMS16, MH17, MW13, MsHCV18, Min03, MM15, MSN^{+}16, MLG^{+}17, MML16, MHPO15, MHL^{+}20, Mur02b, Mur07, MP14, Mur15, NS17, NB13, NDWL20, NR12, Nie15, Nor16, NP15, OS18, OK07, OL09, ORVT15, Ott19, OHC19, Pal18, Pap18, PR17, PH09, PW04, Pen06, PBMSN20, Pfe17, PGCE17, Poi11, PT07, PWD17, PLH^{+}10, Pra20, PAC^{+}17]. Package [PKH17, Qia14, QGC19, RPO^{+}19, RD01, RdC18, RSA10, Roo16, RdV15, SN18, SG16, SS18, SMBTS19, SE19, SB10, SHR17, SHW18, SORS06, SJO^{+}20, Sch07, Sch16, Scr17, SJB20, SP15, SVMMRP16, SCEG19, Sha11, Sha13, SY19, SJLD^{+}17, SP19, SNG18, DD18, SCLV20, Sou11, SFdC18, SC20, SHG14, SL17, TDAP18, TLP07, TD13, Tou16, TF07, UAA18, VZ19, VD13, VB15, VCQ15, Vil17, WV14, Wal16, WS13, WC17, WSK16, War03, War06, WRR18, WM10, WCAR16, WR18, XGSH13, XC08, XY19, Yan02, Yee08, YJBC16, YKT18, You16, YLGT19, Zab14, ZB20, ZC16, ZM17, dVdMF19, dCSPM19, van14, BECDN17, CC19, Cc20, ÇT18, CGG13, Dal02, DSU18, FMMV18, Han05, Han06b, Han06c, HY18, Høj06, HQ16, JP18a, KLWD18, MK19, Mar02, Mev06, Mug08, OGP17, PFMCFC18]. package [PMMEGM19, RPV05, RFGC20, Sai05, Scr04, VFB16, WBS06, WAD^{+}19, Win17, Yoo18]. Packages [AE17, Bat03, Bra02, BKvB^{+}17, CL01b, DBB^{+}16, Fox07, GDRF09, KRJ16, LJ06, NEB^{+}20, Plu11, Rip04, Rip05b, Rou05, SB18, SB19, THL16, ZG04, Hoc19, YR03]. Pages [GDRF09]. Pairwise [VS15]. pampe [VB15]. Panel [PGCE17, VB15]. PanJen [JP18a]. Parallel [AZJH12, GF19, KPBS09, Yu02]. Parallelisation [Scr17]. Parallelization [Boe13]. Parameter [ARA^{+}14, ALB^{+}19, ORC^{+}18, Dal01a]. Parameterized [VS15]. Parameters [Pfe17, SHR17]. Parametric [CDLR09, Koe06, SJO^{+}20, Jar07]. Pareto [Roo16]. Parsimonious [CLLG08]. Part [Lei02, Lei03]. Partial [Gre17, Han17, JGB14]. Partitioning [CFF20]. Partitioning-Based [CFF20]. Party [SHZ09]. Pathways [GCGG04, CC19]. Pattern [BACMÁK17]. Patterns [SKZS17, WW06]. PDF [MR06]. pdp [Gre17]. Peaks [Rib07]. pedigree [Zha06]. Penalized [KLK20]. Penalties [PMZ20]. Penalty [CCY^{+}17]. Peptides [ORVT15, ORVT15]. Perform [OC07]. Performance [AKU17, LK13]. Performing [CPS15, LBCFV16, SP19]. Period [Nie15]. Permutation [WB10, YLGT19]. Persistent [O’N15]. Person [God13, GE14]. Perspective [God13, GE14]. Perspectives [Zab14]. PGEE [IW17]. Pharma [Ric18]. Pharmacometric [GSBT^{+}20]. Phase [Gra14]. phaseR [Gra14]. Phenotypic [DD18]. Phillips [SP19]. physical [FVS^{+}20]. PIN [Zag13]. Pioneering [SCEG19]. Pipe [MZ18]. PITCHf [Sie14]. PITCHf/x [Sie14]. pitchRx [Sie14]. PivotalR [Qia14]. Plagiarism [BG20]. Plane [Gra14]. Planning [LE18]. Platform [GD18, Pet03]. Players [AL16]. Plot [JW06]. plotrix [Lem06]. Plots [ALH18, Gre17, GB20, Lig02]. pls [Mev06]. Plug [Fox07]. Plug-In [Fox07]. PMML [GZLW09]. Point [BK19, EB18, HTM15, LOM20, LRMLLCS17, Mor20, Rou01]. Poisson [CI17, LDC20]. Policy [Dam09]. Polish [BSZWW14]. politeness [YKT18]. Polynomials [Kah13, Ben05, Han08]. Popular [THL16]. Population [BHM^{+}17, DD18, VDK15]. Portable [Plu11]. Portfolio [Pal18]. PortfolioOptim [Pal18]. Portfolios [LK13, KE06]. Porting [Iac01, dL01]. Positioning [SMBTS19]. PostGIS [BB18]. PostScript [MR06]. POT [Rib07]. Potential [MSN^{+}16]. Poverty [BRS18]. powder [WW06]. PPCI [HP19]. practice [GB06]. PracTools [VDK15]. Preclinical [PL05]. predictability [KLWD18]. Predicting [OSS^{+}17b, dVdMF19]. Prediction [BS13, IKR10, Jon18b, Wai16]. Predictions [SB18]. Predictors [Kos08, NR12, PHL02]. Preferences [Roo16]. Preprocessing [Lee16, MLG^{+}17]. Presence [LGSG18]. Presentation [Arm13]. Presenting [Coo11]. Preserving [AWB17]. Primer [Dal01b]. Principal [SG16]. Prinsimp [ZHC^{+}14]. Prior [MK19]. Probabilistic [FRG^{+}11, GW17]. Probabilities [HBG01, MMH09]. Probability [GS16, LDB13, SS16, ÇT18, RFGC20]. Probes [BWM14]. Probit [WdM13]. Problems [CS16, CCdUÁ14, Mar11, MVT19, SN18]. Procedure [SP19]. Procedures [GBB^{+}17, You16]. Process [GGHB18, GPB19, Scr04]. Processes [BKPW20, HTM15, Mor20]. Processing [AZJH12, Arn17, Kom06, VP14, Wic10]. Profile [BR20, LGSG18]. profileModel [Kos08]. Profiling [Kos08, RdV15]. Progeny [HQ16]. progenyClust [HQ16]. Program [VB15]. Programmer [Car08, Fox05, Han08, Lum01, Lum04b, Ven01, Ven02]. Programming [CL01a, GD18, Lan17, Rah18, Rou05, Szk16, WS11]. Progressive [AdlF18]. Project [Fox09, Bio13a, Bio13b, Bio14, The14a, The15a, The15b, Tea16, The16a, Tea17a, Tea17b, Tea18a, Tea18b, Tea19, Tea20]. Projected [Fre17]. Projection [HP19]. Projections [Adä19, IBG18, SC20]. ProjectManagement [GDGJC20]. Projects [GDGJC20]. Propagation [SHW18]. properties [Han06a].

Proportional [KH19, MHL^{+}20, NS17]. Proportions [WS13]. Proposal [Bar18]. Protein [Sch06]. Provenance [Nas20]. Providing [GF19, WV14]. pseval [SG16]. PSF [BACMÁK17]. Pstat [DD18]. Psychometric [Chr18, LLM16]. Psychometrics [MH07, MD18]. PubMed [GG02]. Pure [MHR13]. Purpose [WFRM15]. Pursuit [HP19]. Putting [BC18]. puzzle [GSBT^{+}20].

Q [Zab14]. QCA [OS18, TD13]. qcc [Scr04]. QGIS [MSB17]. qmethod [Zab14]. QPot [MSN^{+}16]. Qtools [Ger16]. Quadratic [Rah18]. Quadtree [LOM20]. Qualitative [TD13]. Quality [RC12, Saw02, Scr04]. QuantifQuantile [CPS15]. Quantifying [GSL19]. Quantile [ALH18, CPS15, Ger16, HW20, LBCFV16]. Quantile-Quantile [ALH18]. Quantitative [BB17a, EZ20]. Quantities [KH19]. Quantization [CPS15]. quantreg.nonpar [LBCFV16]. Quasi [MSN^{+}16]. Quasi-Potential [MSN^{+}16]. Querying [GG02]. Queueing [JM17, JM17]. Queues [O’N15]. quickpsy [LLM16].

R [BND11, BSZWW14, BAB^{+}17, BKJ^{+}18, BPC^{+}20, KLM20, Lei07, LG13, Lia07, Abe15, Adä19, ABS08, AWB17, ALH18, Alp07, AG20, AHLL19, AL16, AB17, AEB17, AZJH12, ABC18, AKU17, Båå12, BB16, BND11, BKPW20, BDH17, BBC^{+}18, Bar18, BG20, BECDN17, Bat04, Bat05, BF14, BK19, Bea16, Bec16, BBACK18, BBTH17, BS12, BS13, BR20, Boe13, BE17, BACMÁK17, BFK^{+}20, BVB13, Bra05, BHW02, BGK18, BH05, BB18, Bür18, BW15, CI17, CA19, Can02, Car08, CRS09, CCdUÁ14, CCT15, ÇT18, CL01a, CL01b, Cha09, Cha20, wCTKL20, CPS15, CC15, CLLG08, CCFVM17, CCFVL20, CE18, CKL17, Chr18, CDLR09, CGG13, Use13, DY16, DDK18, Dal01b, Dal02, Dam09, DB16, Dem16, DZK13, Dja19, DEVK19, DSN17, DSU18, DYB02, Dug17, DSTT06, Dur06, ERK16]. R [ES13, FRL^{+}16, Fei08, FLSNG19, FGS19, Fey16, FMMV18, FVK^{+}13, FKW^{+}18, Fox07, Fox09, FRG^{+}11, FO16, FKP19, FLO20, GK17, GB06, GDF18, GCG^{+}16, GDM^{+}17, GGHB18, GPTM15, GF19, GdC11, GT14, Gil04, GD18, GSL19, GKZK16, GDGJC20, GSBT^{+}20, GB19, GDRF09, Gra14, GS14, Gre17, GP04, GL07, GPB19, GHM16, Guh06, Hah17, HJKK19, Han05, Han06b, Han06c, Han07a, Han08, HHB18, Her11, Hin17, HY18, Hma20, Hoc19, HB19, HUC13, HP19, HMZ12, Hor13a, Hor13b, Hor14, Hot01, HBG01, HL02, Hot16, Hot17a, Hot17b, Hot18a, Hot18b, Hot19a, Hot19b, Hot20a, Hot20b, HQ16, Iac01, ICK18, IAA16, IAA17, IW17, IMA16, IK07, JGB14, JA18, JP18a, JM17, Jon18a, JCWL19, Jur12, Kah13, KMH20, KE06, Kap14, KMKT20, KSLS20, KG06, KRJ16, KZDZ18]. R [KLC19, KPBS09, Kol19, KY15, KGZ14, Kos08, KRM18, KH19, LL09, LVV11, LOM20, LGE^{+}18, Lan17, LDB13, LZN^{+}20, Lee14, Lee16, Lei02, LR01, LDC20, Lig02, Lig03a, Lig03b, LM05, Lig06, LF08, LLM16, LBCFV16, Liu14, LK15, LCT20, LS20, LDM03, LRMLLCS17, Lum01, Lum03, Luq07, MH14, MH15, MK19, Mar02, MQ06, MFCR19, MZB^{+}18, MDDY17, MMN17, MO20, MHR13, MKL20, MKRPC19, MMS16, MH17, MsHCV18, MM15, MSN^{+}16, Mor18, MLG^{+}17, MML16, MHPO15, MBB17, MHL^{+}20, MZ18, MSB17, Mug08, MU09, Mur04, Mur09, Mur11, Mur05b, MSLC15, NS17, NB13, Nas20, NW07, NDWL20, Nie15, Nor16, NTE14, NP15, NEB^{+}20, O’N15, OS18, OK07, OOFdlFS^{+}16, OC07, Oom13, OGP17, OSS^{+}17a, Ott19, OHC19, Pap18, PR17, PB05, PMH16, Peb18, PBMSN20, PFMCFC18, Pet03]. R [Pfe17, PGCE17, PL05, PMM^{+}19, Plu11, Plu15, Poc06, PMMEGM19, PWD17, PLH^{+}10, Pra20, PAC^{+}17, PKH17, Qiu15, RPV05, Rah18, RPO^{+}19, Ric18, RBK18, Rip01b, Rip01c, Rip02, Rip04, Rip05a, Rip05b, RdC18, Rob07, RBB^{+}15, RBS13, RdV15, RFGC20, SN18, SM12, SMBTS19, SE19, SHW18, Sch03, Sch04, Sch07, Sch16, Scr04, SJB20, SP15, SVMMRP16, Sha11, Sha13, SY19, SD13, SNG18, DD18, SPS10, Sól10, SCLV20, Sou11, Spe17, SFdC18, SC20, SHG14, SB19, Sto14, SD20, SL17, Sze18, THL16, TDAP18, The13a, The13b, The14b, The14c, The15d, The15c, The16b, The16c, The17a, The17b, The18a, The18b, The19a, The19b, TKM15, TZ09, Tie03, UPA18, UAA18, Ulr13, VP14, VB15, Ver05, VCQ15, VS15, VFB16, WV14, Wal16, WS13, WFRM15, WC17, WBS06, WSK16]. R [WRR18, WdM13, Wil09, WCAR16, Win17, Woo01, Wri16, XGSH13, Xie05, XY19, YR03, Yan04, YJBC16, dLN^{+}18, Yu02, Zag13, ZB20, ZG04, Zha06, ZC16, ZM17, dL01, dVdMF19, van14, Hol06]. R-Forge [TZ09]. R-Package [Ott19, WBS06]. R-Tcl [Dal01b, Dal02]. R-Tcl/Tk [Dal01b, Dal02]. R/Finance [Ulr13]. R/Sweave [GB06]. R2HTML [Lec03]. rainbow [Sha11]. Random [BS13, Dem16, EHD07, GPTM15, HBH15, IK07, LQ18, Sar13, Sch01, LL05]. randomForest [LW02]. Randomization [YLGT19, Sai05]. Randomized [PWD17, YLGT19]. Randomness [DB16]. Rank [AWB17, BS13, Han17, Hot01, HL02, KM12]. Rank-Based [BS13, KM12]. Rankcluster [JGB14]. Ranked [SCEG19]. Ranking [JP18a, VS15]. Rankings [JGB14]. Raster [Mur11]. Rate [OL09]. rating [KLWD18]. RatingScaleReduction [KLWD18]. Ratios [DSH07]. Rattle [Wil09]. ray [WW06]. RC [COGRB16, Dem16]. rCBA [HJKK19]. Rcgi [Ray01]. RcmdrPlugin.temis [BVB13]. rcosmo [FLO20]. RcppMsgPack [CE18]. rcss [HY18]. rdrobust [CCT15]. Reactions [OGP17]. Read [Nor16]. Reading [MW13, Mur02a]. RealVAMS [BGK18]. Record [MVT19]. RecordLinkage [SB10]. Recreational [Han05]. Recursive [VZ19]. Reduction [EO03, KRM18, ZZZ^{+}19, KLWD18]. Refer [HMZ12]. Reference [Bro16]. References [Mur10]. regimes [WAD^{+}19]. regress [CM06]. Regression [AB17, AdlF18, Bea16, BHM^{+}17, BB17b, CCT15, CFF20, CPS15, CCY^{+}17, GMBL18, GSHB14, HW20, HM20, IAA17, IKR10, JP18a, Jur12, LW02, LBCFV16, LK15, MKRPC19, MMZ16, Mug08, NR12, OSS^{+}17b, Pap18, SVMMRP16, SD20, UAA18, Woo01, YJBC16, Zei01, ZH02, WV06]. Regression-Discontinuity [CCT15]. Regressor [WB10]. Regressors [IAA16, IAA17]. regular [Hoc19]. Relationships [Mug08, Zei01, ZH02]. Removing [BWM14]. Rencontres [SD13]. rentrez [Win17]. Reparameterization [PGCE17]. Repeated [FKP19, HHB18]. Replicating [Dam09]. Report [AdSQ18, BDH17, BSZWW14, BAB^{+}17, BKJ^{+}18, BPC^{+}20, CvDMvE19, Dar18, GT14, LAL18, Ric16, Ric18, RBK18, SD13, Tve15, Ulr13, Ver17, MZB^{+}18, RBK18]. Reports [SFdC18, Str19]. Representation [SJLD^{+}17]. Reproducible [Lee14, NDWL20, RBB^{+}15, TVMZ10]. Resampling [Can02, FKP19]. Resampling-Based [FKP19]. Research [KMKT20, Lee14, TVMZ10, WCAR16]. Residual [HvWDM17, dCSPM19]. Residuals [GMBL18, WB10]. resistor [Han06a]. Resolution [MSC14]. Resources [Dur06]. Response [Adä19, Koe06, KB14, OSS^{+}17a, XY19, PLH^{+}10, Ran06]. Responses [Tou16]. Results [THL16]. Retail [Wie17]. Retrieval [LHKB17]. Retrieve [VFB16]. Retrieving [Bec16]. revengc [DSU18]. Reverse [DSU18, SORS06]. Review [Bar18, Hol06, Lei07, LG13, Lia07, RdV15]. Rfit [KM12]. rFSA [LGE^{+}18]. rggobi [WLLS08]. Rgraphviz [GCGG04]. Rica [Mor20]. Ridge [UAA18, Woo01]. Right [LGSG18]. Risk [ABC18, Mor18, OSS^{+}17b]. riskRegression [OSS^{+}17b]. Risks [AHLL19, KSLS20]. River [VFB16]. rlfsm [MO20]. rlme [BS13]. rmdcev [LS20]. Rmetrics [CMW11]. Rmpi [Yu02]. RNaviCell [DBB^{+}16]. RNetCDF [MW13]. RNGforGPD [LDC20]. rnrfa [VFB16]. roahd [IPRT19]. Robotics [DZK13]. Robust [Ago01, Bea16, BFK^{+}20, CCT15, GPB19, IPRT19, PWD17, Vil17]. RobustGaSP [GPB19]. ROC [GKZK16, PFMCFC18]. Rocker [BE17]. Rockerverse [NEB^{+}20]. role [Poc06]. Roles [HMZ12]. Rolling [JCWL19]. rollmatch [JCWL19]. ROSE [LMT14]. rotations [SHG14]. Routine [MD18]. Routines [Lan01a, MVT19]. rpanel [BCB06]. rpostgis [BB18]. rPref [Roo16]. rpsftm [AWB17]. RPVM [CMM03, LR01]. RQGIS [MSB17]. RSSampling [SCEG19]. RStorm [Kap14]. rstream [LL05]. rTableICC [Dem16]. RTextTools [CJB^{+}13]. rtip [BRS18]. Rules [Hah17]. Rum2020 [FVS^{+}20]. Running [Xie05]. RWiener [WV14]. rworldmap [Sou11]. Rwui [NW07].

S [Cha20]. S3 [Lum04b, MZ18]. S4 [Bat03, Lum04b, RKSC06, SJLD^{+}17]. sae [MM15]. Sample [OL09, War06, BHH18]. Samples [CDLR09, VDK15]. Sampling [Hof13, SCEG19]. SARIMA [SFdC18]. SAS [Dam09]. Saturation [Suc20]. Saunders [LHR06]. sbtools [WCAR16]. Scalable [KSLS20]. scale [DBB^{+}16, KLWD18, SHR17]. Scattered [OK07]. Scatterplots [Poi11]. Schema [Nor16]. Schema-on-Read [Nor16]. SchemaOnRead [Nor16]. Scheme [VS15]. Science [Cha20]. Scientists [BB17a]. scmamp [CS16]. Screens [SH06]. Scripts [NW07]. SDEFSR [GCG^{+}16]. sdpt3r [Rah18]. Search [Lan01a]. Searching [GDRF09]. Seasonal [BBD12]. segmented [Mug08]. Seismology [BH05]. Selected [HL02]. Selecting [MH17]. Selection [CRS09, GDF18, GPTM15, IKR10, MHL^{+}20, Pap18, PMM^{+}19, SVMMRP16, Yoo18, LC15]. Semi [AHLL19, Jar07, Jur12, SJO^{+}20, BBC^{+}18]. Semi-automated [Jur12]. Semi-competing [AHLL19]. semi-Markov [BBC^{+}18]. Semi-Parametric [SJO^{+}20, Jar07]. SemiCompRisks [AHLL19]. Semidefinite [Rah18]. Semiparametric [Pap18, WR18, ZZZ^{+}19]. Sensitivity [WFRM15]. separation [WV06]. Sequence [BACMÁK17, Wri16]. Sequential [Abe15, Liu14, Szk16, PZK06]. SER [LAL18]. Series [BBACK18, BR20, Hma20, LBCFV16, LCT20, MFCR19, Mey02, MML16, MBB17, ORC^{+}18, PF16, PMMEGM19, Rip02, SE19, SS13, Sha11, Sha13, WC17, HWW12, WAD^{+}19]. Service [Wie17]. Services [Kol19]. Session [BS12]. Set [SCEG19]. SetMethods [OS18]. Sets [HB19]. sfCluster [KPBS09]. sgof [CCdUÁ14]. sgr [LP14]. Shadow [DEVK19, DEVK19]. Shapes [WG10]. Sharing [GZLW09]. Shiny [CKL17, KMKT20]. ShinyItemAnalysis [MD18]. Short [DY16, LK06]. showtext [Qiu15]. Shrinker [CI17]. sided [Fay10]. siggenes [SKI06]. SimCorMultRes [Tou16]. SimCorrMix [FT19]. SIMEX [LK06]. SimilaR [BG20]. Similarity [HTM15]. Simple [Arm13, CDLR09, HTM15, LB16, Lum04b, Peb18]. Simulated [XGSH13]. Simulating [GW17, LP14, SKZS17, Tou16]. Simulation [BBC^{+}18, FT19, JP18b, LS20, Pet03, Sch01, Suc13, LL05]. Simulations [ARA^{+}14]. Simultaneous [ARS15, VZ19]. Single [Bos17, JP18b]. Single- [Bos17]. Single-Equation [JP18b]. SIRE [VZ19]. Six [KMKT20]. Size [OL09, War06]. Sizes [BHH18]. Skew [MUX20, Suc13]. Skew- [MUX20]. Slab [IKR10]. Sloth [Mor20]. Small [MM15, Mor18]. SMM [BBC^{+}18]. smoof [Bos17]. Smoothed [BGBV20]. Smoothing [RSR14]. SMR [BF14]. Snowboot [CGLN18]. snowfall [KPBS09]. SNP [Car06]. Social [AL16, BFHM11, Fox09, JM16]. Software [CG06, EZ20, GE14, MSLC15, TZ09]. Solution [BVB13]. Solver [ZB20]. Solving [SPS10]. Some [FR06b, Gen02, Scr17]. Sorted [CCFVL20]. SortedEffects [CCFVL20]. sos [GDRF09]. Source [Mur10]. Sources [Lig06]. Space [ORC^{+}18, Tie03, HWW12]. Spaced [PMMEGM19]. sparkTable [TKM15]. Sparse [Hug14, LC15, NS17]. spartan [ARA^{+}14]. Spatial [ARS15, BGBV20, Biv01, BB18, CR02, GSL19, Hug14, KW13, Mor20, Ott19, Peb18, Sar13, SHW18, WdM13, PB05]. Spatio [GPH16]. Spatio-Temporal [GPH16]. Spatiotemporal [Ott19]. spcadjust [GK17]. spduration [BHM^{+}17]. Special [Han06c]. Species [IBG18, Mor20]. Specific [Suc20]. Specification [Tou16]. Specified [Suc20]. Specifying [HB19]. Spectrum [KO09]. Speed [BKvB^{+}17]. Spell [Fox05]. Spelling [HM11]. spGARCH [Ott19]. Spherical [FLO20, OK07]. SpherWave [OK07]. Spike [IKR10]. spikeslab [IKR10]. spinifex [SC20]. Splines [NR12]. Split [BHM^{+}17]. Split-Population [BHM^{+}17]. Splitting [BHM^{+}17]. spMC [Sar13]. Sports [NDWL20]. Spread [GSL19]. spup [SHW18]. squared [Dal01a]. Squares [Bat04, CFF20, CKL17]. SRCS [VS15]. ssize [War06]. ssize.fdr [OL09]. Stable [ALB^{+}19, MO20]. Stacks [O’N15]. Standard [GZLW09, HvWDM17, MUX20, MR06, PFMCFC18]. Standardized [Peb18]. StarBioTrek [CC19]. Started [Wic11]. Stata [Dam09]. State [Båå12, HWW12]. State-space [HWW12]. Stationarity [NTE14]. Stationary [JP18b]. Statistical [BH05, CS16, CDLR09, God13, GP08, LR01, MSB17, THL16, VS15, XC08, Yu02, Scr04, GE14]. Statisticians [PL05]. Statistics [Ago01, HL02, Lan01b, Liu14, Ray01, Ver05, Lia07]. statmod [GS16]. stay [WBS06]. Stellar [VD13, VD13]. stepwise [KLWD18]. Stilt [ORC^{+}18]. Stochastic [BRS18, DHS19, GPB19, MSN^{+}16, HY18, LL05]. Stoichiometric [OGP17]. Story [HUC13]. stplanr [LE18]. Strategies [Gen02]. Stratified [GSHB14, Jon18a, MHL^{+}20]. Streaming [Kap14]. Streams [LL05]. String [Wic10, van14]. stringdist [van14]. stringr [Wic10]. Strings [TDAP18]. strucchange [Zei01]. Structural [AWB17, Zei01]. Structure [ANRG18]. Structures [Bet05]. Student [LDM03, AHvD09, AH10]. Student- [AHvD09, AH10]. Studentized [BF14]. Students [PKH17]. Studies [FO16, Han07b, Jon18a, PLH^{+}10, Sch16, Sto14, Sai05]. Study [AE17, Mor20]. Stylometry [ERK16]. Subgroup [GCG^{+}16]. Subsets [LGE^{+}18]. successful [FVS^{+}20]. SUDAAN [Dam09]. Sul [SP19]. Sum [HW20, LQ18]. Summarized [DSU18]. Sumo [BS12]. Supervised [CJB^{+}13]. Support [FKW^{+}18, Jur12, Mey01, Peb18]. sure [GMBL18]. Surface [OOFdlFS^{+}16, RBS13]. Surrogate [SG16]. SurvBoost [MHL^{+}20]. Survey [JM16, Lum03]. Surveys [AMRdMR15, RFGC20]. Survival [wCTKL20, FKW^{+}18, IK07, KM08, MMS16, Lum04a]. Sweave [Lei02, Bet05, GB06, Gor08, Kuh06, Lei03, Lun11]. swgee [XY19]. switching [HY18]. SWMPr [Bec16]. Symmetry [LRMLLCS17]. System [GGHB18, MU09, Qiu15]. Systems [Bro16, GCG^{+}16, Gra14, GHM16].

Table [KBBH15, CDT04, DDC07]. Tables [Dem16, KBBH13, Oh14, Str19, TKM15].

Tackling [IBG18]. Tails [dCSPM19]. Talk [Hyn11]. Taming [Sie14]. Task [MH07, MC14, Zei05]. Taskforce [Tur20]. Tcl/Tk [Dal01b, Dal02]. tcltk [BCB06]. Teaching [GDM^{+}17, MD18]. Tears [Bra03]. Techniques [Dam09, JM16]. Technologies [MC14]. Tell [HUC13]. Temporal [GW17, GPH16, HTM15, SS13]. tensor [Rou01]. Term [DY16, GW17]. Terry [Han17]. Test [Bos17, Liu14, WB10, YLGT19]. testforDEP [MsHCV18]. Testing [BZST10, BFK^{+}20, CCdUÁ14, Cc20, DB16, GDF18, JP18b, Kap14, NTE14, SP15, TLP07, VZ19, Vil17, Wic11, Zei01, Ran06]. Tests [AG20, AE11, BHH18, CDLR09, DDK18, Fay10, Fey16, FFW13, Hot01, LRMLLCS17, MD18, MsHCV18]. testthat [Wic11]. Text [BVB13, CJB^{+}13, ERK16, Fei08, Jur12]. Texture [NTE14]. Their [MNT18, Rip05b]. Them [HMZ12]. Theory [BC18]. Thermal [Sch16]. Think [PKH17]. Threshold [CLLG08, Rib07]. Tidy [Arn17, MKL20]. Tier [BB16]. tigris [Wal16]. Tiling [RH06]. Time [AWB17, BBACK18, BHM^{+}17, BR20, GP04, Hma20, HWW12, LCT20, MFCR19, MMS16, Mey02, MML16, MBB17, ORC^{+}18, Pfe17, PF16, PMMEGM19, RH01, Rip02, SE19, SS13, Sha11, Sha13, Spe17, WC17, BBC^{+}18, WAD^{+}19]. Time-Series [SE19, HWW12, WAD^{+}19]. Time-Varying [BHM^{+}17]. timeDate [CMW11]. Tips [Coo11]. Tk [Dal01b, Dal02]. tmvtnorm [WM10]. Tolerance [You16, You16]. Too [Bra03]. Tool [MMO20, RSR14, GKZK16]. Toolbox [PMZ20]. Tools [AKU17, Ger16, KMKT20, Len15, LC06, MKL20, MsHCV18, NtP12, OOFdlFS^{+}16, Rob07, RC12, ZG04]. Tour [SC20]. tours [KLC19]. Toxicity [ZC16, Ran06]. Tracks [VD13]. Trade [EKN^{+}08]. trading [ÇT18]. trafo [MKRPC19]. Training [GF10]. Traits [DD18]. Trajectory [Fox09]. Transcript [RH06]. Transformations [JP18a]. Transforming [MKRPC19]. Transitioning [Dam09]. Translating [LDB13]. TRanslational [ACD^{+}16]. Transparency [Mur04]. Transport [LE18]. Treatment [SY19]. Tree [BW15]. Tree-Based [BW15]. treeClust [BW15]. trends [WAD^{+}19]. Trial [Str19]. Trials [KCK^{+}20, OC07, PWD17, YLGT19, PZK06]. TRONCO [ACD^{+}16]. Trophic [CGG13]. Truncated [MMZ16, PBMSN20, WM10]. TSdist [MML16]. tsfknn [MFCR19]. tsmp [BR20]. Tucker [LS20]. TULIP [PMZ20]. Tuning [NP15]. Turk [Lee16]. turned [FVS^{+}20]. Two [BB16, DDC07, Fay10, Fey16, FO16, PMMEGM19, QGC19, WS13]. Two-Factor [FO16]. Two-level [QGC19]. Two-sided [Fay10]. Two-table [DDC07]. Two-Tier [BB16]. Two-way [Fey16]. Type [DY16, Sze18, SHR17]. Types [FT19].

UK [VFB16]. Unbalanced [Vil17].

Uncertain [MK19]. Uncertainties [IBG18]. Uncertainty [ARA^{+}14, SHW18, WFRM15]. Understand [ARA^{+}14]. unequal [Cc20]. Unevenly [PMMEGM19]. Unfolding [BKPW20]. Unidimensionality [FLSNG19]. Unified [KLK20, THL16]. Unifying [KRM18]. Units [PMH16]. unival [FLSNG19]. Univariate [BBACK18, LDC20]. Universe [SNG18]. Unix [Rob07]. Unknown [Gor07]. Unsupervised [ANRG18]. Update [HvWDM17]. Urban [DEVK19]. Use [Car06, EZ20, MUX20, Sch04]. User [BSZWW14, Use13, NW07, Suc20, Sze18, Hyn11, Ric16, Tve15, Ver17]. User-Friendly [Sze18]. User-Specified [Suc20]. Users [BAB^{+}17]. Using [AZJH12, Bet05, BB17b, BHH18, BH05, CMM03, DYB02, Fal06, FLSNG19, Fey16, FFW13, GB06, GPTM15, GdC11, GKZK16, Gor08, GP08, GJP09, Guh06, HW20, IKR10, KF16, Kol19, Lan01b, LDM03, LJ06, Mor18, OOFdlFS^{+}16, OC07, OL09, OSS^{+}17a, PBMSN20, PGCE17, PKH17, Qiu15, Rip01c, SE19, Sch03, SFMR16, Szk16, TZ09, Ver05, War06, Wri16, dLN^{+}18, Zab14, dVdMF19, Ago01, Arn17, CT06, Dug17, FO16, GDF18, GPH16, HP19, Jar07, KG06, MQ06, Mor20, OSS^{+}17b, PZK06, Pen06, SORS06, Sch06, Smi03, WRR18, Lia07]. Utilize [MK19]. utiml [RdC18].

v2.0 [Wri16]. Validating [JM16]. Validation [CG06, GB19, OGP17]. Validity [FLSNG19]. Value [BGK18, LS20, MBB17, PL05, Ste02]. Value-added [BGK18]. Values [Gor07, RPO^{+}19, Dal01a]. VAM [BGK18]. Variable [FT19, GDF18, GPTM15, GB20, IKR10, MH17, MHL^{+}20, Pap18, SVMMRP16, YJBC16]. Variables [HBH15, KB14, LQ18, VCQ15, dVdMF19]. Variance [CKL17, Gro03, Law12]. variances [Cc20]. Variants [BWM14, LB16, MNT18]. Variation [dCSPM19]. Variety [OC07]. Various [LL09]. Varying [BHM^{+}17]. Vdgraph [Law12]. Vector [FKW^{+}18, Mey01, Peb18]. Version [GF19]. Versioning [Oom13]. Versions [YR03]. Very [Han07a]. Verzani [Lia07]. VGAM [Yee08]. via [AE17, DY16, MHL^{+}20, NS17, VZ19, ZZZ^{+}19]. View [MH07, MC14]. Viewing [Mur07]. Views [Zei05]. Vignettes [Lei03]. vip [GB20]. virtual [FVS^{+}20]. visreg [BB17b]. Visual [GB19]. Visualisation [GdC11]. Visualise [MO17]. Visualising [Abe15, Bro16]. Visualization [BB17b, EZ20, Hah17, KW13, MsHCV18, Mor18, TBH17]. Visualize [THL16, VFB16]. Visualizing [DBB^{+}16, Guh06, Sha11, VS15]. vs [VZ19]. VSURF [GPTM15].

Watch [HM11]. water [OOFdlFS^{+}16]. Wavelets [OK07]. Way [DDK18, MKL20, Fey16]. Weather [FRG^{+}11]. Web [BS12, GKZK16, KMKT20, KRJ16, Kol19, MC14, NW07, Ray01, Smi03]. Web-based [Smi03]. Web-tool [GKZK16]. Websites [TKM15]. wec [NtGP17]. WeDiBaDis [IMA16]. Weibull [GSHB14]. Weighted [Ago01, DSN17, HW20, IMA16, NtGP17]. Weighting [SS16]. welchADF [Vil17]. Welfare [BRS18]. Well [HB19]. Well-fitting [HB19]. Whats [OHC19]. while [OL09]. Who [HMZ12]. Wiener [WV14]. Windows [LM05, Rip01b, YR03]. Wine [Yan04]. Without [Bra03, GJP09, KLWD18]. wle [Ago01]. Work [Coo11, Kol19, LM05, Wal16, BC18]. Working [AEB17, CC15, Gor07]. Writing [MW13].

X [WW06, dL01, Sie14]. X-ray [WW06]. X11 [dL01]. xgrid [AZJH12]. XML [Lan01b]. XML2R [Sie14].

Years [KMKT20].

Zero [BKvB^{+}17, LK15]. Zero-inflated [BKvB^{+}17]. Zero/One [LK15]. zoib [LK15].

References

Anderson-Bergman:2017:BRM

[AB17] Clifford Anderson-Bergman. Bayesian regression models for interval-censored data in R. The R Journal, 9(2):487– 498, December 2017. ISSN 2073-4859. URL https:

                                                     //journal.r-project.org/
                                                     archive/2017/RJ-2017-050.

Ardia:2011:DEP

[ABC^{+}11] David Ardia, Kris Boudt, Peter Carl, Katharine M.

Mullen, and Brian G. Peterson. Differential evolution with DEoptim. The R Journal, 3(1):27–34, June 2011. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2011-1/RJournal_
2011-1_Ardia~et~al.pdf.

Ardia:2018:DRE

[ABC18] David Ardia, Kris Boudt, and Leopoldo Catania. Downside risk evaluation with the R package GAS. The R Journal, 10(2):410–421, December 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-064.

Abel:2015:PRP

[Abe15] Guy Abel. fanplot: An R package for visualising sequential distributions. The R Journal, 7(1):15–23, June 2015. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2015-1/RJournal_
2015-1_abel.pdf.

Allignol:2008:PRP

[ABS08] A. Allignol, J. Beyersmann, and M. Schumacher. mvna: An R package for the Nelson– Aalen estimator in multistate models. R News: the Newsletter of the R Project, 8(2): 48–50, October 2008. CO-DEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Antoniotti:2016:DTB

[ACD^{+}16] Marco Antoniotti, Giulio Caravagna, Luca De Sano, Alex Graudenzi, Giancarlo Mauri, Bud Mishra, and Daniele Ramazzotti. Design of the TRONCO BioConductor package for TRanslational ON-COlogy. The R Journal, 8 (2):39–59, December 2016. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2016/RJ-2016-032.

Adammer:2019:LRP

[Adä19] Philipp Adämmer. lpirfs: an R package to estimate impulse response functions by local projections. The R Journal, 11(2):421–438, December 2019. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2019/RJ-2019-052.

Adler:2012:FLI

[Adl12] Daniel Adler. Foreign library interface. The R Journal, 4(1):30–40, June 2012. CODEN ???? ISSN 2073-4859. URL http:

                                        //journal.r-project.org/
                                        archive/2012-1/RJournal_
                                        2012-1_Adler.pdf.

Azarang:2018:IRM

[AdlF18] Leyla Azarang and Manuel Oviedo de la Fuente. idmTPreg: Regression model for progressive illness death data. The R Journal, 10(2):317– 325, December 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-081.

Acion:2018:CRL

[AdSQ18] Laura Acion, Natalia da Silva, and Riva Quiroga. Conference report: LatinR 2018. The R Journal, 10(2):564– 566, December 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018-2/latinR.pdf.

Arnold:2011:NGF

[AE11] Taylor B. Arnold and John W.

Emerson. Nonparametric goodness-of-fit tests for discrete null distributions. The R Journal, 3(2):34–39, December 2011. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2011-2/RJournal_
2011-2_Arnold+Emerson.pdf.

Anderson:2017:HDP

[AE17] G. Brooke Anderson and Dirk Eddelbuettel. Hosting data packages via drat: a case study with hurricane exposure data. The R Journal, 9(1):486–497, June 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-026.

Anderson:2017:WDC

[AEB17] G. Brooke Anderson, Colin Eason, and Elizabeth A. Barnes. Working with daily climate model output data in

R and the futureheatwaves

package. The R Journal, 9 (1):124–137, June 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-032.

Altunkaynak:2020:NRP

[AG20] Bulent Altunkaynak and Hamza Gamgam. npordtests: an R package of nonparametric tests for equality of location against ordered alternatives. The R Journal, 12(1):147–171, June 2020. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2020/RJ-2020-022.

Agostinelli:2001:PPR

[Ago01] Claudio Agostinelli. wle: A package for robust statistics using weighted likelihood. R News: the Newsletter of the R Project, 1(3):32–38, September 2001. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/
Rnews/.

Ardia:2010:BEG

[AH10] David Ardia and Lennart F. Hoogerheide. Bayesian estimation of the GARCH(1,1) model with Student-t innovations. The R Journal, 2(2):41–47, December 2010. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2010-2/RJournal_
2010-2_Ardia+Hoogerheide.
pdf.

Alvares:2019:SRP

[AHLL19] Danilo Alvares, Sebastien Haneuse, Catherine Lee, and

Kyu Ha Lee. SemiCompRisks:

an R package for the analysis of independent and clustercorrelated semi-competing risks data. The R Journal, 11(1):376–400, June 2019. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2019/RJ-2019-038.

Ardia:2009:PAM

[AHvD09] David Ardia, Lennart F.

Hoogerheide, and Herman K. van Dijk. AdMit: Adaptive mixtures of Studentt distributions. The R Journal, 1(1):25–30, June 2009. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2009-1/RJournal_
2009-1_Ardia+et+al.pdf.

Arnold:2017:IHP

[AKU17] Taylor Arnold, Michael J.

Kane, and Simon Urbanek. iotools: High-performance I/O tools for R. The R Journal, 9(1):6–13, June 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-001.

An:2016:KRP

[AL16] Weihua An and Yu-Hsin Liu.

keyplayer: an R package for locating key players in social networks. The R Journal, 8(1):257–268, August 2016.

ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2016/RJ-2016-018.

An:2019:DSC

[ALB^{+}19] Hyowon An, Justin T. Landis, Aubrey G. Bailey, James S. Marron, and Dirk P. Dittmer. dr4pl: a stable convergence algorithm for the 4 parameter logistic model. The R Journal, 11(2):171– 190, December 2019. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2019/RJ-2019-003.

Almeida:2018:GCQ

[ALH18] Alexandre Almeida, Adam Loy, and Heike Hofmann. ggplot2 compatible quantilequantile plots in R. The R Journal, 10(2):248–261, December 2018. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2018/RJ-2018-051.

Alpert:2007:FJR

[Alp07] Bill Alpert. Financial journalism with R. R News: the Newsletter of the R Project, 7(3):34–36, December 2007. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-
                                        project.org/doc/Rnews/.

Arcos:2015:PPE

[AMRdMR15] Antonio Arcos, David Molina, Maria Giovanna Ranalli, and María del Mar Rueda. Frames2: A package for estimation in dual frame surveys. The R Journal, 7(1):52– 72, June 2015. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2015-1/RJournal_
2015-1_arcos-molina-ranalli-
etal.pdf.

Archimbaud:2018:IUO

[ANRG18] Aurore Archimbaud, Klaus Nordhausen, and Anne Ruiz-Gazen. ICSOutlier: Unsupervised outlier detection for low-dimensional contamination structure. The R Journal, 10(1):234–250, July 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-034.

Alden:2014:APU

[ARA^{+}14] Kieran Alden, Mark Read, Paul S. Andrews, Jon Timmis, and Mark Coles. Applying spartan to understand parameter uncertainty in simulations. The R Journal, 6(2):63–80, December 2014. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2014-2/RJournal_
2014-2_alden-read-andrews-
etal.pdf.

Armstrong:2013:PIP

[Arm13] David A. Armstrong II.

factorplot: Improving presentation of simple contrasts in GLMs. The R Journal, 5(2):4–16, December 2013. CODEN ???? ISSN 2073-4859. URL http:

                                        //journal.r-project.org/
                                        archive/2013-2/RJournal_
                                        2013-2_armstrong.pdf.

Arnold:2017:TDM

[Arn17] Taylor Arnold. A tidy data model for natural language processing using cleanNLP. The R Journal, 9(2):248– 267, December 2017. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2017/RJ-2017-035.

Alam:2015:FCS

[ARS15] Moudud Alam, Lars Rönnegård, and Xia Shen. Fitting conditional and simultaneous autoregressive spatial models in hglm. The R Journal, 7(2): 5–18, December 2015. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2015/RJ-2015-017.

Allison:2017:RRP

[AWB17] Annabel Allison, Ian R. White, and Simon Bond. rpsftm: an R package for rank preserving structural failure time models. The R Journal, 9(2):342– 353, December 2017. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2017/RJ-2017-068.

Anoke:2012:XRP

[AZJH12] Sarah C. Anoke, Yuting Zhao, Rafael Jaeger, and Nicholas J. Horton. xgrid and R: Parallel distributed processing using heterogeneous groups of Apple computers. The R Journal, 4(1):45–55, June 2012. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2012-1/RJournal_
2012-1_Anoke~et~al.pdf.

Baaaath:2012:SNC

[Båå12] Rasmus Bååth. The state of naming conventions in R. The R Journal, 4(2):74–75, December 2012. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2012-2/RJournal_
2012-2_Baaaath.pdf.

Beresewicz:2017:CRE

[BAB^{+}17] Maciej Beresewicz, Adolfo Alvarez, Przemysław Biecek, Marcin K. Dyderski, Marcin Kosinski, Jakub Nowosad, Kamil Rotter, Alicja Szabelska-Beresewicz, Marcin Szymkowiak, Łukasz Wawrowski, and Joanna Zyprych-Walczak. Conference report: European R Users Meeting 2016. The R Journal, 9(1):501– 504, June 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017-1/erum.pdf.

Bokde:2017:PIR

[BACMÁK17] Neeraj Bokde, Gualberto Asencio-Cortés, Francisco Martínez-Álvarez, and Kishore Kulat. PSF: Introduction to R package for pattern sequence based forecasting algorithm. The R Journal, 9 (1):324–333, June 2017. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2017/RJ-2017-021.

Ball:2003:L

[Bal03] Rod Ball. lmeSplines. R News: the Newsletter of the R Project, 3(3):24–28, December 2003. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/
                                        Rnews/.

Barry:2018:CRR

[Bar18]  Timothy Barry.  Collec-

tions in R: Review and proposal. The R Journal, 10(1): 455–471, July 2018. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2018/RJ-2018-037.

Bates:2003:CPS

[Bat03] Douglas Bates. Converting packages to S4. R News: the Newsletter of the R Project, 3(1):6–8, June 2003. CO-DEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-
                                        project.org/doc/Rnews/.

Bates:2004:LSC

[Bat04] Douglas Bates. Least squares calculations in R. R News: the Newsletter of the R Project, 4(1):17–20, June 2004. CO-DEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-
                                        project.org/doc/Rnews/.

Bates:2005:FLM

[Bat05] Douglas Bates. Fitting linear mixed models in R. R News: the Newsletter of the R Project, 5(1):27–30, May 2005. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/
Rnews/.

Bacci:2016:TTL

[BB16] Silvia Bacci and Francesco Bartolucci. Two-tier latent class IRT models in R. The R Journal, 8(2):139– 166, December 2016. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2016/RJ-2016-038.

Barrett:2017:FQS

[BB17a] Tyson S. Barrett and Emily Brignone. Furniture for quantitative scientists. The R Journal, 9(2):142–148, December 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-037.

Breheny:2017:VRM

[BB17b] Patrick Breheny and Woodrow Burchett. Visualization of regression models using visreg. The R Journal, 9(2):56– 71, December 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-046.

Bucklin:2018:RLR

[BB18] David Bucklin and Mathieu Basille. rpostgis: Linking R with a PostGIS spatial database. The R Journal, 10(1):251–268, July 2018.

ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-025.

Beck:2018:RPI

[BBACK18] Marcus W. Beck, Neeraj Bokde, Gualberto Asencio-Cortés, and Kishore Kulat.

R package imputeTestbench

to compare imputation methods for univariate time series. The R Journal, 10(1): 218–233, July 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-024.

Barbu:2018:SRP

[BBC^{+}18] Vlad Stefan Barbu, Caroline Bérard, Dominique Cellier, Mathilde Sautreuil, and Nicolas Vergne. SMM: an R package for estimation and simulation of discrete-time semi-Markov models. The R Journal, 10 (2):226–247, December 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-050.

Barnett:2012:ASD

[BBD12] Adrian G. Barnett, Peter Baker, and Annette J. Dobson. Analysing seasonal data. The R Journal, 4(1):5–10, June 2012. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2012-1/RJournal_
2012-1_G~Barnett-et~al.
pdf.

Benitez:2017:ARP

[BBTH17] Rafael Benítez, Vicente J.

Bolós, and José-Luis Toca-Herrera. afmToolkit: an R package for automated AFM force-distance curves analysis. The R Journal, 9(2):291– 308, December 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-045.

er:2018:EGT

[BC18] Bilge Basȩr and Nalan Cinemre. Epistemic game theory: Putting algorithms to work. The R Journal, 10(1): 370–380, July 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-003.

Bowman:2006:PMG

[BCB06] Adrian Bowman, Ewan Crawford, and Richard Bowman. rpanel: making graphs move with tcltk. R News: the Newsletter of the R Project, 6 (4):12–17, October 2006. CO-DEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Bollmann:2017:FC

[BCD^{+}17] Stella Bollmann, Dianne Cook, Jasmine Dumas, John Fox, Julie Josse, Oliver Keyes, Carolin Strobl, Heather Turner, and Rudolf Debelak. Forwards column. The R Journal, 9(2):541–552, December 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/

                                        archive/2017-2/forwards.
                                        pdf.

Baradel:2017:CRR

[BDH17] Nicolas Baradel, Christophe Dutang, and Caroline Hillairet. Conference report: R in Insurance 2017. The R Journal, 9 (2):539–540, December 2017. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2017-2/insurance.
                                        pdf.

Boettiger:2017:IRD

[BE17] Carl Boettiger and Dirk Eddelbuettel. An introduction to Rocker: Docker containers for R. The R Journal, 9 (2):527–536, December 2017. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2017/RJ-2017-065.

Beath:2016:MRP

[Bea16] Ken J. Beath. metaplus: an R package for the analysis of robust meta-analysis and meta-regression. The R Journal, 8(1):5–16, August 2016. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2016/RJ-2016-001.

Beck:2016:SRP

[Bec16] Marcus W. Beck. SWMPr: an R package for retrieving, organizing, and analyzing environmental data for estuaries. The R Journal, 8(1):219–232, August 2016. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2016/RJ-2016-015.

Basulto-Elias:2017:FRP

[BECDN17] Guillermo Basulto-Elias, Alicia Carriquiry, Kris De Brabanter, and Daniel J. Nordman. fourierin: an R package to compute Fourier integrals. The R Journal, 9 (2):72–83, December 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-044.

Benner:2005:PMF

[Ben05] Axel Benner. mfp: Multivariable fractional polynomials. R News: the Newsletter of the R Project, 5(2): 20–23, November 2005. CO-DEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Betebenner:2005:UCS

[Bet05]  Damian Betebenner.  Us-

ing control structures with Sweave. R News: the Newsletter of the R Project, 5(1): 40–44, May 2005. CO-DEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Batista:2014:PRP

[BF14] Ben Dêivide Oliveira Batista and Daniel Furtado Ferreira. SMR: An R package for computing the externally studentized normal midrange distribution. The R Journal, 6(2):123–136, December 2014. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/

                                        archive/2014-2/RJournal_
                                        2014-2_ferreira-batista.
                                        pdf.

Bohn:2011:CBS

[BFHM11] Angela Bohn, Ingo Feinerer, Kurt Hornik, and Patrick Mair. Content-based social network analysis of mailing lists. The R Journal, 3(1):11– 18, June 2011. CODEN ???? ISSN 2073-4859. URL http:

                                        //journal.r-project.org/
                                        archive/2011-1/RJournal_
                                        2011-1_Bohn~et~al.pdf.

Bose:2020:FRP

[BFK^{+}20] Koushiki Bose, Jianqing Fan, Yuan Ke, Xiaoou Pan, and Wen-Xin Zhou. FarmTest: an R package for factoradjusted robust multiple testing. The R Journal, 12 (2):372–387, December 2020. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2021/RJ-2021-023.

Bartoszuk:2020:SRC

[BG20] Maciej Bartoszuk and Marek Gagolewski. SimilaR: R code clone and plagiarism detection. The R Journal, 12(1): 367–385, June 2020. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2020/RJ-2020-017.

Bai:2020:MSS

[BGBV20] Lu Bai, Daniel L. Gillen, Scott M. Bartell, and Verónica M. Vieira. Mapping smoothed spatial effect estimates from individual-level data: MapGAM. The R Journal, 12(1):32– 48, June 2020. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2020/RJ-2020-001.

Broatch:2018:RRP

[BGK18] Jennifer Broatch, Jennifer Green, and Andrew Karl. RealVAMS: an R package for fitting a multivariate value-added model (VAM). The R Journal, 10(1):22– 30, July 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-033.

Brownrigg:2005:URS

[BH05] Ray Brownrigg and David Harte. Using R for statistical seismology. R News: the Newsletter of the R Project, 5(1):31–35, May 2005. CO-DEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Brill:2018:NIT

[BHH18] Barak Brill, Yair Heller, and Ruth Heller. Nonparametric independence tests and k-sample tests for large sample sizes using package HHG. The R Journal, 10(1): 424–438, July 2018. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2018/RJ-2018-008.

Beger:2017:SIS

[BHM^{+}17] Andreas Beger, Daniel W.

Hill, Jr., Nils. W. Metternich, Shahryar Minhas, and Michael D. Ward. Splitting it up: The spduration splitpopulation duration regression package for time-varying covariates. The R Journal, 9 (2):474–486, December 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017/RJ-2017-056.

Bretz:2002:MCR

[BHW02] Frank Bretz, Torsten Hothorn, and Peter Westfall. On multiple comparisons in R. R News: the Newsletter of the R Project, 2(3):14–17, December 2002. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/
Rnews/.

Brouste:2016:CFG

[BILL16] Alexandre Brouste, Jacques Istas, and Sophie Lambert-Lacroix. Conditional fractional Gaussian fields with the package FieldSim. The R Journal, 8(1):38–47, August 2016. ISSN 2073-4859.

URL https://journal.r-
project.org/archive/2016/
RJ-2016-003.

Team:2013:NBPa

[Bio13a] Bioconductor Team. News from the Bioconductor Project. The R Journal, 5(1):218–220, June 2013. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2013-1/RJournal_
2013-1_bioconductor.pdf.

Team:2013:NBPb

[Bio13b] Bioconductor Team. News from the Bioconductor Project. The R Journal, 5(2):162–164, December 2013. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2013-2/RJournal_
2013-2_bioconductor.pdf.

Team:2014:NBPa

[Bio14] Bioconductor Team. News from the Bioconductor Project. The R Journal, 6(1):184–185, June 2014. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2014-1/RJournal_
2014-1_bioconductor.pdf.

Bivand:2001:MSD

[Biv01] Roger Bivand. More on spatial data. R News: the Newsletter of the R Project, 1(3):13–17, September 2001. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Bivand:2017:Ea

[Biv17a]  Roger Bivand.  Edito-

rial. The R Journal, 9 (1):4–5, June 2017. ISSN 2073-4859. URL https:

//journal.r-project.org/
archive/2017-1/editorial.
pdf.

Bivand:2017:Eb

[Biv17b]  Roger Bivand.  Editorial.

The R Journal, 9(2):4– 5, December 2017. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2017-2/editorial.
                                        pdf.

Baumann:2019:FPA

[BK19] Stuart Baumann and Margaryta Klymak. Fixed point acceleration in R. The R Journal, 11(1):359–375, June 2019. ISSN 2073-4859. URL https:

                                        //journal.r-project.org/
                                        archive/2019/RJ-2019-037.

Burdukiewicz:2018:CRW

[BKJ^{+}18] Michał Burdukiewicz, Marta Karas, Leon Eyrich Jessen, Marcin Kosinśki, BerndBischl, andStefanRödiger.Conferencereport : W hyR?2018.TheRJournal, 10

(2) : 572−−578, December2018.ISSN2073−

                                        4859.URLhttpsjournalrprojectorgarchive20182whyRpdf.

Balafas:2020:MRP

[BKPW20] Spyros E. Balafas, Wim P. Krijnen, Wendy J. Post, and Ernst C. Wit. mudfold: an R package for nonparametric IRT modelling of unfolding processes. The R Journal, 12(1):49–75, June 2020. ISSN 2073-

                           4859. URL https://journal.r-project.
                           org/archive/2020/RJ-2020-002.

Brooks:2017:GBS

[BKvB^{+}17] Mollie E. Brooks, Kasper Kristensen, Koen J. van Benthem, Arni Magnusson, Casper W. Berg, Anders Nielsen, Hans J. Skaug, Martin Mächler, and Benjamin M. Bolker. glmmTMB balances speed and flexibility among packages for zero-inflated generalized linear mixed modeling. The R Journal, 9 (2):378–400, December 2017. ISSN 2073-

                           4859. URL https://journal.r-project.
                           org/archive/2017/RJ-2017-066.

Baier:2011:CDA

[BND11] Thomas Baier, Erich Neuwirth, and Michele De Meo. Creating and deploying an application with (R)Excel and R. The R Journal, 3(2):5–11, December 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-2/RJournal_
2011-2_Baier~et~al.pdf.

Boehringer:2013:DPR

[Boe13] Stefan Boehringer. Dynamic parallelization of R functions. The R Journal, 5 (2):88–97, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_boehringer.pdf.

Bossek:2017:SSM

[Bos17] Jakob Bossek. smoof: Single- and multiobjective optimization test functions. The R Journal, 9(1):103–113, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-004.

Burdukiewicz:2020:CRW

[BPC^{+}20] Michał Burdukiewicz, Filip Pietluch, Jarosław Chilimoniuk, Katarzyna Sidorczuk, Dominik Rafacz, Leon Eyrich Jessen, Stefan Rödiger, Marcin Kosin śki, andP iotrW ójcik.Conferencereport : W hyR?2019.TheRJournal, 12(1) : 484, June2020.ISSN2073−

4859.URLhttpsjournalrprojectorgarchive20201whyRpdf.

Bischoff:2020:TRP

[BR20] Francisco Bischoff and Pedro Pereira Rodrigues. tsmp: an R package for time series with matrix profile. The R Journal, 12(1):76–86, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-021.

Brahm:2002:DDP

[Bra02] David E. Brahm. Delayed data packages. R News: the Newsletter of the R Project, 2(3):

11–12, December 2002. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Bravington:2003:DTM

[Bra03] Mark Bravington. Debugging without (too many) tears. R News: the Newsletter of the R Project, 3(3):29–32, December 2003. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Brazzale:2005:PRP

[Bra05] Alessandra Brazzale. hoa: An R package bundle for higher order likelihood inference. R News: the Newsletter of the R Project, 5 (1):20–27, May 2005. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Brown:2016:MCR

[Bro16] Patrick E. Brown. Maps, coordinate reference systems and visualising geographic data with mapmisc. The R Journal, 8(1):64–91, August 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-005.

Berihuete:2018:WIP

[BRS18] Angel Berihuete, Carmen D. Ramos, and Miguel A. Sordo. Welfare, inequality and poverty analysis with rtip: an approach based on stochastic dominance. The R Journal, 10(1):328–341, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-029.

Bergsma:2012:PAW

[BS12] Timothy T. Bergsma and Michael S. Smith. Sumo: An authenticating Web application with an embedded R session. The R Journal, 4(1):60–63, June 2012. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2012-1/RJournal_
2012-1_Bergsma+Smith.pdf.

Bilgic:2013:PRP

[BS13] Yusuf K. Bilgic and Herb Susmann. rlme:

An R package for rank-based estimation and prediction in random effects nested models. The R Journal, 5(2):71–80, December 2013. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2013-2/
RJournal_2013-2_bilgic-susmann.pdf.

Beresewicz:2014:CRP

[BSZWW14] Maciej Beresewicz, Alicja Szabelska, Joanna Zyprych-Walczak, and Łukasz Wawrowski. Conference report: Polish Academic R User Meeting. The R Journal, 6(2): 187–189, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_beresewicz-szabelska-zyprychwalczak-
etal.pdf.

Baumgartner:2015:ICC

[BT15] Michael Baumgartner and Alrik Thiem. Identifying complex causal dependencies in configurational data with coincidence analysis. The R Journal, 7(1):176–184, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_baumgartner-thiem.pdf.

Bensicacute:2017:LLA

[BTH^{+}17] Mirta Benšic´,P etarT aler, SafetHamedovic´,Emmanuel KarloNyarko, andKristianSabo.LeArEst : Lengthandareaestimationfromdatameasuredwithadditiveerror.TheRJournal, 9

(2) : 461 − −473, December2017.ISSN2073 −

4859.URLhttpsjournalrprojectorgarchive2017RJ2017043.

Burkner:2018:ABM

[Bür18] Paul-Christian Bürkner. Advanced Bayesian multilevel modeling with the R package brms. The R Journal, 10(1):395–411, July 2018.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2018/RJ-2018-017.

Bouchet-Valat:2013:RTG

[BVB13] Milan Bouchet-Valat and Gilles Bastin. RcmdrPlugin.temis, a graphical integrated text mining solution in R. The R Journal, 5(1):188–197, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_bouchetvalat-bastin.pdf.

Buttrey:2015:TRP

[BW15] Samuel E. Buttrey and Lyn R. Whitaker. treeClust: an R package for tree-based clustering dissimilarities. The R Journal, 7 (2):227–236, December 2015. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2015/RJ-2015-032.

Bottomly:2014:PFA

[BWM14] Daniel Bottomly, Beth Wilmot, and Shannon K. McWeeney. oligoMask: A framework for assessing and removing the effect of genetic variants on microarray probes. The R Journal, 6(1):159–163, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2014-1/RJournal_
                                        2014-1_bottomly-wilmot-mcweeney.pdf.

Brown:2010:MGL

[BZ10] Patrick Brown and Lutong Zhou. MCMC for generalized linear mixed models with glmmBUGS. The R Journal, 2(1):13–17, June 2010. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2010-1/RJournal_2010-1_Brown+Zhou.pdf.

Bilder:2010:PPG

[BZST10] Christopher R. Bilder, Boan Zhang, Frank Schaarschmidt, and Joshua M. Tebbs. binGroup: A package for group testing. The R Journal, 2(2):56–60, December 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2010-2/RJournal_
2010-2_Bilder~et~al.pdf.

Cannas:2019:MCD

[CA19] Massimo Cannas and Bruno Arpino. Matching with clustered data: the CMatching package in R. The R Journal, 11(1):7–21, June 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-018.

Canty:2002:RMR

[Can02] Angelo J. Canty. Resampling methods in R: The boot package. R News: the Newsletter of the R Project, 2(3):2–7, December 2002. CO-DEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Carey:2006:SMA

[Car06] Vince Carey. SNP metadata access and use with Bioconductor. R News: the Newsletter of the R Project, 6(5):36–39, December 2006. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Carey:2008:PNR

[Car08] Vince Carey. Programmers’ niche: The Y of R. R News: the Newsletter of the R Project, 8 (2):51–53, October 2008. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Charte:2015:WMD

[CC15] Francisco Charte and David Charte. Working with multilabel datasets in R: The mldr package. The R Journal, 7(2): 149–162, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-027.

Cava:2019:INP

[CC19] Claudia Cava and Isabella Castiglioni. Integration of networks and pathways with StarBioTrek package. The R Journal, 11(1):310–322, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-025.

Cavus:2020:TEN

[Cc20] Mustafa Cavus and Berna Yazıcı. Testing the equality of normal distributed and independent groups’ means under unequal variances by doex package. The R Journal, 12 (2):134–154, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-008.

Castro-Conde:2014:PRP

[CCdUÁ14] Irene Castro-Conde and Jacobo de Uña-Álvarez. sgof: An R package for multiple testing problems. The R Journal, 6 (2):96–113, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_conde-alvarez.pdf.

Chen:2020:SSC

[CCFVL20] Shuowen Chen, Victor Chernozhukov, Iván Fernández-Val, and Ye Luo. SortedEffects: Sorted causal effects in R. The R Journal, 12(1):131–146, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-006.

Chen:2017:CRP

[CCFVM17] Mingli Chen, Victor Chernozhukov, Iván Fernández-Val, and Blaise Melly. Counterfactual: an R package for counterfactual analysis. The R Journal, 9(1):370–384, June 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-033.

Cattaneo:2015:PRP

[CCT15] Matias D. Cattaneo, Sebastian Calonico, and Rocio Titiunik. rdrobust: An R package for robust inference in regressiondiscontinuity designs. The R Journal, 7(1):38–51, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_cattaneo-calonico-titiunik.pdf.

Chen:2017:MMI

[CCY^{+}17] Ping-Yang Chen, Ching-Chuan Chen, Chun-Hao Yang, Sheng-Mao Chang, and Kuo-Jung Lee. milr: Multiple-instance logistic regression with lasso penalty. The R Journal, 9(1):446–457, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-013.

Coeurjolly:2009:PSR

[CDLR09] J. F. Coeurjolly, R. Drouilhet, P. Lafaye de Micheaux, and J. F. Robineau. asympTest: A simple R package for classical parametric statistical tests and confidence intervals in large samples. The R Journal, 1(2):26–30, December 2009. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2009-2/RJournal_2009-2_Coeurjolly~et~
al.pdf.

Chessel:2004:PPO

[CDT04] Daniel Chessel, Anne B. Dufour, and Jean Thioulouse. The ade4 package — I: Onetable methods. R News: the Newsletter of the R Project, 4(1):5–10, June 2004. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Ching:2018:RMH

[CE18] Travers Ching and Dirk Eddelbuettel. RcppMsgPack: MessagePack headers and interface functions for R. The R Journal, 10 (2):516–525, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-068.

Campbell:2007:B

[CEGK07] Kyle Campbell, Jeff Enos, Daniel Gerlanc, and David Kane. Backtests. R News: the Newsletter of the R Project, 7(1):36–41, April 2007. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Cattaneo:2020:LPB

[CFF20] Matias D. Cattaneo, Max H. Farrell, and Yingjie Feng. lspartition: Partitioningbased least squares regression. The R Journal, 12(1):172–187, June 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020/RJ-2020-005.

Cook:2006:BSV

[CG06] Samantha Cook and Andrew Gelman. Bayesian software validation. R News: the Newsletter of the R Project, 6(1):11–12, March 2006. CODEN ???? ISSN 1609-3631. URL

                                        http://CRAN.R-project.org/doc/Rnews/.

Colleter:2013:IER

[CGG13] Mathieu Colleter, Jerome Guitton, and Didier Gascuel. An introduction to the EcoTroph R package: Analyzing aquatic ecosystem trophic network. The R Journal, 5(1):98–108, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_colleter-guitton-gascuel.pdf.

Chen:2018:SBM

[CGLN18] Yuzhou Chen, Yulia R. Gel, Vyacheslav Lyubchich, and Kusha Nezafati. Snowboot: Bootstrap methods for network inference. The R Journal, 10(2):95–113, December 2018.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2018/RJ-2018-056.

Chambers:2009:FR

[Cha09] John M. Chambers. Facets of R. The R Journal, 1(1):5–8, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2009-1/RJournal_
2009-1_Chambers.pdf.

Chambers:2020:RDS

[Cha20] John M. Chambers. S, R, and data science. The R Journal, 12(1):462–476, June 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2020/RJ-2020-028.

Christensen:2018:NMM

[Chr18] Alexander P. Christensen. NetworkToolbox: Methods and measures for brain, cognitive, and psychometric network analysis in R. The R Journal, 10(2):422–439, December 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-065.

Chernozhukov:2016:HHD

[CHS16] Victor Chernozhukov, Chris Hansen, and Martin Spindler. hdm: High-dimensional metrics. The R Journal, 8(2):185–199, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-040.

Canida:2017:ORI

[CI17] Travis Canida and John Ihrie. openEBGM: an R implementation of the gamma-Poisson shrinker data mining model. The R Journal, 9(2):499–519, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-063.

Collingwood:2013:PSL

[CJB^{+}13] Loren Collingwood, Timothy P. Jurka, Amber E. Boydstun, Emiliano Grossman, and Wouter van Atteveldt. RTextTools: A supervised learning package for text classification. The R Journal, 5(1):6–13, June 2013. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/archive/2013-
1/RJournal_2013-1_collingwood-jurka-
boydstun-etal.pdf.

Choe:2017:ERP

[CKL17] Hye-Min Choe, Mijeong Kim, and Eun-Kyung Lee. EMSaov: an R package for the analysis of variance with the expected mean squares and its shiny application. The R Journal, 9(1):252–261, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-011.

Chambers:2001:OOP

[CL01a] John M. Chambers and Duncan Temple Lang. Object-oriented programming in R. R News: the Newsletter of the R Project, 1(3):17–19, September 2001. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Chambers:2001:OPR

[CL01b] John M. Chambers and Duncan Temple Lang. Omegahat packages for R. R News: the Newsletter of the R Project, 1(1):21–24, January 2001. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Chen:2008:BEP

[CLLG08] Cathy W. S. Chen, Edward M. H. Lin, F. C. Liu, and Richard Gerlach. Bayesian estimation for parsimonious threshold autoregressive models in R. R News: the Newsletter of the R Project, 8(1):26–33, May 2008. CODEN ???? ISSN 1609-3631. URL http://CRAN.

                                        R-project.org/doc/Rnews/.

Clifford:2006:RF

[CM06] David Clifford and Peter McCullagh. The regress function. R News: the Newsletter of the R Project, 6(2):6–10, May 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Carson:2003:CGU

[CMM03] Brett Carson, Robert Murison, and Ian A. Mason. Computational gains using RPVM on a Beowulf cluster. R News: the Newsletter of the R Project, 3(1):21–26, June 2003. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Chalabi:2011:PPP

[CMW11] Yohan Chalabi, Martin Mächler, and Diethelm Würtz. Rmetrics — timeDate package. The R Journal, 3(1):19–24, June 2011. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2011-1/
RJournal_2011-1_Chalabi~et~al.pdf.

Collingwood:2016:ECE

[COGRB16] Loren Collingwood, Kassra Oskooii, Sergio Garcia-Rios, and Matt Barreto. eiCompare: Comparing ecological inference estimates across EI and EI:RC. The R Journal, 8(2):92– 101, December 2016. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2016/RJ-2016-035.

Cook:2011:TPY

[Coo11] Dianne Cook. Tips for presenting your work. The R Journal, 3(1):72–74, June 2011. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2011-1/
RJournal_2011-1_Cook.pdf.

Charlier:2015:QRP

[CPS15] Isabelle Charlier, Davy Paindaveine, and Jérôme Saracco. QuantifQuantile: an R package for performing quantile regression through optimal quantization. The R Journal, 7(2):65–80, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-021.

Christensen:2002:PPG

[CR02] Ole F. Christensen and Paulo J. Ribeiro. geoRglm: A package for generalised linear spatial models. R News: the Newsletter of the R Project, 2(2):26–28, June 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.

                                        R-project.org/doc/Rnews/.

Carpenter:2009:PRP

[CRS09] J. Carpenter, G. Rücker, and G. Schwarzer. copas: An R package for fitting the Copas selection model. The R Journal, 1 (2):31–36, December 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-2/RJournal_
                                        2009-2_Carpenter~et~al.pdf.

Calvo:2016:SSC

[CS16] Borja Calvo and Guzmán Santafé. scmamp: Statistical comparison of multiple algorithms in multiple problems. The R Journal, 8 (1):248–256, August 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-017.

Culhane:2006:MAI

[CT06] Aedín C. Culhane and Jean Thioulouse. A multivariate approach to integrating datasets using made4 and ade4. R News: the Newsletter of the R Project, 6(5):54–58, December 2006. CODEN ???? ISSN 1609-3631. URL

                                        http://CRAN.R-project.org/doc/Rnews/.

Celik:2018:IRP

[ÇT18] Duygu Çelik and Murat Tiniç. InfoTrad: an R package for estimating the probability of informed trading. The R Journal, 10(1):31–42, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-013.

Cordoba:2019:CRC

[CvDMvE19] Marcela Alfaro Córdoba, Frans van Dunné, Agustín Gómez Meléndez, and Jacob van Etten. Conference: Report ConectaR 2019. The R Journal, 11(2):439–442, December 2019. ISSN 2073-4859. URL https://journal.

r-project.org/archive/2019-2/conectaR.
pdf.

Dalgaard:2001:DNC

[Dal01a] Peter Dalgaard. The density of the noncentral chi-squared distribution for large values of the noncentrality parameter. R News: the Newsletter of the R Project, 1(1):14–15, January 2001. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Dalgaard:2001:PRT

[Dal01b] Peter Dalgaard. A primer on the R-Tcl/Tk package. R News: the Newsletter of the R Project, 1(3):27–31, September 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Dalgaard:2002:CRT

[Dal02] Peter Dalgaard. Changes to the R-Tcl/Tk package. R News: the Newsletter of the R Project, 2(3):25–27, December 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Dalgaard:2007:NFM

[Dal07] Peter Dalgaard. New functions for multivariate analysis. R News: the Newsletter of the R Project, 7(2):2–7, October 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Damico:2009:TRR

[Dam09] Anthony Damico. Transitioning to R: Replicating SAS, Stata, and SUDAAN analysis techniques in health policy data. The R Journal, 1(2):37–44, December 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2009-2/RJournal_
2009-2_Damico.pdf.

Daroczi:2018:CRE

[Dar18] Gergely Daróczi. Conference report: eRum 2018. The R Journal, 10(1):549–550, July 2018. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2018-1/
erums.pdf.

Demirhan:2016:CRP

[DB16] Haydar Demirhan and Nihan Bitirim. CryptRndTest: an R package for testing the cryptographic randomness. The R Journal, 8(1):233–247, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-016.

Deveau:2016:CBM

[DBB^{+}16] Paul Deveau, Emmanuel Barillot, Valentina Boeva, Andrei Zinovyev, and Eric Bonnet. Calculating biological module enrichment or depletion and visualizing data on large-scale molecular maps with ACSNMineR and RNaviCell packages. The R Journal, 8(2):293–306, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-047.

delCastillo:2019:FTE

[dCSPM19] Joan del Castillo, Isabel Serra, Maria Padilla, and David Moriña. Fitting tails by the empirical residual coefficient of variation: The ercv package. The R Journal, 11 (2):56–68, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-044.

Silva:2018:PRP

[DD18] Stéphane Blondeau Da Silva and Anne Da Silva. Pstat: an R package to assess population differentiation in phenotypic traits. The R Journal, 10(1):447–454, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-010.

Dray:2007:PPI

[DDC07] Stéphane Dray, Anne B. Dufour, and Daniel Chessel. The ade4 package — II: Two-table and K-table methods. R News: the Newsletter of the R Project, 7(2):47–52, October 2007. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Dag:2018:ORP

[DDK18] Osman Dag, Anil Dolgun, and Naime Meric Konar. onewaytests: an R package for one-way tests in independent groups designs. The R Journal, 10(1):175–199, July 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-022.

Demirhan:2016:RRP

[Dem16] Haydar Demirhan. rTableICC: an R package for random generation of 22K and RC contingency tables. The R Journal, 8(1):48–63, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-004.

Dorman:2019:SRP

[DEVK19] Michael Dorman, Evyatar Erell, Adi Vulkan, and Itai Kloog. shadow: R package for geometric shadow calculations in an urban environment. The R Journal, 11(1):287–309, June 2019. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2019/RJ-
2019-024.

Dormann:2008:IPP

[DGF08] Carsten F. Dormann, Bernd Gruber, and Jochen Fründ. Introducing the bipartite package: Analysing ecological networks. R News: the Newsletter of the R Project, 8(2): 8–11, October 2008. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Dion:2019:MPF

[DHS19] Charlotte Dion, Simone Hermann, and Adeline Samson. mixedsde: a package to fit mixed stochastic differential equations. The R Journal, 11(1):44–66, June 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-009.

Djatschenko:2019:BRP

[Dja19] Wadim Djatschenko. BondValuation: an R package for fixed coupon bond analysis. The R Journal, 11(2):124–141, December 2019.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2019/RJ-2019-055.

deLeeuw:2001:PRD

[dL01] Jan de Leeuw. Porting R to Darwin/X11 and Mac OS X. R News: the Newsletter of the R Project, 1(3):2–4, September 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Yokota:2018:ACC

[dLN^{+}18] Renata Tiene de Carvalho Yokota, Caspar W. N. Looman, Wilma Johanna Nusselder, Herman Van Oyen, and Geert Molenberghs. addhaz: Contribution of chronic diseases to the disability burden using R. The R Journal, 10(2):75–94, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-055.

Dilba:2007:IRN

[DSH07] Gemechis Dilba, Frank Schaarschmidt, and Ludwig A. Hothorn. Inferences for ratios of normal means. R News: the Newsletter of the R Project, 7(1):20–23, April 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Downey:2017:ARP

[DSN17] Sean S. Downey, Guowei Sun, and Peter Norquest. alineR: an R package for optimizing feature-weighted alignments and linguistic distances. The R Journal, 9(1):138–152, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-005.

Dunning:2006:PRP

[DSTT06] Mark Dunning, Mike Smith, Natalie Thorne, and Simon Tavaré. beadarray: An R package to analyse Illumina BeadArrays. R News: the Newsletter of the R Project, 6(5):17–23, December 2006. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Duchscherer:2018:RRP

[DSU18] Samantha Duchscherer, Robert Stewart, and Marie Urban. revengc: an R package to reverse engineer summarized data. The R Journal, 10(2):114–123, December 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-044.

Duggan:2017:IMB

[Dug17] Jim Duggan. Implementing a metapopulation Bass diffusion model using the R package deSolve. The R Journal, 9(1):153–163, June 2017. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2017/RJ-
2017-006.

Durinck:2006:IBD

[Dur06] Steffen Durinck. Integrating biological data resources into R with biomaRt. R News: the Newsletter of the R Project, 6(5):40–45, December 2006. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Dusa:2018:CCF

[Dus18]  Adrian Dusa.  Consistency cubes: a

fast, efficient method for exact Boolean minimization. The R Journal, 10(2): 357–370, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-080.

deValk:2019:NRP

[dVdMF19] Serge de Valk, Daiane de Mattos, and Pedro Ferreira. Nowcasting: an R package for predicting economic variables using dynamic factor models. The R Journal, 11(1):230–244, June 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-020.

Dag:2016:GRP

[DY16] Osman Dag and Ceylan Yozgatligil. GMDH: an R package for short term forecasting via GMDH-type neural network algorithms. The R Journal, 8(1):379–386, August 2016.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2016/RJ-2016-028.

Dudoit:2002:URA

[DYB02] Sandrine Dudoit, Yee Hwa Yang, and Ben Bolstad. Using R for the analysis of DNA microarray data. R News: the Newsletter of the R Project, 2(1):24–32, March 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Dietrich:2013:RR

[DZK13] André Dietrich, Sebastian Zug, and Jörg Kaiser. The R in robotics. The R Journal, 5(2):117–129, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_dietrich-zug-kaiser.pdf.

Evangelista:2018:GPD

[EB18] Paul F. Evangelista and David Beskow. Geospatial point density. The R Journal, 10(2):347–356, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-061.

Einbeck:2007:NPF

[EHD07] Jochen Einbeck, John Hinde, and Ross Darnell. A new package for fitting random effect models. R News: the Newsletter of the R Project, 7(1):26–30, April 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Enos:2008:TC

[EKN^{+}08] Jeff Enos, David Kane, Arjun Ravi Narayan, Aaron Schwartz, Daniel Suo, and Luyi Zhao. Trade costs. R News: the Newsletter of the R Project, 8(1):10–13, May 2008. CODEN ???? ISSN 1609-3631. URL http://CRAN.

R-project.org/doc/Rnews/.

Edwards:2003:DRD

[EO03] Jonathan Edwards and Paul Oman. Dimensional reduction for data mapping. R News: the Newsletter of the R Project, 3(3):2–7, December 2003. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Eder:2016:SRP

[ERK16] Maciej Eder, Jan Rybicki, and Mike Kestemont. Stylometry with R: a package for computational text analysis. The R Journal, 8(1):107–121, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-007.

Eugster:2013:PPR

[ES13] Manuel J. A. Eugster and Thomas Schlesinger.

osmar: OpenStreetMap and R. The R Jour-

nal, 5(1):53–64, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_eugster-schlesinger.pdf.

Exavier:2020:OSQ

[EZ20] Reginal Exavier and Peter Zeilhofer. OpenLand: Software for quantitative analysis and visualization of land use and cover change. The R Journal, 12(2):359–371, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2021/RJ-2021-021.

Falcon:2006:MPD

[Fal06] Seth Falcon. Modeling package dependencies using graphs. R News: the Newsletter of the R Project, 6(5):8–12, December 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Fay:2010:TSE

[Fay10] Michael P. Fay. Two-sided exact tests and matching confidence intervals for discrete data. The R Journal, 2(1):53–58, June 2010. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2010-1/
RJournal_2010-1_Fay.pdf.

Feinerer:2008:ITM

[Fei08] Ingo Feinerer. An introduction to text mining in R. R News: the Newsletter of the R Project, 8(2):19–22, October 2008. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

FerreiradaSilva:2010:PBC

[Fer10] Adelino Ferreira da Silva. cudaBayesreg: Bayesian computation in CUDA. The R Journal, 2(2):48–55, December 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2010-2/RJournal_
2010-2_Ferreira~da-Silva.pdf.

Feys:2016:NTI

[Fey16] Jos Feys. Nonparametric tests for the interaction in two-way factorial designs using R. The R Journal, 8(1):367–378, August 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-027.

Fox:2013:HTM

[FFW13] John Fox, Michael Friendly, and Sanford Weisberg. Hypothesis tests for multivariate linear models using the car package. The R Journal, 5(1):39–53, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_fox-friendly-weisberg.pdf.

Ferraro:2019:FRP

[FGS19] Maria Brigida Ferraro, Paolo Giordani, and Alessio Serafini. fclust: an R package for fuzzy clustering. The R Journal, 11(1):198–210, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-017.

Friedrich:2019:RBA

[FKP19] Sarah Friedrich, Frank Konietschke, and Markus Pauly. Resampling-based analysis of multivariate data and repeated measures designs with the R package MANOVA.RM. The R Journal, 11(2):380–400, December 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-051.

Fouodo:2018:SVM

[FKW^{+}18] Césaire J. K. Fouodo, Inke R. König, Claus Weihs, Andreas Ziegler, and Marvin N. Wright. Support vector machines for survival analysis with R. The R Journal, 10(1):412–423, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-005.

Fryer:2020:RRP

[FLO20] Daniel Fryer, Ming Li, and Andriy Olenko. rcosmo: R package for analysis of spherical, HEALPix and cosmological data. The R Journal, 12(1):206–225, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-012.

Ferrando:2019:UFB

[FLSNG19] Pere J. Ferrando, Urbano Lorenzo-Seva, and David Navarro-Gonzalez. unival: an FAbased R package for assessing essential unidimensionality using external validity information. The R Journal, 11(1):427–436, June 2019. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2019/RJ-
2019-040.

Fonseca:2018:ARP

[FMMV18] Yuri R. Fonseca, Ricardo P. Masini, Marcelo C. Medeiros, and Gabriel F. R. Vasconcelos. ArCo: an R package to estimate artificial counterfactuals. The R Journal, 10(1): 91–108, July 2018. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2018/RJ-2018-016.

Franck:2016:EIE

[FO16] Christopher T. Franck and Jason A. Osborne. Exploring interaction effects in twofactor studies using the hiddenf package in R. The R Journal, 8(1):159–172, August 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-011.

Fox:2005:PNH

[Fox05] John Fox. Programmer’s niche: How do you spell that number? R News: the Newsletter of the R Project, 5(1):51–55, May 2005. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Fox:2007:ERC

[Fox07] John Fox. Extending the R Commander by “plug-in” packages. R News: the Newsletter of the R Project, 7(3):46–52, December 2007. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Fox:2009:ASO

[Fox09] John Fox. Aspects of the social organization and trajectory of the R Project. The R Journal, 1(2):5–13, December 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2009-2/RJournal_
2009-2_Fox.pdf.

Fraley:2006:MBM

[FR06a] Chris Fraley and Adrian E. Raftery. Modelbased microarray image analysis. R News: the Newsletter of the R Project, 6(5):60–63, December 2006. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Fraley:2006:SAM

[FR06b] Chris Fraley and Adrian E. Raftery. Some applications of model-based clustering in chemistry. R News: the Newsletter of the R Project, 6(3):17–23, August 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

French:2017:AMH

[Fre17] Joshua P. French. autoimage: Multiple heat maps for projected coordinates. The R Journal, 9(1):284–297, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-025.

Fraley:2011:PWF

[FRG^{+}11] Chris Fraley, Adrian Raftery, Tilmann Gneiting, McLean Sloughter, and Veronica Berrocal. Probabilistic weather forecasting in R. The R Journal, 3(1):55–63, June 2011. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2011-1/
RJournal_2011-1_Fraley~et~al.pdf.

Fachada:2016:MRP

[FRL^{+}16] Nuno Fachada, João Rodrigues, Vitor V. Lopes, Rui C. Martins, and Agostinho C. Rosa. micompr: an R package for multivariate independent comparison of observations. The R Journal, 8(2):405–420, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-055.

Fialkowski:2019:SSC

[FT19] Allison Fialkowski and Hemant Tiwari. SimCorrMix: Simulation of correlated data with multiple variable types including continuous and count mixture distributions. The R Journal, 11(1):250–286, June 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-022.

Forcheh:2013:BRP

[FVK^{+}13] Anyiawung Chiara Forcheh, Geert Verbeke, Adetayo Kasim, Dan Lin, Ziv Shkedy, Willem Talloen, Hinrich W. H. Goehlmann, and Lieven Clement. The beadarrayFilter: An R package to filter beads. The R Journal, 5(1):171–181, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_forcheh-verbeke-kasim-etal.pdf.

Fortuna:2020:RHW

[FVS^{+}20] Mariachiara Fortuna, Francesca Vitalini, Mirko Signorelli, Emanuela Furfaro, Federico Marini, Gert Janssenswillen’, Riccardo Porreca, Riccardo L. Rossi, Andrea Guzzo, Roberta Sirovich, Andrea Melloncelli, Lorenzo Salvi, Serena Signorelli, and Filippo Chiarello. e-Rum2020: how we turned a physical conference into a successful virtual event. The R Journal, 12(2):417–425, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2020-2/erum.pdf.

Gaure:2013:PFL

[Gau13] Simen Gaure. lfe: Fitting linear models with multiple factors with many levels. The R Journal, 5(2):104–117, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_gaure.pdf.

Garbade:2006:URS

[GB06] Sven Garbade and Peter Burgard. Using R/Sweave in everyday clinical practice. R News: the Newsletter of the R Project, 6(2): 26–31, May 2006. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Gosiewska:2019:ARP

[GB19] Alicja Gosiewska and Przemysław Biecek. auditor: an R package for model-agnostic visual validation and diagnostics. The R Journal, 11(2):85–98, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-036.

Greenwell:2020:VIP

[GB20] Brandon M. Greenwell and Bradley C. Boehmke. Variable importance plots — an introduction to the vip package. The R Journal, 12(1):343–366, June 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020/RJ-2020-013.

Gutierrez:2017:AIA

[GBB^{+}17] Robert J. Gutierrez, Bradley C. Boehmke, Kenneth W. Bauer, Cade M. Saie, and Trevor J. Bihl. anomalyDetection: Implementation of augmented network log anomaly detection procedures. The R Journal, 9 (2):354–365, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-039.

Gentleman:2002:B

[GC02] Robert Gentleman and Vincent Carey. Bioconductor. R News: the Newsletter of the R Project, 2(1):11–16, March 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Garcia:2016:SDE

[GCG^{+}16] Ángel M. García, Francisco Charte, Pedro González, Cristóbal J. Carmona, and María J. del Jesus. Subgroup discovery with evolutionary fuzzy systems in R: The SDEFSR package. The R Journal, 8(2):307–323, December 2016.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2016/RJ-2016-048.

Gentry:2004:LPP

[GCGG04] Jeff Gentry, Vincent Carey, Emden Gansner, and Robert Gentleman. Laying out pathways with Rgraphviz. R News: the Newsletter of the R Project, 4(2):14–18, September 2004. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Gilbert:2018:JJP

[GD18] Floid R. Gilbert and David B. Dahl. jsr223: a Java platform integration for R with programming languages Groovy, JavaScript, JRuby, Jython, and Kotlin. The R Journal, 10 (2):440–454, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-066.

Gesmann:2011:UGV

[GdC11] Markus Gesmann and Diego de Castillo. Using the Google visualisation API with R. The R Journal, 3(2):40–44, December 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-2/RJournal_
2011-2_Gesmann+de~Castillo.pdf.

Garcia-Donato:2018:BTV

[GDF18] Gonzalo Garcia-Donato and Anabel Forte. Bayesian testing, variable selection and model averaging in linear models using R with BayesVarSel. The R Journal, 10(1):155–174, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-021.

Goncalves-Dosantos:2020:PRP

[GDGJC20] Juan Carlos Gonçalves-Dosantos, Ignacio García-Jurado, and Julián Costa. ProjectManagement: an R package for managing projects. The R Journal, 12(1):419–436, June 2020. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2020/RJ-
2020-009.

Gehrke:2017:RTC

[GDM^{+}17] Matthias Gehrke, Reed Davis, Norman Matloff, Paul Thompson, Tiffany Chen, Emily Watkins, and Laurel Beckett. R teaching column. The R Journal, 9(2): 553–562, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017-2/teaching.pdf.

Graves:2009:PSH

[GDRF09] Spencer Graves, Sundar Dorai-Raj, and Romain François. sos: Searching help pages of R packages. The R Journal, 1 (2):56–59, December 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-2/RJournal_
                                        2009-2_Graves~et~al.pdf.

Godfrey:2014:ASS

[GE14] A. Jonathan R. Godfrey and Robert Erhardt. Addendum to “Statistical software from a blind person’s perspective”. The R Journal, 6(1):182, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2014-1/RJournal_
                                        2014-1_godfrey-erhardt.pdf. See [God13].

Gentleman:2002:SSD

[Gen02] Robert Gentleman. Some strategies for dealing with genomic data. R News: the Newsletter of the R Project, 2(3):22–25, December 2002. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Geraci:2016:QCM

[Ger16] Marco Geraci. Qtools: a collection of models and tools for quantile inference. The R Journal, 8(2):117–138, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-037.

Gunther:2010:PTN

[GF10] Frauke Günther and Stefan Fritsch. neuralnet: Training of neural networks. The R Journal, 2(1):30–38, June 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2010-1/RJournal_
                                        2010-1_Guenther+Fritsch.pdf.

Gerber:2019:ORP

[GF19] Florian Gerber and Reinhard Furrer. optimParallel: an R package providing a parallel version of the L-BFGS-B optimization method. The R Journal, 11(1):352–358, June 2019.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2019/RJ-2019-030.

Gentleman:2002:QP

[GG02] Robert Gentleman and Jeff Gentry. Querying PubMed. R News: the Newsletter of the R Project, 2(2):28–31, June 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.

R-project.org/doc/Rnews/.

Gelfond:2018:SAD

[GGHB18] Jonathan Gelfond, Martin Goros, Brian Hernandez, and Alex Bokov. A system for an accountable data analysis process in R. The R Journal, 10(1):6–21, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-001.

Guevara:2016:DRP

[GHM16] Miguel R. Guevara, Dominik Hartmann, and Marcelo Mendoza. diverse: an R package to analyze diversity in complex systems. The R Journal, 8(2):60–78, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-033.

Gilbert:2004:RPM

[Gil04] Paul Gilbert. R package maintenance. R News: the Newsletter of the R Project, 4(2): 21–24, September 2004. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Goulet:2009:PMD

[GJP09] Vincent Goulet, Michel Jacques, and Mathieu Pigeon. expert: Modeling without data using expert opinion. The R Journal, 1(1):31–36, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2009-1/RJournal_
2009-1_Goulet+et+al.pdf.

Gandy:2017:SRP

[GK17] Axel Gandy and Jan Terje Kvaløy. spcadjust: an R package for adjusting for estimation error in control charts. The R Journal, 9(1):458–476, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-014.

Goksuluk:2016:EIW

[GKZK16] Dincer Goksuluk, Selcuk Korkmaz, Gokmen Zararsiz, and A. Ergun Karaagaoglu. easyROC: an interactive Web-tool for ROC curve analysis using R language environment. The R Journal, 8(2):213–230, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-042.

Grun:2007:PRP

[GL07] Bettina Grün and Friedrich Leisch. FlexMix: An R package for finite mixture modelling. R News: the Newsletter of the R Project, 7 (1):8–13, April 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Greenwell:2018:RDB

[GMBL18] Brandon M. Greenwell, Andrew J. Mc-Carthy, Bradley C. Boehmke, and Dungang Liu. Residuals and diagnostics for binary and ordinal regression models: an introduction to the sure package. The R Journal, 10(1):381–394, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-004.

Godfrey:2013:SAB

[God13] A. Jonathan R. Godfrey. Statistical analysis from a blind person’s perspective. The R Journal, 5(1):73–80, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_

2013-1_godfrey.pdf. See addendum [GE14].

Gorjanc:2007:WUV

[Gor07] Gregor Gorjanc. Working with unknown values. R News: the Newsletter of the R Project, 7(1):24–26, April 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Gorjanc:2008:USL

[Gor08] Gregor Gorjanc. Using Sweave with LyX. R News: the Newsletter of the R Project, 8 (1):2–9, May 2008. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Grothendieck:2004:RHD

[GP04] Gabor Grothendieck and Thomas Petzoldt. R Help Desk: Date and time classes in R. R News: the Newsletter of the R Project, 4 (1):29–32, June 2004. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Goulet:2008:SML

[GP08] Vincent Goulet and Mathieu Pigeon. Statistical modeling of loss distributions using actuar. R News: the Newsletter of the R Project, 8(1):34–40, May 2008. CODEN ???? ISSN 1609-3631. URL http://CRAN.

R-project.org/doc/Rnews/.

Gu:2019:RRG

[GPB19] Mengyang Gu, Jesus Palomo, and James O. Berger. RobustGaSP: Robust Gaussian stochastic process emulation in R. The R Journal, 11(1):112–136, June 2019. ISSN

2073-4859.  URL https://journal.r-
project.org/archive/2019/RJ-2019-011.

Graler:2016:STI

[GPH16] Benedikt Gräler, Edzer Pebesma, and Gerard Heuvelink. Spatio-temporal interpolation using gstat. The R Journal, 8 (1):204–218, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-014.

Genuer:2015:VRP

[GPTM15] Robin Genuer, Jean-Michel Poggi, and Christine Tuleau-Malot. VSURF: an R package for variable selection using random forests. The R Journal, 7(2):19–33, December 2015.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2015/RJ-2015-018.

Grayling:2014:PRP

[Gra14] Michael J. Grayling. phaseR: An R package for phase plane analysis of autonomous ODE systems. The R Journal, 6(2):43–51, December 2014. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2014-2/RJournal_2014-2_grayling.pdf.

Greenwell:2017:PRP

[Gre17] Brandon M. Greenwell. pdp: an R package for constructing partial dependence plots. The R Journal, 9(1):421–436, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-016.

Gross:2003:VIF

[Gro03] Jürgen Groß. Variance inflation factors. R News: the Newsletter of the R Project, 3(1): 13–15, June 2003. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Grun:2015:Ea

[Grü15a] Bettina Grün. Editorial. The R Journal, 7(1):3, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2015-1/RJournal_
                                        2015-1_editorial.pdf.

Grun:2015:Eb

[Grü15b] Bettina Grün. Editorial. The R Journal, 7(2):4, December 2015. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2015-2/editorial.pdf.

Greenwell:2014:PRP

[GS14] Brandon M. Greenwell and Christine M. Schubert Kabban. investr: An R package for inverse estimation. The R Journal, 6(1):90–100, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_greenwell-kabban.pdf.

Giner:2016:SPC

[GS16] Göknur Giner and Gordon K. Smyth. statmod: Probability calculations for the inverse Gaussian distribution. The R Journal, 8(1):339–351, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-024.

Gonzalez-Sales:2020:APD

[GSBT^{+}20] Mario González-Sales, Olivier Barrière, Pierre Olivier Tremblay, Guillaume Bonnefois, Julie Desrochers, and Fahima Nekka. Assembling pharmacometric datasets in R — the puzzle package. The R Journal, 12(2):107– 119, December 2020. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2021/RJ-2021-006.

Gu:2014:SWR

[GSHB14] Xiangdong Gu, David Shapiro, Michael D. Hughes, and Raji Balasubramanian. Stratified Weibull regression model for intervalcensored data. The R Journal, 6(1):31–40, June 2014. CODEN ???? ISSN 2073-

4859. URL http://journal.r-project.
org/archive/2014-1/RJournal_2014-1_gu-
shapiro-hughes-etal.pdf.

Giles:2019:IRP

[GSL19] John R. Giles, Henrik Salje, and Justin Lessler. The IDSpatialStats R package: Quantifying spatial dependence of infectious disease spread. The R Journal, 11 (2):308–327, December 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-043.

Gesmann:2014:CRR

[GT14] Markus Gesmann and Andreas Tsanakas. Conference report: R in insurance 2014. The R Journal, 6(2):185–186, December 2014. CO-DEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2014-2/
                                        RJournal_2014-2_gesmann-tsanakas.pdf.

Guha:2006:GUV

[Guh06] Rajarshi Guha. Generating, using and visualizing molecular information in R. R News: the Newsletter of the R Project, 6(3):28–33, August 2006. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Gandrud:2017:SPL

[GW17] Christopher Gandrud and Laron K. Williams. Simulating probabilistic long-term effects in models with temporal dependence. The R Journal, 9(2):401–408, December 2017.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2017/RJ-2017-059.

Guazzelli:2009:POS

[GZLW09] Alex Guazzelli, Michael Zeller, Wen-Ching Lin, and Graham Williams. PMML: An open standard for sharing models. The R Journal, 1(1):60–65, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-1/RJournal_
                                        2009-1_Guazzelli+et+al.pdf.

Hahsler:2017:AIV

[Hah17] Michael Hahsler. arulesViz: Interactive visualization of association rules with R. The R Journal, 9(2):163–175, December 2017.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2017/RJ-2017-047.

Hankin:2005:RMR

[Han05] Robin K. S. Hankin. Recreational mathematics with R: introducing the “magic” package. R News: the Newsletter of the R Project, 5 (1):48–51, May 2005. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Hankin:2006:EPR

[Han06a] Robin K. S. Hankin. Electrical properties of resistor networks. R News: the Newsletter of the R Project, 6(2):52–53, May 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Hankin:2006:NDA

[Han06b] Robin K. S. Hankin. Normed division algebras with R: Introducing the onion package. R News: the Newsletter of the R Project, 6 (2):49–52, May 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Hankin:2006:SFR

[Han06c] Robin K. S. Hankin. Special functions in R: introducing the gsl package. R News: the Newsletter of the R Project, 6(4):24–26, October 2006. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Hankin:2007:VLN

[Han07a] Robin K. S. Hankin. Very large numbers in R: Introducing package Brobdingnag. R News: the Newsletter of the R Project, 7(3): 15–16, December 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Hansen:2007:PFO

[Han07b] Ben B. Hansen. Optmatch: Flexible, optimal matching for observational studies. R News: the Newsletter of the R Project, 7(2):

18–24, October 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Hankin:2008:PNM

[Han08] Robin K. S. Hankin. Programmers’ niche: Multivariate polynomials in R. R News: the Newsletter of the R Project, 8(1):41–45, May 2008. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Hankin:2015:CMG

[Han15a] Robin Hankin. The complex multivariate Gaussian distribution. The R Journal, 7(1):73–80, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_hankin.pdf.

Hankin:2015:NEG

[Han15b] Robin K. S. Hankin. Numerical evaluation of the Gauss hypergeometric function with the hypergeo package. The R Journal, 7(2):81–88, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-022.

Hankin:2017:PRD

[Han17] Robin K. S. Hankin. Partial rank data with the hyper2 package: Likelihood functions for generalized Bradley–Terry models. The R Journal, 9(2):429–439, December 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-061.

Hoeltgebaum:2019:HRP

[HB19] Henrique Hoeltgebaum and Heather Battey. HCmodelSets: an R package for specifying sets of well-fitting models in high dimensions. The R Journal, 11(2):370–379, December 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-057.

Hothorn:2001:MGP

[HBG01] Torsten Hothorn, Frank Bretz, and Alan Genz. On multivariate t and Gauß probabilities in R. R News: the Newsletter of the R Project, 1(2):27–29, June 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.

R-project.org/doc/Rnews/.

Hare:2015:MDR

[HBH15] Eric Hare, Andreas Buja, and Heike Hofmann. Manipulation of discrete random variables with discreteRV. The R Journal, 7(1):185–194, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_hare-buja-hofmann.pdf.

Herve:2011:PMG

[Her11] Maxime Hervé. GrapheR: a multiplatform GUI for drawing customizable graphs in R. The R Journal, 3(2):45–53, December 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-2/RJournal_
2011-2_Herve.pdf.

Hoff:2014:FLI

[HGF14] Rune Hoff, Jon Michael Gran, and Daniel Farewell. Farewell’s linear increments model for missing data: The FLIM package. The R Journal, 6(2):137–150, December 2014. CO-DEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/archive/2014-
2/RJournal_2014-2_hoff-gran-farewell.
pdf.

Happ:2018:HRP

[HHB18] Martin Happ, Solomon W. Harrar, and Arne C. Bathke. HRM: an R package for analysing high-dimensional multi-factor repeated measures. The R Journal, 10(1): 534–548, July 2018. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2018/RJ-2018-032.

Hino:2017:IID

[Hin17] Hideitsu Hino. ider: Intrinsic dimension estimation with R. The R Journal, 9 (2):329–341, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-054.

Hahsler:2019:ACR

[HJKK19] Michael Hahsler, Ian Johnson, Tomáš Kliegr, and Jaroslav Kucharˇ.AssociativeclassificationinR : arc, arulesCBA, andrCBA. TheRJournal, 11

(2) : 254 − −290, December2019.ISSN2073 −

                                        4859.URLhttpsjournalrprojectorgarchive2019RJ2019048.

Hothorn:2002:MSR

[HL02] Torsten Hothorn and Berthold Lausen. Maximally selected rank statistics in R. R News: the Newsletter of the R Project, 2(1):3–5, March 2002. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Hornik:2017:CCa

[HLZ17a] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 9(1):505–507, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017-1/cran.pdf.

Hornik:2017:CCb

[HLZ17b] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 9 (2):564–566, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017-2/cran.pdf.

Hornik:2018:CCa

[HLZ18a] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 10(1):556–559, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018-1/cran.pdf.

Hornik:2018:CCb

[HLZ18b] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 10 (2):584–586, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018-2/cran.pdf.

Hornik:2019:CCa

[HLZ19a] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 11(1):438–440, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019-1/cran.pdf.

Hornik:2019:CCb

[HLZ19b] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 11 (2):444–446, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019-2/cran.pdf.

Hornik:2020:CCa

[HLZ20a] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 12(1):485–497, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020-1/cran.pdf.

Hornik:2020:CCb

[HLZ20b] Kurt Hornik, Uwe Ligges, and Achim Zeileis. Changes on CRAN. The R Journal, 12 (2):407–409, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020-2/cran.pdf.

Hornik:2011:WYS

[HM11] Kurt Hornik and Duncan Murdoch. Watch your spelling! The R Journal, 3(2): 22–28, December 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-2/RJournal_
2011-2_Hornik+Murdoch.pdf.

Hladka:2020:DGL

[HM20] Adéla Hladká and Patrícia Martinková. difNLR: Generalized logistic regression models for DIF and DDF detection. The R Journal, 12 (1):300–323, June 2020. ISSN 2073-4859. URL

                                        https://journal.r-project.org/archive/
                                        2020/RJ-2020-014.

Hmamouche:2020:NRP

[Hma20] Youssef Hmamouche. NlinTS: an R package for causality detection in time series. The R Journal, 12(1):21–31, June 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020/RJ-2020-016.

Hornik:2012:WDW

[HMZ12] Kurt Hornik, Duncan Murdoch, and Achim Zeileis. Who did what? The roles of R package authors and how to refer to them. The R Journal, 4(1):64–69, June 2012. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2012-1/RJournal_
                                        2012-1_Hornik~et~al.pdf.

Ha:2012:PPF

[HNL12] Il Do Ha, Maengseok Noh, and Youngjo Lee. frailtyHL: A package for fitting frailty models with H-likelihood. The R Journal, 4(2):28– 37, December 2012. CODEN ???? ISSN 2073-

                                        4859. URL http://journal.r-project.
                                        org/archive/2012-2/RJournal_2012-2_Do~
                                        Ha-et~al.pdf.

Hocking:2019:CNO

[Hoc19] Toby Dylan Hocking. Comparing namedCapture with other R packages for regular expressions. The R Journal, 11(2):328–346, December 2019. ISSN 2073-4859. URL https://

                                        journal.r-project.org/archive/2019/RJ-
                                        2019-050.

Hofert:2013:SMD

[Hof13] Marius Hofert. On sampling from the multivariate t distribution. The R Journal, 5 (2):129–137, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_hofert.pdf.

Hojsgaard:2006:PP

[Høj06] Søren Højsgaard. The doBy package. R News: the Newsletter of the R Project, 6(2): 47–49, May 2006. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Holmes:2006:RFM

[Hol06] Susan Holmes. Review of Fionn Murtagh’s book: Correspondence Analysis and Data Coding with Java and R. R News: the Newsletter of the R Project, 6(4):41–43, October 2006. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Hornik:2013:RFNa

[Hor13a] Kurt Hornik. R Foundation news. The R Journal, 5(1):220–221, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_r-foundation.pdf.

Hornik:2013:RFNb

[Hor13b] Kurt Hornik. R Foundation news. The R Journal, 5(2):161–162, December 2013. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2013-2/
RJournal_2013-2_r-foundation.pdf.

Hornik:2014:RFN

[Hor14] Kurt Hornik. R Foundation news. The R Journal, 6(1):183, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_r-foundation.pdf.

Hothorn:2001:ERT

[Hot01] Torsten Hothorn. On exact rank tests in R. R News: the Newsletter of the R Project, 1(1): 11–12, January 2001. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Hothorn:2016:RFN

[Hot16] Torsten Hothorn. R Foundation news. The R Journal, 8(2):506, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016-2/foundation.pdf.

Hothorn:2017:RFNa

[Hot17a] Torsten Hothorn. R Foundation news. The R Journal, 9(1):498–500, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017-1/foundation.pdf.

Hothorn:2017:RFNb

[Hot17b] Torsten Hothorn. R Foundation news. The R Journal, 9(2):563, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017-2/foundation.pdf.

Hothorn:2018:RFNa

[Hot18a] Torsten Hothorn. R Foundation news. The R Journal, 10(1):555, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018-1/foundation.pdf.

Hothorn:2018:RFNb

[Hot18b]  Torsten Hothorn.  R Foundation news.

The R Journal, 10(2):583, December 2018.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2018-2/foundation.
                                        pdf.

Hothorn:2019:RFNa

[Hot19a] Torsten Hothorn. R Foundation news. The R Journal, 11(1):583, June 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019-1/foundation.pdf.

Hothorn:2019:RFNb

[Hot19b]  Torsten Hothorn.  R Foundation news.

The R Journal, 11(2):443, December 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019-2/foundation.
pdf.

Hothorn:2020:RFNa

[Hot20a] Torsten Hothorn. R Foundation news. The R Journal, 12(1):498, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020-1/foundation.pdf.

Hothorn:2020:RFNb

[Hot20b] Torsten Hothorn. R Foundation news. The R Journal, 12(2):413–416, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2020-2/foundation.
pdf.

Hofmeyr:2019:PRP

[HP19] David P. Hofmeyr and Nicos G. Pavlidis. PPCI: an R package for cluster identification using projection pursuit. The R Journal, 11 (2):152–170, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-046.

Hu:2016:PRP

[HQ16] Chenyue W. Hu and Amina A. Qutub. progenyClust: an R package for progeny clustering. The R Journal, 8(1):328–338, August 2016. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2016/RJ-
2016-023.

Hayfield:2007:PPK

[HR07] Tristen Hayfield and Jeffrey S. Racine. The np package: Kernel methods for categorical and continuous data. R News: the Newsletter of the R Project, 7(2):35–43, October 2007. CO-DEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Hino:2015:MPC

[HTM15] Hideitsu Hino, Ken Takano, and Noboru Murata. mmpp: a package for calculating similarity and distance metrics for simple and marked temporal point processes. The R Journal, 7(2):237–248, December 2015. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2015/RJ-2015-033.

Hofmann:2013:LGT

[HUC13] Heike Hofmann, Antony Unwin, and Dianne Cook. Let graphics tell the story — Datasets in R. The R Journal, 5(1):117–130, June 2013. CODEN ???? ISSN 2073-4859. URL http:

                                        //journal.r-project.org/archive/2013-
                                        1/RJournal_2013-1_hofmann-unwin-cook.
                                        pdf.

Hughes:2014:PPF

[Hug14] John Hughes. ngspatial: A package for fitting the centered autologistic and sparse spatial generalized linear mixed models for areal data. The R Journal, 6(2):81–95, December 2014. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2014-2/RJournal_2014-2_hughes.pdf.

Heisterkamp:2017:UNP

[HvWDM17] Simon H. Heisterkamp, Engelbertus van Willigen, Paul-Matthias Diderichsen, and John Maringwa. Update of the nlme package to allow a fixed standard deviation of the residual error. The R Journal, 9(1):239–251, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-010.

Hargarten:2020:MMI

[HW20] Paul M. Hargarten and David C. Wheeler. miWQS: Multiple imputation using weighted quantile sum regression. The R Journal, 12 (2):226–250, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-014.

Holmes:2012:PMA

[HWW12] Elizabeth E. Holmes, Eric J. Ward, and Kellie Wills. MARSS: Multivariate autoregressive state-space models for analyzing time-series data. The R Journal, 4(1):11–19, June 2012. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2012-1/
RJournal_2012-1_Holmes~et~al.pdf.

Hinz:2018:RRP

[HY18] Juri Hinz and Jeremy Yee. rcss: R package for optimal convex stochastic switching. The R Journal, 10(2):38–54, December 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-054.

Hyndman:2011:GUT

[Hyn11] Rob J. Hyndman. Giving a useR! talk. The R Journal, 3(1):69–71, June 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-1/RJournal_
2011-1_Hyndman.pdf.

Hornik:2013:CCa

[HZ13a] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 5(1):239–264, June 2013. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2013-1/RJournal_2013-1_cran.pdf.

Hornik:2013:CCb

[HZ13b] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 5(2):166–192, December 2013. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2013-2/RJournal_2013-2_cran.pdf.

Hornik:2014:CCa

[HZ14a] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 6(1):186–220, June 2014. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2014-1/RJournal_2014-1_cran.pdf.

Hornik:2014:CCb

[HZ14b] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 6(2):192–223, December 2014. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2014-2/RJournal_2014-2_cran.pdf.

Hornik:2015:CCa

[HZ15a] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 7(1):201–226, June 2015. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2015-1/RJournal_2015-1_cran.pdf.

Hornik:2015:CCb

[HZ15b] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 7(2):298–339, December 2015. ISSN 2073-4859. URL https:/

                                        /journal.r-project.org/archive/2015-2/
                                        cran.pdf.

Hornik:2016:CCa

[HZ16a] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 8(1):402–403, August 2016. ISSN 2073-4859. URL https:/

                                        /journal.r-project.org/archive/2016-1/
                                        cran.pdf.

Hornik:2016:CCb

[HZ16b] Kurt Hornik and Achim Zeileis. Changes on CRAN. The R Journal, 8(2):507–509, December 2016. ISSN 2073-4859. URL https:/

                                        /journal.r-project.org/archive/2016-2/
                                        cran.pdf.

Imdadullah:2016:MRP

[IAA16] Muhammad Imdadullah, Muhammad Aslam, and Saima Altaf. mctest: an R package for detection of collinearity among regressors. The R Journal, 8(2):495–505, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-062.

Imdadullah:2017:LCR

[IAA17] Muhammad Imdadullah, Muhammad Aslam, and Saima Altaf. liureg: a comprehensive R package for the Liu estimation of linear regression model with collinear regressors. The R Journal, 9(2):232–247, December 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-048.

Iacus:2001:PRM

[Iac01] Stefano M. Iacus. Porting R to the Macintosh. R News: the Newsletter of the R Project, 1 (1):13, January 2001. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Iturbide:2018:TUS

[IBG18] M. Iturbide, J. Bedia, and J. M. Gutiérrez. Tackling uncertainties of species distribution model projections with package mopa. The R Journal, 10(1):122–139, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-019.

Im:2018:FRP

[ICK18] Jongho Im, In Ho Cho, and Jae Kwang Kim. FHDI: an R package for fractional hot deck imputation. The R Journal, 10(1):140–154, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-020.

Ishwaran:2007:RSF

[IK07] Hemant Ishwaran and Udaya B. Kogalur. Random survival forests for R. R News: the Newsletter of the R Project, 7(2):25–31, October 2007. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Ishwaran:2010:PPV

[IKR10] Hemant Ishwaran, Udaya B. Kogalur, and J. Sunil Rao. spikeslab: Prediction and variable selection using spike and slab regression. The R Journal, 2(2):68–73, December 2010. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2010-2/
RJournal_2010-2_Ishwaran~et~al.pdf.

Irigoien:2016:WDB

[IMA16] Itziar Irigoien, Francesc Mestres, and Concepcion Arenas. Weighted distance based discriminant analysis: The R package WeDiBaDis. The R Journal, 8(2):434–450, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-057.

Ieva:2019:RPR

[IPRT19] Francesca Ieva, Anna Maria Paganoni, Juan Romo, and Nicholas Tarabelloni. roahd package: Robust analysis of high dimensional data. The R Journal, 11(2):291–307, December 2019. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2019/RJ-
2019-032.

Inan:2017:PRP

[IW17] Gul Inan and Lan Wang. PGEE: an R package for analysis of longitudinal data with high-dimensional covariates. The R Journal, 9(1):393–402, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-030.

Jelihovschi:2018:LRP

[JA18] Enio G. Jelihovschi and Ivan Bezerra Allaman. lba: an R package for latent budget analysis. The R Journal, 10(1):269–287, July 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-026.

Jara:2007:ABN

[Jar07] Alejandro Jara. Applied Bayesian non- and semi-parametric inference using DPpackage. R News: the Newsletter of the R Project, 7(3): 17–26, December 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Jones:2019:RRP

[JCWL19] Kasey Jones, Rob Chew, Allison Witman, and Yiyan Liu. rollmatch: an R package for rolling entry matching. The R Journal, 11(2):243–253, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-005.

Jacques:2014:PRP

[JGB14] Julien Jacques, Quentin Grimonprez, and Christophe Biernacki. Rankcluster: An R package for clustering multivariate partial rankings. The R Journal, 6(1): 101–110, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_jacques-grimonprez-biernacki.
pdf.

Joblin:2016:ISA

[JM16] Mitchell Joblin and Wolfgang Mauerer. An interactive survey application for validating social network analysis techniques. The R Journal, 8(1):149–158, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-010.

Jimenez:2017:QPA

[JM17] Pedro Cañadilla Jiménez and Yolanda Román Montoya. queueing: a package for analysis of queueing networks and models in R. The R Journal, 9(2):116–126, December 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-051.

Jones:2018:CRP

[Jon18a] Edmund Jones. cchs: an R package for stratified case-cohort studies. The R Journal, 10(1):484–494, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-012.

Jones:2018:MMC

[Jon18b] Zachary M. Jones. mmpf: Monte-Carlo methods for prediction functions. The R Journal, 10(1):56–60, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-038.

Jensen:2018:PRP

[JP18a] Cathrine Ulla Jensen and Toke Emil Panduro. PanJen: an R package for ranking transformations in a linear regression. The R Journal, 10(1):109–121, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-018.

Jordan:2018:DST

[JP18b] Soren Jordan and Andrew Q. Philips. Dynamic simulation and testing for singleequation cointegrating and stationary autoregressive distributed lag models. The R Journal, 10(2):469–488, December 2018.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2018/RJ-2018-076.

Jurka:2012:PRP

[Jur12] Timothy P. Jurka. maxent: An R package for low-memory multinomial logistic regression with support for semi-automated text classification. The R Journal, 4(1):56–59, June 2012. CODEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2012-1/
                                        RJournal_2012-1_Jurka.pdf.

Jain:2006:BP

[JW06] Nitin Jain and Gregory R. Warnes. Balloon plot. R News: the Newsletter of the R Project, 6(2):35–38, May 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Kahle:2013:PMP

[Kah13] David Kahle. mpoly: Multivariate polynomials in R. The R Journal, 5(1):162–171, June 2013. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2013-1/RJournal_2013-1_kahle.pdf.

Kane:2019:E

[Kan19] Michael J. Kane. Editorial. The R Journal, 11(2):4–5, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019-2/editorial.pdf.

Kane:2020:Ea

[Kan20a] Michael J. Kane. Editorial. The R Journal, 12(1):4–5, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020-1/editorial.pdf.

Kane:2020:Eb

[Kan20b] Michael J. Kane. Editorial. The R Journal, 12(2):4–5, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020-2/editorial.pdf.

Kaptein:2014:PDT

[Kap14] Maurits Kaptein. RStorm: Developing and testing streaming algorithms in R. The R Journal, 6(1):123–132, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_kaptein.pdf.

Koziol:2014:PPA

[KB14] Natalie A. Koziol and Christopher R. Bilder. MRCV: A package for analyzing categorical variables with multiple response options. The R Journal, 6(1):144–150, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_koziol-bilder.pdf.

Kostov:2013:MFA

[KBBH13] Belchin Kostov, Mónica Bécue-Bertaut, and François Husson. Multiple factor analysis on contingency tables in the FactoMineR package. The R Journal, 5(1):29–39, June 2013. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/archive/2013-
1/RJournal_2013-1_kostov-becuebertaut-
husson.pdf.

Kostov:2015:CAG

[KBBH15] Belchin Kostov, Mónica Bécue-Bertaut, and François Husson. Correspondence analysis on generalised aggregated lexical table (CA-GALT) in the FactoMineR package. The R Journal, 7(1):109–117, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_kostov-becuebertaut-husson.pdf.

Kane:2020:ABT

[KCK^{+}20] Michael J. Kane, Nan Chen, Alexander M. Kaizer, Xun Jiang, H. Amy Xia, and Brian P. Hobbs. Analyzing basket trials under multisource exchangeability assumptions. The R Journal, 12(2):342–358, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2021/RJ-2021-020.

Kane:2006:AEP

[KE06] David Kane and Jeff Enos. Analysing equity portfolios in R. R News: the Newsletter of the R Project, 6(2):13–21, May 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Koitka:2016:NGA

[KF16] Sven Koitka and Christoph M. Friedrich. nmfgpu4R: GPU-accelerated computation of the non-negative matrix factorization (NMF) using CUDA capable hardware. The R Journal, 8(2):382–392, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-053.

Kerman:2006:BDA

[KG06] Jouni Kerman and Andrew Gelman. Bayesian data analysis using R. R News: the Newsletter of the R Project, 6(1):21–24, March 2006. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Korkmaz:2014:PRP

[KGZ14] Selcuk Korkmaz, Dincer Goksuluk, and Gokmen Zararsiz. MVN: An R package for assessing multivariate normality. The R Journal, 6(2):151–162, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_korkmaz-goksuluk-zararsiz.pdf.

Kropko:2019:CRP

[KH19] Jonathan Kropko and Jeffrey J. Harden. coxed: an R package for computing durationbased quantities from the Cox proportional hazards model. The R Journal, 11 (2):38–45, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-042.

Kipp:2019:CRD

[KLC19] Michael Kipp, Ursula Laa, and Dianne Cook. Connecting R with D3 for dynamic graphics, to explore multivariate data with tours. The R Journal, 11(1):245–249, June 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-002.

Kim:2020:UAN

[KLK20] Dongshin Kim, Sangin Lee, and Sunghoon Kwon. A unified algorithm for the non-convex penalized estimation: The ncpen package.

The R Journal, 12(2):43–60, December 2020.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2021/RJ-2021-003.

Kross:2020:AAR

[KLM20] Sean Kross, Jeffrey T. Leek, and John Muschelli. ari: The Automated R Instructor. The R Journal, 12(1):258–265, June 2020.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2020/RJ-2020-023.

Kane:2011:AEL

[KLN11] David Kane, Andrew Liu, and Khanh Nguyen. Analyzing an electronic limit order book. The R Journal, 3(1):64–68, June 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2011-1/RJournal_
                                        2011-1_Kane~et~al.pdf.

Koczkodaj:2018:RPS

[KLWD18] Waldemar W. Koczkodaj, Feng Li, and Alicja

                                        Wolny-Dominiak. RatingScaleReduction

package: stepwise rating scale item reduction without predictability loss. The R Journal, 10(1):43–55, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-035.

Katki:2008:SAC

[KM08] Hormuzd A. Katki and Steven D. Mark. Survival analysis for cohorts with missing covariate information. R News: the Newsletter of the R Project, 8(1):14–19, May 2008. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Kloke:2012:PRB

[KM12] John D. Kloke and Joseph W. McKean. Rfit: Rank-based estimation for linear models. The R Journal, 4(2):57–64, December 2012. CO-DEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2012-2/
                                        RJournal_2012-2_Kloke+McKean.pdf.

Kalibera:2020:CR

[KMH20] Tomas Kalibera, Sebastian Meyer, and Kurt Hornik. Changes in R 3.6–4.0. The R Journal, 12(2):402–406, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020-2/core.pdf.

Kasprzak:2020:SYS

[KMKT20] Peter Kasprzak, Lachlan Mitchell, Olena Kravchuk, and Andy Timmins. Six years of Shiny in research — collaborative development of Web tools in R. The R Journal, 12(2):155–162, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-004.

Kim:2009:PPE

[KO09] Donghoh Kim and Hee-Seok Oh. EMD: A package for empirical mode decomposition and Hilbert spectrum. The R Journal, 1(1):40–46, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2009-1/RJournal_
2009-1_Kim+Oh.pdf.

Koenker:2006:PLB

[Koe06] Roger Koenker. Parametric links for binary response. R News: the Newsletter of the R Project, 6(4):32–34, October 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Kolb:2019:UWS

[Kol19] Jan-Philipp Kolb. Using Web services to work with geodata in R. The R Journal, 11(2):6–23, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-041.

Komsta:2006:PDO

[Kom06] Łukasz Komsta. Processing data for outliers. R News: the Newsletter of the R Project, 6 (2):10–13, May 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Kosmidis:2008:PRP

[Kos08] Ioannis Kosmidis. The profileModel R package: Profiling objectives for models with linear predictors. R News: the Newsletter of the R Project, 8(2):12–18, October 2008. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Knaus:2009:EPC

[KPBS09] Jochen Knaus, Christine Porzelius, Harald Binder, and Guido Schwarzer. Easier parallel computing in R with snowfall and sfCluster. The R Journal, 1 (1):54–59, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2009-1/RJournal_
2009-1_Knaus+et+al.pdf.

Keyes:2016:RPA

[KRJ16] Oliver Keyes, Bob Rudis, and Jay Jacobs. R packages to aid in handling Web access logs. The R Journal, 8(1):360–366, August 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-026.

Kraemer:2018:DCU

[KRM18] Guido Kraemer, Markus Reichstein, and Miguel D. Mahecha. dimRed and coRanking — unifying dimensionality reduction in R. The R Journal, 10(1):342–358, July 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-039.

Kawaguchi:2020:FSI

[KSLS20] Eric S. Kawaguchi, Jenny I. Shen, Gang Li, and Marc A. Suchard. A fast and scalable implementation method for competing risks data with the R package fastcmprsk. The R Journal, 12(2):163–172, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2021/RJ-2021-010.

Kuhn:2006:SOD

[Kuh06] Max Kuhn. Sweave and the Open Document Format — the odfWeave package. R News: the Newsletter of the R Project, 6(4):2–8, October 2006. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Kahle:2013:PSV

[KW13] David Kahle and Hadley Wickham. ggmap: Spatial visualization with ggplot2. The R Journal, 5(1):144–162, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_kahle-wickham.pdf.

Koohafkan:2015:OCC

[KY15] Michael C. Koohafkan and Bassam A. Younis. Open-channel computation with R. The R Journal, 7(2):249–262, December 2015.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2015/RJ-2015-034.

Kim:2018:MRP

[KZDZ18] Juhyun Kim, Yiwen Zhang, Joshua Day, and Hua Zhou. MGLM: an R package for multivariate categorical data analysis. The R Journal, 10(1):73–90, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-015.

Levy:2018:CRS

[LAL18] Ariel Levy, Luciane F. Alcoforado, and Orlando Celso Longo. Conference report: SER III. The R Journal, 10 (2):567–571, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018-2/serIII.pdf.

Lang:2001:SCC

[Lan01a] Duncan Temple Lang. In search of C/C++ & FORTRAN routines. R News: the Newsletter of the R Project, 1(3):20–23, September 2001. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Lang:2001:UXS

[Lan01b] Duncan Temple Lang. Using XML for statistics: The XML package. R News: the Newsletter of the R Project, 1(1):24–27, January 2001. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Lang:2017:CFA

[Lan17] Michel Lang. checkmate: Fast argument checks for defensive R programming. The R Journal, 9(1):437–445, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-028.

Lawson:2012:PPC

[Law12] John Lawson. Vdgraph: A package for creating variance dispersion graphs. The R Journal, 4(1):41–44, June 2012. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2012-1/RJournal_
                                        2012-1_Lawson.pdf.

Lawrence:2016:Ea

[Law16a] Michael Lawrence. Editorial. The R Journal, 8(1):4, August 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016-1/editorial.pdf.

Lawrence:2016:Eb

[Law16b] Michael Lawrence. Editorial. The R Journal, 8(2):4, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016-2/editorial.pdf.

Lombardo:2016:VSC

[LB16] Rosaria Lombardo and Eric J. Beh. Variants of simple correspondence analysis. The R Journal, 8(2):167–184, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-039.

Lipsitz:2016:QNR

[LBCFV16] Michael Lipsitz, Alexandre Belloni, Victor Chernozhukov, and Iván Fernández-Val. quantreg.nonpar: an R package for performing nonparametric series quantile regression. The R Journal, 8(2):370–381, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-052.

Long:2006:GNT

[LC06] Li Long and Vince Carey. Graphs and networks: Tools in Bioconductor. R News: the Newsletter of the R Project, 6(5):2–7, December 2006. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Lee:2015:BBS

[LC15] Kuo-Jung Lee and Ray-Bing Chen. BSGS: Bayesian Sparse Group Selection. The R Journal, 7(2):122–133, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-025.

Liu:2020:NRP

[LCT20] Xialu Liu, Rong Chen, and Ruey Tsay. NTS: an R package for nonlinear time series analysis. The R Journal, 12(2): 266–292, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-016.

LeBauer:2013:TPD

[LDB13] David LeBauer, Michael Dietze, and Ben Bolker. Translating probability distributions: From R to BUGS and back again. The R Journal, 5(1):207–210, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_lebauer-dietze-bolker.pdf.

Li:2020:RRP

[LDC20] Hesen Li, Hakan Demirtas, and Ruizhe Chen. RNGforGPD: an R package for generation of univariate and multivariate generalized Poisson data. The R Journal, 12 (2):120–133, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-007.

Lockwood:2003:URE

[LDM03] J. R. Lockwood, Harold Doran, and Daniel F. McCaffrey. Using R for estimating longitudinal student achievement models. R News: the Newsletter of the R Project, 3(3):17–23, December 2003. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Lovelace:2018:SPT

[LE18] Robin Lovelace and Richard Ellison. stplanr: a package for transport planning. The R Journal, 10(2):7–23, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-053.

Lecoutre:2003:PP

[Lec03] Eric Lecoutre. The R2HTML package. R News: the Newsletter of the R Project, 3(3):33–36, December 2003. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Leeper:2014:ARR

[Lee14] Thomas J. Leeper. Archiving reproducible research with R and Dataverse. The R Journal, 6(1):151–158, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_leeper.pdf.

Leeper:2016:CDP

[Lee16] Thomas J. Leeper. Crowdsourced data preprocessing with R and Amazon Mechanical Turk. The R Journal, 8(1):276–288, August 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-020.

Leisch:2002:SPM

[Lei02] Friedrich Leisch. Sweave, Part I: Mixing R and LaTEX. R News: the Newsletter of the R Project, 2(3):28–31, December 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Leisch:2003:PPI

[Lei03] Friedrich Leisch. Sweave, Part II: Package vignettes. R News: the Newsletter of the R Project, 3(2):21–24, October 2003. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Leisch:2007:RRB

[Lei07] Friedrich Leisch. Review of “The R Book”. R News: the Newsletter of the R Project, 7(2): 53–54, October 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Lemon:2006:P

[Lem06] Jim Lemon. plotrix. R News: the Newsletter of the R Project, 6(4):8–12, October 2006. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Lenth:2015:ETP

[Len15] Russell V. Lenth. Estimability tools for package developers. The R Journal, 7 (1):195–199, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_lenth.pdf.

Ligges:2008:RHD

[LF08] Uwe Ligges and John Fox. R Help Desk: How can I avoid this loop or make it faster? R News: the Newsletter of the R Project, 8 (1):46–50, May 2008. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Leng:2013:CRC

[LG13] Jing Leng and Jingjing Guan. Conference review: The 6th Chinese R Conference. The R Journal, 5(1):210–212, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_chinese-r-conf.pdf.

Lambert:2018:RRP

[LGE^{+}18] Joshua Lambert, Liyu Gong, Corrine F. Elliott, Katherine Thompson, and Arnold Stromberg. rFSA: an R package for finding best subsets and interactions. The R Journal, 10(2):295–308, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-059.

Li:2018:PLE

[LGSG18] Yanming Li, Brenda W. Gillespie, Kerby Shedden, and John A. Gillespie. Profile likelihood estimation of the correlation coefficient in the presence of left, right or interval censoring and missing data. The R Journal, 10(2):159–179, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-040.

LeMeur:2006:AFC

[LH06] Nolwenn Le Meur and Florian Hahne. Analyzing flow cytometry data with Bioconductor. R News: the Newsletter of the R Project, 6(5): 27–32, December 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Lahti:2017:RAE

[LHKB17] Leo Lahti, Janne Huovari, Markus Kainu, and Przemysław Biecek. Retrieval and analysis of Eurostat Open Data with the eurostat package. The R Journal, 9(1):385–392, June 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-019.

Leiva:2006:NPB

[LHR06] Víctor Leiva, Hugo Hernández, and Marco Riquelme. A new package for the Birnbaum– Saunders distribution. R News: the Newsletter of the R Project, 6(4):35–40, October 2006. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Liaw:2007:RJV

[Lia07] Andy Liaw. Review of John Verzani’s book: Using R for Introductory Statistics. R News: the Newsletter of the R Project, 7(1):41–42, April 2007. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Ligges:2002:RHD

[Lig02] Uwe Ligges. R Help Desk: Automation of mathematical annotation in plots. R News: the Newsletter of the R Project, 2(3):32–34, December 2002. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Ligges:2003:RHDa

[Lig03a] Uwe Ligges. R Help Desk: Getting help — R’s help facilities and manuals. R News: the Newsletter of the R Project, 3(1):26–28, 2003. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Ligges:2003:RHDb

[Lig03b] Uwe Ligges. R Help Desk: Package management. R News: the Newsletter of the R Project, 3(3):37–39, December 2003. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Ligges:2006:RHD

[Lig06] Uwe Ligges. R Help Desk: Accessing the sources. R News: the Newsletter of the R Project, 6(4):43–45, October 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Liu:2014:PRP

[Liu14] Fang Liu. gset: An R package for exact sequential test of equivalence hypothesis based on bivariate non-central t-statistics. The R Journal, 6(2):174–184, December 2014. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2014-2/
RJournal_2014-2_liu.pdf.

Lucas:2006:UPP

[LJ06] Antoine Lucas and Sylvain Jasson. Using amap and ctc packages for huge clustering. R News: the Newsletter of the R Project, 6(5): 58–60, December 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Lederer:2006:SIS

[LK06] Wolfgang Lederer and Helmut Küchenhoff. A short introduction to the SIMEX and MC-SIMEX. R News: the Newsletter of the R Project, 6(4):26–31, October 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Lu:2013:PAE

[LK13] Yang Lu and David Kane. Performance attribution for equity portfolios. The R Journal, 5(2):53–62, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_lu-kane.pdf.

Liu:2015:ZRP

[LK15] Fang Liu and Yunchuan Kong. zoib: an R package for Bayesian inference for beta regression and zero/one inflated beta regression. The R Journal, 7(2):34–51, December 2015.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2015/RJ-2015-019.

LEcuyer:2005:PSR

[LL05] Pierre L’Ecuyer and Josef Leydold. rstream: Streams of random numbers for stochastic simulation. R News: the Newsletter of the R Project, 5(2):16–20, November 2005. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

LafayedeMicheaux:2009:PRP

[LL09] Pierre Lafaye de Micheaux and Benoît Li-

quet. ConvergenceConcepts: an R package

to investigate-various-modes-of-convergence. The R Journal, 1(2):18–25, December 2009. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/archive/2009-
2/RJournal_2009-2_Lafaye~de-Micheaux+
Liquet.pdf.

Linares:2016:QRP

[LLM16] Daniel Linares and Joan López-Moliner. quickpsy: an R package to fit psychometric functions for multiple groups. The R Journal, 8(1):122–131, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-008.

Ligges:2005:RHD

[LM05] Uwe Ligges and Duncan Murdoch. R Help Desk: Make ‘R CMD’ work under Windows — an example. R News: the Newsletter of the R Project, 5(2):27–28, November 2005. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Lau:2007:PEI

[LMK07] Olivia Lau, Ryan T. Moore, and Michael Kellermann. eiPack: R × C ecological inferences and higher-dimension data management. R News: the Newsletter of the R Project, 7(2): 43–47, October 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Lunardon:2014:PPC

[LMT14] Nicola Lunardon, Giovanna Menardi, and Nicola Torelli. ROSE: A package for class imbalance learning. The R Journal, 6(1):79–89, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_menardi-lunardon-torelli.pdf.

Lagonigro:2020:ARP

[LOM20] Raymond Lagonigro, Ramon Oller, and Joan Carles Martori. AQuadtree: an R package for quadtree anonymization of point data. The R Journal, 12(2):209–225, December 2020. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2021/RJ-
2021-013.

Lombardi:2014:PPS

[LP14] Luigi Lombardi and Massimiliano Pastore. sgr: A package for simulating conditional fake ordinal data. The R Journal, 6 (1):164–177, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_lombardi-pastore.pdf.

Liu:2018:ASI

[LQ18] Boxiang Liu and Thomas Quertermous. Approximating the sum of independent nonidentical binomial random variables. The R Journal, 10(1):472–483, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-011.

Li:2001:PCS

[LR01] Michael Na Li and A. J. Rossini. RPVM: Cluster statistical computing in R. R News: the Newsletter of the R Project, 1(3):4–7, September 2001. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Lopez-Raton:2017:GRP

[LRMLLCS17] Mónica López-Ratón, Elisa M. Molanes-López, Emilio Letón, and Carmen Cadarso-Suárez. GsymPoint: an R package to estimate the generalized symmetry point, an optimal cut-off point for binary classification in continuous diagnostic tests. The R Journal, 9(1):262–283, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-015.

Lloyd-Smith:2020:KTM

[LS20] Patrick Lloyd-Smith. Kuhn–Tucker and multiple discrete-continuous extreme value model estimation and simulation in R: The rmdcev package. The R Journal, 12(2): 251–265, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-015.

Lumley:2001:PNM

[Lum01] Thomas Lumley. Programmer’s niche: Macros in R. R News: the Newsletter of the R Project, 1(3):11–13, September 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Lumley:2003:ASD

[Lum03] Thomas Lumley. Analysing survey data in R. R News: the Newsletter of the R Project, 3 (1):17–20, June 2003. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Lumley:2004:PP

[Lum04a] Thomas Lumley. The survival package. R News: the Newsletter of the R Project, 4(1): 26–28, June 2004. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Lumley:2004:PNS

[Lum04b] Thomas Lumley. Programmers’ niche: A simple class, in S3 and S4. R News: the Newsletter of the R Project, 4(1):33–36, June 2004. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Lundholm:2011:ICC

[Lun11] Michael Lundholm. Implementing the compendium concept with Sweave and DOC-STRIP. The R Journal, 3(2):16–21, December 2011. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2011-2/RJournal_2011-2_Lundholm.pdf.

Luque:2007:DBA

[Luq07] Sebastián P. Luque. Diving behaviour analysis in R. R News: the Newsletter of the R Project, 7(3):8–14, December 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Lafitte:2011:CBF

[LVV11] Frédéric Lafitte, Dirk Van Heule, and Julien Van hamme. Cryptographic Boolean functions with R. The R Journal, 3(1):44–47, June 2011. CODEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2011-1/
                                        RJournal_2011-1_Lafitte~et~al.pdf.

Liaw:2002:CRP

[LW02] Andy Liaw and Matthew Wiener. Classification and regression by randomForest. R News: the Newsletter of the R Project, 2(3): 18–22, December 2002. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Lee:2020:CRP

[LZN^{+}20] Chi Hyun Lee, Heng Zhou, Jing Ning, Diane D. Liu, and Yu Shen. CoxPhLb: an R package for analyzing length biased data under Cox model. The R Journal, 12(1):118–130, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-024.

Marchini:2002:PRP

[Mar02] Jonathan Marchini. AnalyzeFMRI: An R package for the exploration and analysis of MRI and fMRI datasets. R News: the Newsletter of the R Project, 2(1):17–23, March 2002. CO-DEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Marschner:2011:PFG

[Mar11] Ian C. Marschner. glm2: Fitting generalized linear models with convergence problems. The R Journal, 3(2):12–15, December 2011. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2011-2/
RJournal_2011-2_Marschner.pdf.

Matloff:2019:E

[Mat19] Norm Matloff. Editorial. The R Journal, 11(1):4–6, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019-1/editorial.pdf.

Moritz:2017:ITS

[MBB17] Steffen Moritz and Thomas Bartz-Beielstein. imputeTS: Time series missing value imputation in R. The R Journal, 9(1):207–218, June 2017. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2017/RJ-
2017-009.

Mihaljevicacute:2018:BLB

[MBL18] Bojan Mihaljevic´,ConchaBielza, andP edroLarrañaga.bnclassify : LearningBayesiannetworkclassifiers.TheRJournal, 10

(2) : 455 − −468, December2018.ISSN2073 −

                                        4859.URLhttpsjournalrprojectorgarchive2018RJ2018073.

Mair:2014:WTT

[MC14] Patrick Mair and Scott Chamberlain. Web technologies task view. The R Journal, 6(1):178–181, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2014-1/RJournal_
                                        2014-1_mair-chamberlain.pdf.

Martinkova:2018:STP

[MD18] Patrícia Martinková and Adéla Drabinová.

ShinyItemAnalysis for teaching psychometrics and to enforce routine analysis of educational tests. The R Journal, 10(2): 503–515, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-074.

Mazoyer:2017:FRP

[MDDY17] Adrien Mazoyer, Rémy Drouilhet, Stéphane Despréaux, and Bernard Ycart. flan: an R package for inference on mutation models. The R Journal, 9(1):334–351, June 2017.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2017/RJ-2017-029.

Mevik:2006:PP

[Mev06] Bjørn-Helge Mevik. The pls package. R News: the Newsletter of the R Project, 6(3): 12–17, August 2006. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Meyer:2001:SVM

[Mey01] David Meyer. Support vector machines. R News: the Newsletter of the R Project, 1(3): 23–26, September 2001. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Meyer:2002:NTS

[Mey02] David Meyer. Naive time series forecasting methods. R News: the Newsletter of the R Project, 2(2):7–10, June 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.

R-project.org/doc/Rnews/.

Martinez:2019:TSF

[MFCR19] Francisco Martínez, María P. Frías, Francisco Charte, and Antonio J. Rivera. Time series forecasting with KNN in R: the tsfknn package. The R Journal, 11(2): 229–242, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-004.

Mair:2007:PTV

[MH07] Patrick Mair and Reinhold Hatzinger. Psychometrics task view. R News: the Newsletter of the R Project, 7(3):38–40, December 2007. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Machler:2014:RFN

[MH14] Martin Mächler and Kurt Hornik. R Foundation news. The R Journal, 6(2):190–191, December 2014. CODEN ???? ISSN 2073-

4859. URL http://journal.r-project.
org/archive/2014-2/RJournal_2014-2_r-
foundation.pdf.

Machler:2015:RFN

[MH15] Martin Mächler and Kurt Hornik. R Foundation news. The R Journal, 7(1):200, June 2015. CODEN ???? ISSN 2073-

4859. URL http://journal.r-project.
org/archive/2015-1/RJournal_2015-1_r-
foundation.pdf.

Melita:2017:DRP

[MH17] Nicolae Teodor Melita and Stefan Holban. dGAselID: an R package for selecting a variable number of features in high dimensional data. The R Journal, 9(2):18–34, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-040.

Morris:2020:SRP

[MHL^{+}20] Emily Morris, Kevin He, Yanming Li, Yi Li, and Jian Kang. SurvBoost: an R package for high-dimensional variable selection in the stratified proportional hazards model via gradient boosting. The R Journal, 12(1):105–117, June 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020/RJ-2020-018.

Morina:2015:GHD

[MHPO15] David Moriña, Manuel Higueras, Pedro Puig, and María Oliveira. Generalized Hermite distribution modelling with the R package hermite. The R Journal, 7(2): 263–274, December 2015. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2015/RJ-2015-035.

McDaniel:2013:FPR

[MHR13] Lee S. McDaniel, Nicholas C. Henderson, and Paul J. Rathouz. Fast pure R implementation of GEE: Application of the Matrix package. The R Journal, 5(1):181–188, June 2013. CODEN ???? ISSN 2073-4859. URL http:

                                        //journal.r-project.org/archive/2013-
                                        1/RJournal_2013-1_mcdaniel-henderson-
                                        rathouz.pdf.

Mineo:2003:NPG

[Min03] Angelo Mineo. A new package for the general error distribution. R News: the Newsletter of the R Project, 3(2):13–16, October 2003. CO-DEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Mainzer:2019:CRP

[MK19] Rheanna Mainzer and Paul Kabaila. ciuupi: an R package for computing confidence intervals that utilize uncertain prior information. The R Journal, 11(1):323–336, June 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-026.

McGowan:2020:TAR

[MKL20] Lucy D’Agostino McGowan, Sean Kross, and Jeffrey Leek. Tools for analyzing R code the tidy way. The R Journal, 12(1):226–242, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-011.

Medina:2019:RPT

[MKRPC19] Lily Medina, Ann-Kristin Kreutzmann, Natalia Rojas-Perilla, and Piedad Castro. The R package trafo for transforming linear regression models. The R Journal, 11 (2):99–123, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-054.

Murrell:2012:DGG

[ML12] Paul Murrell and Velvet Ly. Debugging grid graphics. The R Journal, 4(2):19–27, December 2012. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2012-2/RJournal_2012-2_Murrell+Ly.pdf.

Morales:2017:NPL

[MLG^{+}17] Pablo Morales, Julián Luengo, Luís P. F. Garcia, Ana C. Lorena, André C. P. L. F. de Carvalho, and Francisco Herrera. The NoiseFiltersR package: Label noise preprocessing in R. The R Journal, 9(1):219–228, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-027.

Molina:2015:PRP

[MM15] Isabel Molina and Yolanda Marhuenda. sae: An R package for small area estimation. The R Journal, 7(1):81–98, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2015-1/RJournal_
                                        2015-1_molina-marhuenda.pdf.

Mi:2009:PNN

[MMH09] Xuefei Mi, Tetsuhisa Miwa, and Torsten Hothorn. mvtnorm: New numerical algorithm for multivariate normal probabilities. The R Journal, 1(1):37–39, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-1/RJournal_
                                        2009-1_Mi+et+al.pdf.

Mori:2016:DMT

[MML16] Usue Mori, Alexander Mendiburu, and Jose A. Lozano. Distance measures for time series in R: The TSdist package. The R Journal, 8(2):451–459, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-058.

Mazucheli:2017:MTR

[MMN17] Josmar Mazucheli, André Felipe B. Menezes, and Saralees Nadarajah. mle.tools: an R package for maximum likelihood bias correction. The R Journal, 9(2):268–290, December 2017. ISSN 2073-4859. URL https://

                                        journal.r-project.org/archive/2017/RJ-
                                        2017-055.

Millan-Martinez:2020:GET

[MMO20] Pere Millán-Martínez and Ramon Oller. A graphical EDA tool with ggplot2: brinton. The R Journal, 12(2):311–320, December 2020. ISSN 2073-4859. URL https://

                                        journal.r-project.org/archive/2021/RJ-
                                        2021-018.

Meira-Machado:2016:CRP

[MMS16] Luis Meira-Machado and Marta Sestelo. condSURV: an R package for the estimation of the conditional survival function for ordered multivariate failure time data. The R Journal, 8(2):460–473, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-059.

Messner:2016:HCT

[MMZ16] Jakob W. Messner, Georg J. Mayr, and Achim Zeileis. Heteroscedastic censored and truncated regression with crch. The R Journal, 8(1):173–181, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-012.

Miettinen:2018:FFA

[MNT18] Jari Miettinen, Klaus Nordhausen, and Sara Taskinen. fICA: FastICA algorithms and their improved variants. The R Journal, 10 (2):148–158, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-046.

Margreitter:2017:MVM

[MO17] Christian Margreitter and Chris Oostenbrink. MDplot: Visualise molecular dynamics. The R Journal, 9(1):164–186, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-007.

Mazur:2020:LFS

[MO20] Stepan Mazur and Dmitry Otryakhin. Linear fractional stable motion with the rlfsm R package. The R Journal, 12(1):386–405, June 2020. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2020/RJ-
2020-008.

Moraga:2018:SAD

[Mor18] Paula Moraga. Small area disease risk estimation and visualization using R. The R Journal, 10(1):495–506, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-036.

Moraga:2020:SDM

[Mor20] Paula Moraga. Species distribution modeling using spatial point processes: a case study of sloth occurrence in Costa Rica. The R Journal, 12(2):293–310, December 2020.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2021/RJ-2021-017.

Murrell:2014:PP

[MP14] Paul Murrell and Simon Potter. The gridSVG package. The R Journal, 6(1):133–143, June 2014. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2014-1/RJournal_2014-1_murrell-potter.
                                        pdf.

Martin:2006:ABI

[MQ06] Andrew D. Martin and Kevin M. Quinn. Applied Bayesian inference in R using MCMCpack. R News: the Newsletter of the R Project, 6 (1):2–7, March 2006. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Murrell:2006:NSF

[MR06] Paul Murrell and Brian Ripley. Non-standard fonts in PostScript and PDF graphics. R News: the Newsletter of the R Project, 6(2): 41–47, May 2006. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Medina-Rodriguez:2017:AIH

[MRS17] Nathan Medina-Rodríguez and Ángelo Santana. Allele imputation and haplotype determination from databases composed of nuclear families. The R Journal, 9(2):35–55, December 2017. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2017/RJ-
2017-057.

Muenchow:2017:RIR

[MSB17] Jannes Muenchow, Patrick Schratz, and Alexander Brenning. RQGIS: Integrating R with QGIS for statistical geocomputing. The R Journal, 9(2):409–428, December 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-067.

Muschelli:2014:PII

[MSC14] John Muschelli, Elizabeth Sweeney, and Ciprian Crainiceanu. brainR: Interactive 3 and 4D images of high resolution neuroimage data. The R Journal, 6(1):41–48, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_muschelli-sweeney-crainiceanu.
pdf.

Miecznikowski:2018:TRP

[MsHCV18] Jeffrey C. Miecznikowski, En shuo Hsu, Yanhua Chen, and Albert Vexler. testforDEP: an R package for modern distribution-free tests and visualization tools for independence. The R Journal, 10(2):282–294, December 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-057.

Muschelli:2015:PCF

[MSLC15] John Muschelli, Elizabeth M. Sweeney, Martin Lindquist, and Ciprian M. Crainiceanu. fslr: Connecting the FSL software with R. The R Journal, 7(1):163–175, June 2015. CO-DEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/archive/2015-
1/RJournal_2015-1_muschelli-sweeney-
lindquist-etal.pdf.

Moore:2016:QRP

[MSN^{+}16] Christopher M. Moore, Christopher R. Stieha, Ben C. Nolting, Maria K. Cameron, and Karen C. Abbott. QPot: an R package for stochastic differential equation quasipotential analysis. The R Journal, 8 (2):19–38, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-031.

Murdoch:2009:NRH

[MU09] Duncan Murdoch and Simon Urbanek. The new R help system. The R Journal, 1 (2):60–65, December 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-2/RJournal_
                                        2009-2_Murdoch+Urbanek.pdf.

Muggeo:2008:PRP

[Mug08] Vito M. R. Muggeo. segmented: An R package to fit regression models with broken-line relationships. R News: the Newsletter of the R Project, 8(1):20–25, May 2008. CODEN ???? ISSN 1609-3631. URL http://CRAN.

                                        R-project.org/doc/Rnews/.

Murdoch:2002:RFF

[Mur02a] Duncan Murdoch. Reading foreign files. R News: the Newsletter of the R Project, 2(1): 2–3, March 2002. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Murrell:2002:PGP

[Mur02b] Paul Murrell. The grid graphics package. R News: the Newsletter of the R Project, 2 (2):14–19, June 2002. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Murrell:2003:IGG

[Mur03] Paul Murrell. Integrating grid graphics output with base graphics output. R News: the Newsletter of the R Project, 3(2):7–12, October 2003. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Murrell:2004:FLT

[Mur04] Paul Murrell. Fonts, lines, and transparency in R graphics. R News: the Newsletter of the R Project, 4(2):5–9, September 2004. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Murrell:2005:RCG

[Mur05a] Paul Murrell. Recent changes in grid graphics. R News: the Newsletter of the R Project, 5(1):12–20, May 2005. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Murtagh:2005:CAD

[Mur05b]  Fionn Murtagh.  Correspondence analy-

sis and data coding with Java and R. Computer science and data analysis series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-528-

9. 230 pp. LCCN QA76.73.J38 M877

2005. URL http://www.loc.gov/catdir/
enhancements/fy0648/2005041353-d.html.

Murrell:2007:VBF

[Mur07] Paul Murrell. Viewing binary files with the hexView package. R News: the Newsletter of the R Project, 7(1):2–8, April 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Murrell:2008:CNI

[Mur08] Paul Murrell. Comparing non-identical objects. R News: the Newsletter of the R Project, 8(2):40–47, October 2008. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Murrell:2009:DDR

[Mur09] Paul Murrell. Drawing diagrams with R. The R Journal, 1(1):15–21, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-1/RJournal_
                                        2009-1_Murrell.pdf.

Murdoch:2010:SR

[Mur10] Duncan Murdoch. Source references. The R Journal, 2(2):16–19, December 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2010-2/RJournal_
                                        2010-2_Murdoch.pdf.

Murrell:2011:RIR

[Mur11] Paul Murrell. Raster images in R graphics. The R Journal, 3(1):48–54, June 2011. CO-DEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2011-1/
                                        RJournal_2011-1_Murrell.pdf.

Murrell:2012:WYD

[Mur12a] Paul Murrell. It’s not what you draw, it’s what you don’t draw. The R Journal, 4 (2):13–18, December 2012. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2012-2/RJournal_
                                        2012-2_Murrell2.pdf.

Murrell:2012:WN

[Mur12b] Paul Murrell. What’s in a name? The R Journal, 4(2):5–12, December 2012. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2012-2/RJournal_
                                        2012-2_Murrell.pdf.

Murrell:2013:CP

[Mur13] Paul Murrell. Changes to grid from 3.0.0. The R Journal, 5(2):148–161, December 2013. CODEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2013-2/
                                        RJournal_2013-2_murrell.pdf.

Murrell:2015:PP

[Mur15] Paul Murrell. The gridGraphics package. The R Journal, 7(1):151–162, June 2015. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2015-1/
RJournal_2015-1_murrell.pdf.

Martin:2020:SEI

[MUX20] R. Douglas Martin, Chindhanai Uthaisaad, and Daniel Z. Xia. Skew-t expected information matrix evaluation and use for standard error calculations. The R Journal, 12(1):188–205, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-019.

Moretti:2019:ORE

[MVT19] Diego Moretti, Luca Valentino, and Tiziana Tuoto. Optimization routines for enforcing one-to-one matches in record linkage problems. The R Journal, 11(1):185–197, June 2019. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2019/RJ-
2019-008.

Michna:2013:RPR

[MW13] Pavel Michna and Milton Woods. RNetCDF — a package for reading and writing NetCDF datasets. The R Journal, 5(2): 29–37, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_michna-woods.pdf.

Mount:2018:DPS

[MZ18] John Mount and Nina Zumel. Dot-pipe: an S3 extensible pipe for R. The R Journal, 10(2):309–316, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-042.

Mayer:2018:RDR

[MZB^{+}18] Fernando P. Mayer, Walmes M. Zeviani, Wagner H. Bonat, Elias T. Krainski, and Paulo J. Ribeiro, Jr. R Day report. The R Journal, 10(1):551–554, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018-1/rday.pdf.

Nash:2020:PRG

[Nas20] John C. Nash. Provenance of R’s gradient optimizers. The R Journal, 12(1):477–483, June 2020. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2020/RJ-
2020-029.

Nadarajah:2013:PRP

[NB13] S. Nadarajah and S. A. A. Bakar. CompLognormal: An R package for composite lognormal distributions. The R Journal, 5(2):97– 104, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_nadarajah-bakar.pdf.

Nguyen:2020:FRP

[NDWL20] Robert Nguyen, James Day, David Warton, and Oscar Lane. fitzRoy — an R package to encourage reproducible sports analysis. The R Journal, 12(2):82–106, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2021/RJ-2021-005.

Nust:2020:RPA

[NEB^{+}20] Daniel Nüst, Dirk Eddelbuettel, Dom Bennett, Robrecht Cannoodt, Dav Clark, Gergely Daróczi, Mark Edmondson, Colin Fay, Ellis Hughes, Lars Kjeldgaard, Sean Lopp, Ben Marwick, Heather Nolis, Jacqueline Nolis, Hong Ooi, Karthik Ram, Noam Ross, Lori Shepherd, Péter Sólymos, Tyson Lee Swetnam, Nitesh Turaga, Charlotte Van Petegem, Jason Williams, Craig Willis, and Nan Xiao.

The Rockerverse: Packages and applications for containerisation with R. The R Journal, 12(1):437–461, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-007.

Nielsen:2015:ARP

[Nie15] Bent Nielsen. apc: an R package for age–period–cohort analysis. The R Journal, 7(2):52–64, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-020.

North:2016:SPS

[Nor16] Michael J. North. SchemaOnRead: a package for schema-on-read in R. The R Journal, 8(1):269–275, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-019.

Nunes:2015:ARP

[NP15] Matthew A. Nunes and Dennis Prangle. abctools: an R package for tuning approximate Bayesian computation analyses. The R Journal, 7(2):189–205, December 2015.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2015/RJ-2015-030.

Nie:2012:PPN

[NR12] Zhenghua Nie and Jeffrey S. Racine. The crs package: Nonparametric regression splines for continuous and categorical predictors. The R Journal, 4(2):48–56, December 2012. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2012-2/RJournal_
2012-2_Nie+S~Racine.pdf.

Nabi:2017:CRP

[NS17] Razieh Nabi and Xiaogang Su. coxphMIC: an R package for sparse estimation of Cox proportional hazards models via approximated information criteria. The R Journal, 9(1):229–238, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-018.

Nunes:2014:STT

[NTE14] Matthew A. Nunes, Sarah L. Taylor, and Idris A. Eckley. Stationarity testing and texture counting in R. The R Journal, 6(1):20–30, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2014-1/RJournal_
                                        2014-1_nunes-taylor-eckley.pdf.

Nieuwenhuis:2017:WEC

[NtGP17] Rense Nieuwenhuis, Manfred te Grotenhuis, and Ben Pelzer. Weighted effect coding for observational data with wec. The R Journal, 9(1):477–485, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-017.

Nieuwenhuis:2012:PMT

[NtP12] Rense Nieuwenhuis, Manfred te Grotenhuis, and Ben Pelzer. influence.ME: Tools for detecting influential data in mixed effects models. The R Journal, 4(2):38–47, December 2012. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2012-2/RJournal_2012-2_Nieuwenhuis~et~
                                        al.pdf.

Newton:2007:PWA

[NW07] Richard Newton and Lorenz Wernisch. Rwui: A Web application to create user friendly Web interfaces for R scripts. R News: the Newsletter of the R Project, 7(2):32–35, October 2007. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Onofri:2007:URP

[OC07] Andrea Onofri and Egidio Ciriciofolo. Using R to perform the AMMI analysis on agriculture variety trials. R News: the Newsletter of the R Project, 7(1):14–19, April 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Osorio:2017:MRP

[OGP17] Daniel Osorio, Janneth González, and Andrés Pinzón. minval: an R package for MINimal VALidation of stoichiometric reactions. The R Journal, 9(1):114–123, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-031.

Oh:2014:ACT

[Oh14] Jimmy Oh. Automatic conversion of tables to oLongForm dataframes. The R Journal, 6(2):16–26, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_oh.pdf.

Ou:2019:WDP

[OHC19] Lu Ou, Michael D. Hunter, and Sy-Miin Chow. Whats for dynr: a package for linear and nonlinear dynamic modeling in R. The R Journal, 11(1):91–111, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-012.

Oh:2007:PRP

[OK07] Hee-Seok Oh and Donghoh Kim. SpherWave: An R package for analyzing scattered spherical data by spherical wavelets. R News: the Newsletter of the R Project, 7(3):2–7, December 2007. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Orr:2009:SSE

[OL09] Megan Orr and Peng Liu. Sample size estimation while controlling false discovery rate for microarray experiments using the ssize.fdr package. The R Journal, 1(1):47–53, June 2009. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2009-1/RJournal_2009-1_Orr+Liu.pdf.

ONeil:2015:IPS

[O’N15] Shawn T. O’Neil. Implementing persistent O(1) stacks and queues in R. The R Journal, 7(1):118–126, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2015-1/RJournal_
                                        2015-1_oneil.pdf.

Olmedo:2016:WTF

[OOFdlFS^{+}16] Guillermo Federico Olmedo, Samuel Ortega-Farías, Daniel de la Fuente-Sáiz, David Fonseca-Luego, and Fernando Fuentes-Peñailillo. water: Tools and functions to estimate actual evapotranspiration using land surface energy balance models in R. The R Journal, 8(2):352–369, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-051.

Ooms:2013:DID

[Oom13] Jeroen Ooms. Directions for improved dependency versioning in R. The R Journal, 5(1):197–207, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_ooms.pdf.

Olson:2018:SEE

[ORC^{+}18] Roman Olson, Kelsey L. Ruckert, Won Chang, Klaus Keller, Murali Haran, and Soon-Il An. Stilt: Easy emulation of time series AR(1) computer model output in multidimensional parameter space. The R Journal, 10(2):209–225, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-049.

Osorio:2015:PPD

[ORVT15] Daniel Osorio, Paola Rondón-Villarreal, and Rodrigo Torres. Peptides: A package for data mining of antimicrobial peptides. The R Journal, 7(1):4–14, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_osorio-rondonvillarreal-torres.
pdf.

Oana:2018:SAR

[OS18] Ioana-Elena Oana and Carsten Q. Schneider. SetMethods: an add-on R package for advanced QCA. The R Journal, 10(1):507–533, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-031.

Otava:2017:IMA

[OSS^{+}17a] Martin Otava, Rudradev Sengupta, Ziv Shkedy, Dan Lin, Setia Pramana, Tobias Verbeke, Philippe Haldermans, Ludwig A. Hothorn, Daniel Gerhard, Rebecca M. Kuiper, Florian Klinglmueller, and Adetayo Kasim. IsoGeneGUI: Multiple approaches for doseresponse analysis of microarray data using R. The R Journal, 9(1):14–26, June 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-002.

Ozenne:2017:RPR

[OSS^{+}17b] Brice Ozenne, Anne Lyngholm Sørensen, Thomas Scheike, Christian Torp-Pedersen, and Thomas Alexander Gerds. riskRegression: Predicting the risk of an event using Cox regression models. The R Journal, 9(2):440– 460, December 2017. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2017/RJ-2017-062.

Otto:2019:SRP

[Ott19]  Philipp Otto.  spGARCH: an R-package

for spatial and spatiotemporal ARCH and GARCH models. The R Journal, 11(2): 401–420, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-053.

Probst:2017:MCR

[PAC^{+}17] Philipp Probst, Quay Au, Giuseppe Casalicchio, Clemens Stachl, and Bernd Bischl. Multilabel classification with R package mlr. The R Journal, 9(1):352–369, June 2017.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2017/RJ-2017-012.

Palczewski:2018:LAP

[Pal18] Andrzej Palczewski. LP algorithms for portfolio optimization: The PortfolioOptim package. The R Journal, 10(1):308–327, July 2018.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2018/RJ-2018-028.

Papageorgiou:2018:BRP

[Pap18] Georgios Papageorgiou. BNSP: an R package for fitting Bayesian semiparametric regression models and variable selection. The R Journal, 10(2):526–548, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-069.

Pebesma:2005:CMS

[PB05] Edzer J. Pebesma and Roger S. Bivand. Classes and methods for spatial data in R. R News: the Newsletter of the R Project, 5(2): 9–13, November 2005. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Plummer:2006:PCD

[PBCV06] Martyn Plummer, Nicky Best, Kate Cowles, and Karen Vines. CODA: Convergence diagnosis and output analysis for MCMC. R News: the Newsletter of the R Project, 6(1):7–11, March 2006. CODEN ???? ISSN 1609-3631. URL

                                        http://CRAN.R-project.org/doc/Rnews/.

Perez-Bernabe:2020:MRP

[PBMSN20] Inmaculada Pérez-Bernabé, Ana D. Maldonado, Antonio Salmerón, and Thomas D.

Nielsen. MoTBFs: an R package for learning hybrid Bayesian networks using mixtures of truncated basis functions. The R Journal, 12(2):321–341, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-019.

Pebesma:2018:SFR

[Peb18] Edzer Pebesma. Simple features for R: Standardized support for spatial vector data. The R Journal, 10(1):439–446, July 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-009.

Peng:2006:IDU

[Pen06] Roger D. Peng. Interacting with data using the filehash package. R News: the Newsletter of the R Project, 6(4):19–24, October 2006. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Petzoldt:2003:RSP

[Pet03] Thomas Petzoldt. R as a simulation platform in ecological modelling. R News: the Newsletter of the R Project, 3(3):8–16, December 2003. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Pitsillou:2016:DDC

[PF16] Maria Pitsillou and Konstantinos Fokianos. dCovTS: Distance covariance/correlation for time series. The R Journal, 8(2):324–340, December 2016. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2016/RJ-
2016-049.

Pfeuffer:2017:CRP

[Pfe17] Marius Pfeuffer. ctmcd: an R package for estimating the parameters of a continuoustime Markov chain from discrete-time data. The R Journal, 9(2):127–141, December 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-038.

Perez-Fernandez:2018:NRP

[PFMCFC18] Sonia Pérez-Fernández, Pablo Martínez-Camblor, Peter Filzmoser, and Norberto Corral. nsROC: an R package for non-standard ROC curve analysis. The R Journal, 10(2): 55–74, December 2018. ISSN 2073-4859. URL

                                        https://journal.r-project.org/archive/
                                        2018/RJ-2018-043.

Pickup:2017:ORP

[PGCE17] Mark Pickup, Paul Gustafson, Davor Cubranic, and Geoffrey Evans. OrthoPanels: an R package for estimating a dynamic panel model with fixed effects using the orthogonal reparameterization approach. The R Journal, 9(1):60–76, June 2017. ISSN 2073-4859. URL

                                        https://journal.r-project.org/archive/
                                        2017/RJ-2017-003.

Pau:2009:PP

[PH09] Gregoire Pau and Wolfgang Huber. The hwriter package. The R Journal, 1 (1):22–24, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-1/RJournal_
                                        2009-1_Pau+Huber.pdf.

Peters:2002:PIP

[PHL02] Andrea Peters, Torsten Hothorn, and Berthold Lausen. ipred: Improved predictors. R News: the Newsletter of the R Project, 2(2):33–36, June 2002. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Pruim:2017:MPH

[PKH17] Randall Pruim, Daniel T. Kaplan, and Nicholas J. Horton. The mosaic package: Helping students to think with data using R. The R Journal, 9(1):77–102, June 2017.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2017/RJ-2017-024.

Pikounis:2005:VRP

[PL05] Bill Pikounis and Andy Liaw. The value of R for preclinical statisticians. R News: the Newsletter of the R Project, 5(1):44–47, May 2005. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Pramana:2010:PRP

[PLH^{+}10] Setia Pramana, Dan Lin, Philippe Haldermans, Ziv Shkedy, Tobias Verbeke, Hinrich Göhlmann, An De Bondt, Willem Talloen, and Luc Bijnens. IsoGene: An R package for analyzing dose-response studies in microarray experiments. The R Journal, 2(1):5–12, June 2010. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2010-1/RJournal_2010-1_Pramana~et~al.
pdf.

Plummer:2011:PCR

[Plu11] Martyn Plummer. Portable C++ for R packages. The R Journal, 3(2):60–63, December 2011. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2011-2/RJournal_2011-2_Plummer.pdf.

Plummer:2015:RCR

[Plu15]  Martyn Plummer.  The R Consortium

and the R Foundation. The R Journal, 7(2):288, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015-2/plummer.pdf.

Pebesma:2016:MUR

[PMH16] Edzer Pebesma, Thomas Mailund, and James Hiebert. Measurement units in R. The R Journal, 8(2):486–494, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-061.

Piliszek:2019:MMF

[PMM^{+}19] Radosław Piliszek, Krzysztof Mnich, Szymon Migacz, Paweł Tabaszewski, Andrzej Sułecki, Aneta Polewko-Klim, and Witold Rudnicki. MDFS: MultiDimensional feature selection in R. The R Journal, 11(1):198–229, June 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-019.

Polanco-Martinez:2019:BRP

[PMMEGM19] Josue M. Polanco-Martinez, Martin A. Medina-Elizalde, Maria Fernanda Sanchez Goni, and Manfred Mudelsee. BINCOR: an R package for estimating the correlation between two unevenly spaced time series. The R Journal, 11(1):170–184, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-035.

Pan:2020:TTL

[PMZ20] Yuqing Pan, Qing Mai, and Xin Zhang. TULIP: a toolbox for linear discriminant analysis with penalties. The R Journal, 12 (2):61–81, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-025.

Pocernich:2006:RRC

[Poc06] Matthew Pocernich. R’s role in the climate change debate. R News: the Newsletter of the R Project, 6(4):17–18, October 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Poisot:2011:PPE

[Poi11] Timothée Poisot. The digitize package: Extracting numerical data from scatterplots. The R Journal, 3(1):25–26, June 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-1/RJournal_
2011-1_Poisot.pdf.

Papastamoulis:2017:BRP

[PR17] Panagiotis Papastamoulis and Magnus Rattray. BayesBinMix: an R package for model based clustering of multivariate binary data.

The R Journal, 9(1):403–420, June 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-022.

Prangle:2020:GRP

[Pra20] Dennis Prangle. gk: an R package for the g-and-k and generalised g-and-h distributions. The R Journal, 12(1):7–20, June 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2020/RJ-2020-010.

Pritikin:2016:MBI

[PS16] Joshua N. Pritikin and Karen M. Schmidt. Model builder for item factor analysis with OpenMx. The R Journal, 8(1):182–203, August 2016. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2016/RJ-
2016-013.

Polzehl:2007:PPA

[PT07] Jörg Polzehl and Karten Tabelow. fmri: A package for analyzing fmri data. R News: the Newsletter of the R Project, 7(2):13–17, October 2007. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Peng:2004:NP

[PW04] Roger D. Peng and Leah J. Welty. The NMMAPSdata package. R News: the Newsletter of the R Project, 4(2):10–14, September 2004. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Prague:2017:CRP

[PWD17] Melanie Prague, Rui Wang, and Victor De Gruttola. CRTgeeDR: an R package for doubly robust generalized estimating equations estimations in cluster randomized trials with missing data. The R Journal, 9 (2):105–115, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-041.

Pahl:2006:PDC

[PZK06] Roman Pahl, Andreas Ziegler, and Inke R. König. GroupSeq: Designing clinical trials using group sequential designs. R News: the Newsletter of the R Project, 6(2):21–26, May 2006. CODEN ???? ISSN 1609-3631. URL

                                        http://CRAN.R-project.org/doc/Rnews/.

Quartagno:2019:JFP

[QGC19] Matteo Quartagno, Simon Grund, and James Carpenter. jomo: a flexible package for twolevel joint modelling multiple imputation. The R Journal, 11(2):205–228, December 2019.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2019/RJ-2019-028.

Qian:2014:PPM

[Qia14] Hai Qian. PivotalR: A package for machine learning on big data. The R Journal, 6(1):57–67, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2014-1/RJournal_
                                        2014-1_qian.pdf.

Qiu:2015:PUS

[Qiu15] Yixuan Qiu. showtext: Using system fonts in R graphics. The R Journal, 7(1):99–108, June 2015. CODEN ???? ISSN 2073-4859. URL

                                        http://journal.r-project.org/archive/
                                        2015-1/RJournal_2015-1_qiu.pdf.

Rahman:2018:SSQ

[Rah18]  Adam Rahman.  sdpt3r: Semidefinite

quadratic linear programming in R. The R Journal, 10(2):371–394, December 2018.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2018/RJ-2018-063.

Ranke:2006:FDR

[Ran06] Johannes Ranke. Fitting dose-response curves from bioassays and toxicity testing. R News: the Newsletter of the R Project, 3(3):7–12, August 2006. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Ray:2001:RMW

[Ray01] M. J. Ray. Rcgi 4: Making Web statistics even easier. R News: the Newsletter of the R Project, 1(1):20–21, January 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Rodiger:2015:RER

[RBB^{+}15] Stefan Rödiger, Michał Burdukiewicz, Konstantin Blagodatskikh, Michael Jahn, and Peter Schierack. R as an environment for reproducible analysis of DNA amplification experiments. The R Journal, 7(1):127–150, June 2015. CODEN ???? ISSN 2073-4859. URL http://

journal.r-project.org/archive/2015-1/
RJournal_2015-1_rodiger-burdukiewicz-
blagodatskikh-etal.pdf.

Rickert:2018:CRRb

[RBK18] Joseph Rickert, Naras Balasubramanian, and Michael Kane. Conference report: R / Medicine Report. The R Journal, 10 (2):581–582, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018-2/R_Medicine.pdf.

Roediger:2013:SMC

[RBS13] Stefan Roediger, Alexander Bohm, and Ingolf Schimke. Surface melting curve analysis with R. The R Journal, 5(2):37–53, December 2013. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2013-2/
RJournal_2013-2_roediger-bohm-schimke.
pdf.

Ropkins:2012:ODA

[RC12] Karl Ropkins and David C. Carslaw. openair — data analysis tools for the air quality community. The R Journal, 4(1):20–29, June 2012. CODEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2012-1/
                                        RJournal_2012-1_Ropkins+Carslaw.pdf.

Ribeiro:2001:PPG

[RD01] Paulo J. Ribeiro, Jr. and Peter J. Diggle. geoR: A package for geostatistical analysis. R News: the Newsletter of the R Project, 1 (2):14–18, June 2001. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Rivolli:2018:UPM

[RdC18] Adriano Rivolli and Andre C. P. L. F. de Carvalho. The utiml package: Multilabel classification in R. The R Journal, 10(2):24–37, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-041.

Rubio:2015:CPR

[RdV15] Angel Rubio and Fernando de Villar. Code profiling in R: a review of existing methods and an introduction to package GUIProfiler. The R Journal, 7(2):275–287, December 2015.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2015/RJ-2015-036.

Rueda:2020:RPN

[RFGC20] M. Rueda, R. Ferri-García, and L. Castro. The R package NonProbEst for estimation in non-probability surveys. The R Journal, 12(1):406–418, June 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020/RJ-2020-015.

Ripley:2001:DTC

[RH01] Brian D. Ripley and Kurt Hornik. Date– time classes. R News: the Newsletter of the R Project, 1(2):8–11, June 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.

                                        R-project.org/doc/Rnews/.

Ritchie:2006:TMH

[RH06] Matthew Ritchie and Wolfgang Huber. Transcript mapping with high-density tiling arrays. R News: the Newsletter of the R Project, 6(5): 23–27, December 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Ribatet:2007:PMP

[Rib07] Mathieu Ribatet. POT: Modelling peaks over a threshold. R News: the Newsletter of the R Project, 7(1):34–36, April 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Rickert:2016:CRU

[Ric16] Joe Rickert. Conference report: useR! 2016. The R Journal, 8(1):399–401, August 2016. ISSN 2073-4859. URL https://journal.

r-project.org/archive/2016-1/user2016.
pdf.

Rickert:2018:CRRa

[Ric18] Joseph Rickert. Conference report: R / Pharma 2018. The R Journal, 10 (2):579–580, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018-2/R_Pharma.pdf.

Ripley:2001:C

[Rip01a] Brian D. Ripley. Connections. R News: the Newsletter of the R Project, 1(1):16–17, January 2001. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Ripley:2001:IRU

[Rip01b] Brian D. Ripley. Installing R under Windows. R News: the Newsletter of the R Project, 1 (2):11–14, June 2001. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Ripley:2001:UDR

[Rip01c] Brian D. Ripley. Using databases with R. R News: the Newsletter of the R Project, 1(1): 18–20, January 2001. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Ripley:2002:TSR

[Rip02] Brian D. Ripley. Time series in R 1.5.0. R News: the Newsletter of the R Project, 2(2): 2–7, June 2002. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Ripley:2004:LLP

[Rip04] Brian D. Ripley. Lazy loading and packages in R 2.0.0. R News: the Newsletter of the R Project, 4(2):2–4, September 2004. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Ripley:2005:IFR

[Rip05a] Brian D. Ripley. Internationalization features of R 2.1.0. R News: the Newsletter of the R Project, 5(1):2–7, May 2005. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Ripley:2005:PTM

[Rip05b] Brian D. Ripley. Packages and their management in R 2.1.0. R News: the Newsletter of the R Project, 5(1):8–11, May 2005. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Ruckdeschel:2006:SCD

[RKSC06] Peter Ruckdeschel, Matthias Kohl, Thomas Stabla, and Florian Camphausen. S4 classes for distributions. R News: the Newsletter of the R Project, 6(2):2–6, May 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Robinson:2007:ARU

[Rob07] Andrew Robinson. Augmenting R with Unix tools. R News: the Newsletter of the R Project, 7(1):30–33, April 2007. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Roocks:2016:CPF

[Roo16] Patrick Roocks. Computing Pareto frontiers and database preferences with the rPref package. The R Journal, 8(2): 393–404, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-054.

Rougier:2001:WPT

[Rou01] Jonathan Rougier. What’s the point of ‘tensor’? R News: the Newsletter of the R Project, 1(2):26–27, June 2001. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Rougier:2005:LPC

[Rou05] Jonathan Rougier. Literate programming for creating and maintaining packages. R News: the Newsletter of the R Project, 5(1):35–39, May 2005. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Reisner:2019:BRP

[RPO^{+}19] John Reisner, Hieu Pham, Sigurdur Olafsson, Stephen Vardeman, and Jing Li. biclustermd: an R package for biclustering with missing values. The R Journal, 11(2):69–84, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-045.

Raftery:2005:PRP

[RPV05] Adrian E. Raftery, Ian S. Painter, and Christopher T. Volinsky. BMA: An R package for Bayesian model averaging. R News: the Newsletter of the R Project, 5(2):2–8, November 2005. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Ronnegaard:2010:PPF

[RSA10] Lars Rönnegård, Xia Shen, and Moudud Alam. hglm: A package for fitting hierarchical generalized linear models. The R Journal, 2(2):20–28, December 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2010-2/RJournal_
2010-2_Roennegaard~et~al.pdf.

Rebora:2014:PFT

[RSR14] Paola Rebora, Agus Salim, and Marie Reilly. bshazard: A flexible tool for nonparametric smoothing of the hazard function. The R Journal, 6(2):114–122, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_rebora-salim-reilly.pdf.

Sailer:2005:PPD

[Sai05] Oliver Sailer. crossdes: A package for design and randomization in crossover studies. R News: the Newsletter of the R Project, 5(2): 24–27, November 2005. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Sarkar:2002:L

[Sar02] Deepayan Sarkar. Lattice. R News: the Newsletter of the R Project, 2(2):19–23, June 2002. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Sartore:2013:PMS

[Sar13] Luca Sartore. spMC: Modelling spatial random fields with continuous lag Markov chains. The R Journal, 5(2):16–29, December 2013. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2013-2/
RJournal_2013-2_sartore.pdf.

Sarkar:2014:Ea

[Sar14a] Deepayan Sarkar. Editorial. The R Journal, 6(1):3, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_editorial.pdf.

Sarkar:2014:Eb

[Sar14b] Deepayan Sarkar. Editorial. The R Journal, 6(2):3, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_editorial.pdf.

Sawitzki:2002:QCE

[Saw02] Günther Sawitzki. Quality control and early diagnostics for cDNA microarrays. R News: the Newsletter of the R Project, 2(1):6–10, March 2002. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Sariyar:2010:PPD

[SB10] Murat Sariyar and Andreas Borg. The RecordLinkage package: Detecting errors in data. The R Journal, 2(2): 61–67, December 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2010-2/RJournal_
2010-2_Sariyar+Borg.pdf.

Staniak:2018:EMP

[SB18] Mateusz Staniak and Przemysław Biecek. Explanations of model predictions with live and breakDown packages. The R Journal, 10(2):395–409, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-072.

Staniak:2019:LRP

[SB19] Mateusz Staniak and Przemysław Biecek. The landscape of R packages for automated exploratory data analysis. The R Journal, 11 (2):347–369, December 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-033.

Spyrison:2020:SRP

[SC20] Nicholas Spyrison and Dianne Cook. spinifex: an R package for creating a manual tour of low-dimensional projections of multivariate data. The R Journal, 12(1):243–257, June 2020. ISSN 2073-4859. URL https://

                                        journal.r-project.org/archive/2020/RJ-
                                        2020-027.

Sevinc:2019:RPP

[SCEG19] Busra Sevinc, Bekir Cetintav, Melek Esemen, and Selma Gurler. RSSampling: a pioneering package for ranked set sampling. The R Journal, 11(1):401–415, June 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-039.

Schlather:2001:SAR

[Sch01] Martin Schlather. Simulation and analysis of random fields. R News: the Newsletter of the R Project, 1(2):18–20, June 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.

                                        R-project.org/doc/Rnews/.

Schwartz:2003:RHD

[Sch03] Marc Schwartz. R Help Desk: An introduction to using R’s base graphics. R News: the Newsletter of the R Project, 3(2):2–6, October 2003. CODEN ???? ISSN 1609-3631. URL

                                        http://CRAN.R-project.org/doc/Rnews/.

Schwartz:2004:DUR

[Sch04] Marc Schwartz. The decision to use R. R News: the Newsletter of the R Project, 4(1):

2–5, June 2004. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Scholtens:2006:PCM

[Sch06] Denise Scholtens. Protein complex membership estimation using apComplex. R News: the Newsletter of the R Project, 6(5):32–36, December 2006. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Schwarzer:2007:PRP

[Sch07] Guido Schwarzer. meta: An R package for meta-analysis. R News: the Newsletter of the R Project, 7(3):40–45, December 2007. CO-DEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Schweiker:2016:CRP

[Sch16] Marcel Schweiker. comf: an R package for thermal comfort studies. The R Journal, 8(2):341–351, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-050.

Sorensen:2020:BRP

[SCLV20] Øystein Sørensen, Marta Crispino, Qinghua Liu, and Valeria Vitelli. BayesMallows: an R package for the Bayesian Mallows model. The R Journal, 12(1):324–342, June 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2020/RJ-2020-026.

Scrucca:2004:PRP

[Scr04] Luca Scrucca. qcc: An R package for quality control charting and statistical process control. R News: the Newsletter of the R Project, 4(1):11–17, June 2004. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Scrucca:2017:SEG

[Scr17] Luca Scrucca. On some extensions to GA package: Hybrid optimisation, parallelisation and islands evolution. The R Journal, 9(1):187–206, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-008.

Siberchicot:2013:CRD

[SD13] Aurelie Siberchicot and Stephane Dray. Conference report: Deuxièmes rencontres R. The R Journal, 5(2):164–166, December 2013. CO-DEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2013-2/
                                        RJournal_2013-2_siberchicot-dray.pdf.

Sun:2020:CCB

[SD20] Tao Sun and Ying Ding. CopulaCenR: Copula based regression models for bivariate censored data in R. The R Journal, 12(1):266–282, June 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020/RJ-2020-025.

Sarda-Espinosa:2019:TSC

[SE19] Alexis Sardá-Espinosa. Time-series clustering in R using the dtwclust package. The R Journal, 11(1):22–43, June 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-023.

Speranza:2018:SAA

[SFdC18] Talitha F. Speranza, Pedro C. Ferreira, and Jonatha A. da Costa. SARIMA analysis and automated model reports with BETS, an R package. The R Journal, 10 (2):133–147, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-070.

Scrucca:2016:MCC

[SFMR16] Luca Scrucca, Michael Fop, T. Brendan Murphy, and Adrian E. Raftery. mclust 5:

Clustering, classification and density estimation using Gaussian finite mixture models. The R Journal, 8(1):289–317, August 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-021.

Sachs:2016:IPS

[SG16] Michael C. Sachs and Erin E. Gabriel. An introduction to principal surrogate evaluation with the pseval package. The R Journal, 8(2):277–292, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-046.

Sklyar:2006:IAM

[SH06] Oleg Sklyar and Wolfgang Huber. Image analysis for microscopy screens. R News: the Newsletter of the R Project, 6(5):12–16, December 2006. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Sablica:2020:MCF

[SH20] Lukas Sablica and Kurt Hornik. mistr: a computational framework for mixture and composite distributions. The R Journal, 12(1):283–299, June 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020/RJ-2020-003.

Shang:2011:PRP

[Sha11] Han Lin Shang. rainbow: An R package for visualizing functional time series. The R Journal, 3(2):54–59, December 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-2/RJournal_
2011-2_Lin~Shang.pdf.

Shang:2013:PRP

[Sha13] Han Shang. ftsa: An R package for analyzing functional time series. The R Journal, 5(1):64–73, June 2013. CODEN ????

ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_shang.pdf.

Stanfill:2014:PRP

[SHG14] Bryan Stanfill, Heike Hofmann, and Ulrike Genschel. rotations: An R package for SO(3) data. The R Journal, 6(1):68–78, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2014-1/RJournal_
                                        2014-1_stanfill-hofmann-genschel.pdf.

Savoy:2017:APB

[SHR17] Heather Savoy, Falk Heße, and Yoram Rubin. anchoredDistr: a package for the Bayesian inversion of geostatistical parameters with multi-type and multi-scale data. The R Journal, 9(2):6–17, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-034.

Sawicka:2018:SUP

[SHW18] Kasia Sawicka, Gerard B. M. Heuvelink, and Dennis J. J. Walvoort. Spatial uncertainty propagation analysis with the spup R package. The R Journal, 10(2):180–199, December 2018. ISSN 2073-4859. URL https://

                                        journal.r-project.org/archive/2018/RJ-
                                        2018-047.

Strobl:2009:P

[SHZ09] Carolin Strobl, Torsten Hothorn, and Achim Zeileis. Party on! The R Journal, 1 (2):14–17, December 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2009-2/RJournal_
                                        2009-2_Strobl~et~al.pdf.

Sievert:2014:TPX

[Sie14] Carson Sievert. Taming PITCHf/x data with XML2R and pitchRx. The R Journal, 6(1):5–19, June 2014. CODEN ????

ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_sievert.pdf.

Selosse:2020:ORP

[SJB20] Margot Selosse, Julien Jacques, and Christophe Biernacki. ordinalClust: an R package to analyze ordinal data. The R Journal, 12(2):173– 188, December 2020. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2021/RJ-2021-011.

Siberchicot:2017:ASL

[SJLD^{+}17] Aurélie Siberchicot, Alice Julien-Laferrière, Anne-Béatrice Dufour, Jean Thioulouse, and Stéphane Dray. adegraphics: an S4 lattice-based package for the representation of multivariate data. The R Journal, 9 (2):198–212, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-042.

Schramm:2020:KPK

[SJO^{+}20] Catherine Schramm, Sébastien Jacquemont, Karim Oualkacha, Aurélie Labbe, and Celia M. T. Greenwood. KSPM: a package for kernel semi-parametric models. The R Journal, 12(2):189–208, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2021/RJ-2021-012.

Schwender:2006:IIG

[SKI06] Holger Schwender, Andreas Krause, and Katja Ickstadt. Identifying interesting genes with siggenes. R News: the Newsletter of the R Project, 6(5):45–50, December 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Sharma:2017:SNN

[SKZS17] Ruby Sharma, Sajal Kumar, Hua Zhong, and Mingzhou Song. Simulating noisy, nonparametric, and multivariate discrete patterns.

The R Journal, 9(2):366–377, December 2017.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2017/RJ-2017-053.

Syring:2017:BRP

[SL17] Nicholas Syring and Meng Li. BayesBD: an R package for Bayesian inference on image boundaries. The R Journal, 9 (2):149–162, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017/RJ-2017-052.

Sadeghi:2012:GMM

[SM12] Kayvan Sadeghi and Giovanni M. Marchetti. Graphical Markov models with mixed graphs in R. The R Journal, 4(2): 65–73, December 2012. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2012-2/RJournal_
                                        2012-2_Sadeghi+Marchetti.pdf.

Sansano:2019:IPF

[SMBTS19] Emilio Sansano, Raúl Montoliu, Óscar Belmonte, and Joaquín Torres-Sospedra. Indoor positioning and fingerprinting: The R package ipft. The R Journal, 11(1):67–90, June 2019.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2019/RJ-2019-010.

Smith:2003:WBM

[Smi03] Colin A. Smith. Web-based microarray analysis using Bioconductor. R News: the Newsletter of the R Project, 3(2):17–20, October 2003. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/Rnews/.

Saavedra-Nieves:2018:IRP

[SN18] Alejandro Saavedra-Nieves. Inventorymodel: an R package for centralized inventory problems. The R Journal, 10(1):200–217, July 2018. ISSN 2073-4859. URL https://

                                        journal.r-project.org/archive/2018/RJ-
                                        2018-023.

Silge:2018:NRP

[SNG18] Julia Silge, John C. Nash, and Spencer Graves. Navigating the R package universe. The R Journal, 10(2):558–563, December 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-058.

Solymos:2010:PDC

[Sól10] Péter Sólymos. dclone: Data cloning in R. The R Journal, 2(2):29–37, December 2010. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2010-2/
RJournal_2010-2_Solymos.pdf.

Schafer:2006:REG

[SORS06] Juliane Schäfer, Rainer Opgen-Rhein, and Korbinian Strimmer. Reverse engineering genetic networks using the GeneNet package. R News: the Newsletter of the R Project, 6(5): 50–53, December 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

South:2011:PNR

[Sou11] Andy South. rworldmap: A new R package for mapping global data. The R Journal, 3(1):35–43, June 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-1/RJournal_
2011-1_South.pdf.

Seo:2015:ARP

[SP15] Kangwon Seo and Rong Pan. ALTopt: an R package for optimal experimental design of accelerated life testing. The R Journal, 7(2):177–188, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-029.

Sichera:2019:CPP

[SP19] Roberto Sichera and Pietro Pizzuto. ConvergenceClubs: a package for performing the Phillips and Sul’s club convergence clustering procedure. The R Journal, 11(2):142–151, December 2019.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2019/RJ-2019-021.

Spedicato:2017:DTM

[Spe17] Giorgio Alfredo Spedicato. Discrete time Markov chains with R. The R Journal, 9(2):84–104, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-036.

Soetaert:2010:SDE

[SPS10] Karline Soetaert, Thomas Petzoldt, and R. Woodrow Setzer. Solving differential equations in R. The R Journal, 2(2):5–15, December 2010. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2010-2/RJournal_2010-2_Soetaert~et~al.
pdf.

Sax:2013:TDT

[SS13] Christoph Sax and Peter Steiner. Temporal disaggregation of time series. The R Journal, 5(2):80–88, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_sax-steiner.pdf.

Stoer:2016:MIP

[SS16] Nathalie C. Støer and Sven Ove Samuelsen. multipleNCC: Inverse probability weighting of nested case-control data. The R Journal, 8(2):5–18, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-030.

Salfran:2018:GAM

[SS18] Daniel Salfran and Martin Spiess. Generalized additive model multiple imputation by chained equations with package ImputeRobust. The R Journal, 10(1):61–72, July 2018. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2018/RJ-
2018-014.

Stephenson:2002:PEV

[Ste02] Alec Stephenson. evd: Extreme value distributions. R News: the Newsletter of the R Project, 2(2):31–32, June 2002. CODEN ???? ISSN 1609-3631. URL http://CRAN.

R-project.org/doc/Rnews/.

Stocco:2014:CBM

[Sto14] Andrea Stocco. Coordinate-based metaanalysis of fMRI studies with R. The R Journal, 6(2):5–15, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_stocco.pdf.

Strobel:2019:ACT

[Str19] Armin Ströbel. atable: Create tables for clinical trial reports. The R Journal, 11(1):137–148, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-001.

Sucarrat:2013:PSE

[Suc13] Genaro Sucarrat. betategarch: Simulation, estimation and forecasting of Betaskew-t-EGARCH models. The R Journal, 5 (2):137–148, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_sucarrat.pdf.

Sucarrat:2020:USG

[Suc20] Genaro Sucarrat. User-specified general-to-specific and indicator saturation methods. The R Journal, 12(2):388–401, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2021/RJ-2021-024.

Sestelo:2016:FRP

[SVMMRP16] Marta Sestelo, Nora M. Villanueva, Luis Meira-Machado, and Javier Roca-Pardiñas. FWDselect: an R package for variable selection in regression models. The R Journal, 8(1):132–148, August 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-009.

Schloerke:2016:EB

[SWCH16] Barret Schloerke, Hadley Wickham, Dianne Cook, and Heike Hofmann. Escape from boxland. The R Journal, 8 (2):243–257, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-044.

Shu:2019:IRP

[SY19] Di Shu and Grace Y. Yi. ipwErrorY: an R package for estimation of average treatment effect with misclassified binary outcome. The R Journal, 11(1):337–351, June 2019.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2019/RJ-2019-029.

Szepannek:2018:CUF

[Sze18] Gero Szepannek. clustMixType: Userfriendly clustering of mixed-type data in R. The R Journal, 10(2):200–208, December 2018. ISSN 2073-4859. URL https://

                                        journal.r-project.org/archive/2018/RJ-
                                        2018-048.

Szkaliczki:2016:CSD

[Szk16] Tibor Szkaliczki. clustering.sc.dp: Optimal clustering with sequential constraint by using dynamic programming. The R Journal, 8(1):318–327, August 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-022.

Tyner:2017:NVG

[TBH17] Sam Tyner, François Briatte, and Heike Hofmann. Network visualization with ggplot2. The R Journal, 9(1):27–59, June 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-023.

Thiem:2013:PPQ

[TD13] Alrik Thiem and Adrian Dusa. QCA: A package for qualitative comparative analysis. The R Journal, 5(1):87–98, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_thiem-dusa.pdf.

Tang:2018:GRP

[TDAP18] Hui Tang, Elizabeth L. Day, Molly B. Atkinson, and Norbert J. Pienta. GrpString: an R package for analysis of groups of strings. The R Journal, 10(1):359–369, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-002.

Team:2016:NBPa

[Tea16] The Bioconductor Team. News from the Bioconductor Project. The R Journal, 8 (1):404–405, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016-1/bioconductor.pdf.

Team:2017:NBPa

[Tea17a]  Bioconductor Core Team.  News from

the Bioconductor Project. The R Journal, 9(1):508, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017-1/bioc.pdf.

Team:2017:NBPb

[Tea17b]  Bioconductor Core Team.  News from

the Bioconductor Project. The R Journal, 9(2):567, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017-2/bioc.pdf.

Team:2018:NBPa

[Tea18a]  Bioconductor Core Team.  News from

the Bioconductor Project. The R Journal, 10(1):560, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018-1/bioc.pdf.

Team:2018:NBPb

[Tea18b] Bioconductor Core Team. News from the Bioconductor Project. The R Journal, 10(2):587, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018-2/bioc.pdf.

Team:2019:NBP

[Tea19] Bioconductor Core Team. News from the Bioconductor Project. The R Journal, 11(2):447, December 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019-2/bioc.pdf.

Team:2020:NBP

[Tea20] Bioconductor Core Team. News from the Bioconductor Project. The R Journal, 12(2):410, December 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020-2/bioc.pdf.

Turner:2007:PPG

[TF07] Heather Turner and David Firth. gnm: A package for generalized nonlinear models. R News: the Newsletter of the R Project, 7(2): 8–12, October 2007. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Team:2013:CRa

[The13a] The R.Core Team. Changes in R. The R Journal, 5(1):221–239, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_r-changes.pdf.

Team:2013:CRb

[The13b] The R.Core Team. Changes in R. The R Journal, 5(2):192–198, December 2013. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2013-2/
RJournal_2013-2_r-changes.pdf.

Team:2014:NBPb

[The14a] The Bioconductor Team. News from the Bioconductor Project. The R Journal, 6 (2):227–228, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_bioconductor.pdf.

Team:2014:CRa

[The14b] The R.Core Team. Changes in R. The R Journal, 6(1):221–235, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_r-changes.pdf.

Team:2014:CRb

[The14c] The R.Core Team. Changes in R. The R Journal, 6(2):224–226, December 2014. CO-DEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2014-2/
RJournal_2014-2_r-changes.pdf.

Team:2015:NBPa

[The15a] The Bioconductor Team. News from the Bioconductor Project. The R Journal, 7 (1):239–240, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2015-1/RJournal_
2015-1_bioconductor.pdf.

Team:2015:NBPb

[The15b] The Bioconductor Team. News from the Bioconductor Project. The R Journal, 7 (2):291–292, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015-2/bioconductor.pdf.

Team:2015:CRb

[The15c] The R Core Team. Changes in R. The R Journal, 7(2):293–297, December 2015. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2015-2/r-changes.pdf.

Team:2015:CRa

[The15d] The R.Core Team. Changes in R. The R Journal, 7(1):227–238, June 2015. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2015-1/RJournal_
                                        2015-1_r-changes.pdf.

Team:2016:NBPb

[The16a] The Bioconductor Core Team. News from the Bioconductor Project. The R Journal, 8(2):510, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016-2/bioc.pdf.

Team:2016:CRa

[The16b] The R Core Team. Changes in R. The R Journal, 8(1):406–415, August 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016-1/r-changes.pdf.

Team:2016:CRb

[The16c] The R Core Team. Changes in R. The R Journal, 8(2):511–514, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016-2/ch.pdf.

Team:2017:CRa

[The17a] The R Core Team. Changes in R. The R Journal, 9(1):509–521, June 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017-1/ch.pdf.

Team:2017:CRb

[The17b] The R Core Team. Changes in R. The R Journal, 9(2):568–570, December 2017. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2017-2/ch.pdf.

Team:2018:CRa

[The18a] The R Core Team. Changes in R. The R Journal, 10(1):561–570, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018-1/ch.pdf.

Team:2018:CRb

[The18b] The R Core Team. Changes in R. The R Journal, 10(2):588–590, December 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018-2/ch.pdf.

Team:2019:RNa

[The19a] The R Core Team. R news. The R Journal, 11(1):442–582, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019-1/ch.pdf.

Team:2019:RNb

[The19b] The R Core Team. R news. The R Journal, 11(2):448–554, December 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019-2/ch.pdf.

Tang:2016:GUI

[THL16] Yuan Tang, Masaaki Horikoshi, and Wenxuan Li. ggfortify: Unified interface to visualize statistical results of popular R packages. The R Journal, 8(2):474–485, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-060.

Thomas:2006:BL

[Tho06] Andrew Thomas. The BUGS language. R News: the Newsletter of the R Project, 6(1): 17–21, March 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Tierney:2001:UNM

[Tie01] Luke Tierney. Under new memory management. R News: the Newsletter of the R Project, 1(1):10–11, January 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Tierney:2003:NSM

[Tie03] Luke Tierney. Name space management for R. R News: the Newsletter of the R Project, 3(1):2–6, June 2003. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Templ:2015:PGG

[TKM15] Matthias Templ, Alexander Kowarik, and Bernhard Meindl. sparkTable: Generating graphical tables for Websites and documents with R. The R Journal, 7(1):24–37, June 2015. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2015-1/
RJournal_2015-1_templ-kowarik-meindl.
pdf.

Taylor:2007:IMT

[TLP07] Sandra L. Taylor, Duncan Temple Lang, and Katherine S. Pollard. Improvements to the multiple testing package multtest. R News: the Newsletter of the R Project, 7(3):52–55, December 2007. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Thomas:2006:MBO

[TOLS06] Andrew Thomas, Bob O’Hara, Uwe Ligges, and Sibylle Sturtz. Making BUGS open. R News: the Newsletter of the R Project, 6(1): 12–17, March 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Touloumis:2016:SCB

[Tou16]  Anestis Touloumis.  Simulating corre-

lated binary and multinomial responses under marginal model specification: The SimCorMultRes package. The R Journal, 8(2):

79–91, December 2016. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2016/RJ-2016-034.

Turner:2020:NFT

[Tur20]  Heather Turner.  News from the For-

wards Taskforce. The R Journal, 12(2): 411–412, December 2020. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2020-2/forwards-news.pdf.

Tvedebrink:2015:CRU

[Tve15] Torben Tvedebrink. Conference report: useR! 2015. The R Journal, 7(2):289–290, December 2015. ISSN 2073-4859. URL https:/

/journal.r-project.org/archive/2015-2/
tvedebrink.pdf.

Thioulouse:2010:ORR

[TVMZ10] Jean Thioulouse, Claire Valiente-Moro, and Lionel Zenner. Online reproducible research: An application to multivariate analysis of bacterial DNA fingerprint data. The R Journal, 2(1):44–52, June 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2010-1/RJournal_
2010-1_Thioulouse~et~al.pdf.

Theussl:2009:CSD

[TZ09] Stefan Theußl and Achim Zeileis. Collaborative software development using R-Forge. The R Journal, 1(1):9–14, June 2009. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2009-1/RJournal_
2009-1_Theussl+Zeileis.pdf.

Ullah:2018:LCR

[UAA18] Muhammad Imdad Ullah, Muhammad Aslam, and Saima Altaf. lmridge: a comprehensive R package for ridge regression. The R Journal, 10(2):326–346, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-060.

Ulrich:2013:CRR

[Ulr13] Joshua Ulrich. Conference report: R/Finance 2013. The R Journal, 5(1):212–215, June 2013. CODEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2013-1/
                                        RJournal_2013-1_r-finance.pdf.

Ucar:2018:MER

[UPA18] Iñaki Ucar, Edzer Pebesma, and Arturo Azcorra. Measurement errors in R. The R Journal, 10(2):549–557, December 2018.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2018/RJ-2018-075.

Committee:2013:RUC

[Use13] UseR 2013 Organising Committee. The R User Conference 2013. The R Journal, 5(1):215–218, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_user2013.pdf.

V:2020:ILM

[VAD20] Vineetha Warriyar K. V., Waleed Almutiry, and Rob Deardon. Individual-level modelling of infectious disease data: EpiILM. The R Journal, 12(1):87–104, June 2020. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2020/RJ-2020-020.

vanderLoo:2014:PRP

[van14] Mark van der Loo. stringdist: an R package for approximate string matching. The R Journal, 6(1):111–122, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2014-1/RJournal_
                                        2014-1_loo.pdf.

Vega-Bayo:2015:RPP

[VB15] Ainhoa Vega-Bayo. An R package for the panel approach method for program evaluation: pampe. The R Journal, 7 (2):105–121, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-024.

Vigneau:2015:CRP

[VCQ15] Evelyne Vigneau, Mingkun Chen, and El Mostafa Qannari. ClustVarLV: an R package for the clustering of variables around latent variables. The R Journal, 7(2):134– 148, December 2015. ISSN 2073-4859. URL

https://journal.r-project.org/archive/
2015/RJ-2015-026.

Valle:2013:PPM

[VD13] Giada Valle and Matteo Dell’Omodarme. stellaR: a package to manage stellar evolution tracks and isochrones. The R Journal, 5(1):108–117, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_valle-dellomodarme.pdf.

Valliant:2015:PCD

[VDK15] Richard Valliant, Jill A. Dever, and Frauke Kreuter. PracTools: Computations for design of finite population samples. The R Journal, 7(2):163–176, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-028.

Venables:2001:PN

[Ven01] Bill Venables. Programmer’s niche. R News: the Newsletter of the R Project, 1(1):27–30, January 2001. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Venables:2002:PN

[Ven02] Bill Venables. Programmer’s niche. R News: the Newsletter of the R Project, 2(2):24–26, June 2002. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Verzani:2005:URI

[Ver05] John Verzani. Using R for introductory statistics. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-450-9. xvi + 414 pp. LCCN QA276.4 .V47 2005.

Verzani:2007:IP

[Ver07] John Verzani. An introduction to gWidgets. R News: the Newsletter of the R Project, 7(3): 26–33, December 2007. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Verbeke:2017:CRU

[Ver17] Tobias Verbeke. Conference report: useR!2017. The R Journal, 9(2):537–538, December 2017. ISSN 2073-4859. URL https://journal.

                                        r-project.org/archive/2017-2/useR2017.
                                        pdf.

Verzani:2018:Ea

[Ver18a] John Verzani. Editorial. The R Journal, 10(1):4–5, July 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018-1/editorial.pdf.

Verzani:2018:Eb

[Ver18b] John Verzani. Editorial. The R Journal, 10(2):4–6, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018-2/editorial.pdf.

Vitolo:2016:RRP

[VFB16] Claudia Vitolo, Matthew Fry, and Wouter Buytaert. rnrfa: an R package to retrieve, filter and visualize data from the UK National River Flow Archive. The R Journal, 8(2):102–116, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-036.

Villacorta:2017:WPR

[Vil17] Pablo J. Villacorta. The welchADF package for robust hypothesis testing in unbalanced multivariate mixed models with heteroscedastic and non-normal data. The R Journal, 9(2):309–328, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-049.

VanDomelen:2014:FRF

[VP14] Dane R. Van Domelen and W. Stephen Pittard. Flexible R functions for processing accelerometer data, with emphasis on NHANES 2003–2006. The R Journal, 6 (2):52–62, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_domelen-pittard.pdf.

Villacorta:2015:SSR

[VS15] Pablo J. Villacorta and José A. Sáez. SRCS: Statistical ranking color scheme for visualizing parameterized multiple pairwise comparisons with R. The R Journal, 7 (2):89–104, December 2015. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2015/RJ-2015-023.

Vacca:2019:ITR

[VZ19] Gianmarco Vacca and Maria Grazia Zoia. Identifying and testing recursive vs. interdependent links in simultaneous equation models via the SIRE package. The R Journal, 11(1):149–169, June 2019. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2019/RJ-2019-016.

Ward:2019:MRE

[WAD^{+}19] Eric J. Ward, Sean C. Anderson, Luis A. Damiano, Mary E. Hunsicker, and Michael A. Litzow. Modeling regimes with extremes: the bayesdfa package for identifying and forecasting common trends and anomalies in multivariate time-series data. The R Journal, 11(2):46–55, December 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-007.

Wais:2016:GPM

[Wai16] Kamil Wais. Gender prediction methods based on first names with genderizeR. The R Journal, 8(1):17–37, August 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-002.

Walker:2016:TRP

[Wal16] Kyle Walker. tigris: an R package to access and work with geographic data from the US Census Bureau. The R Journal, 8 (2):231–242, December 2016. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2016/RJ-2016-043.

Warnes:2003:PP

[War03] Gregory R. Warnes. The genetics package. R News: the Newsletter of the R Project, 3 (1):9–13, June 2003. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Warnes:2006:SSE

[War06] Gregory R. Warnes. Sample size estimation for microarray experiments using the ssize package. R News: the Newsletter of the R Project, 6(5):64–68, December 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Werft:2010:PPR

[WB10] Wiebke Werft and Axel Benner. glmperm: A permutation of regressor residuals test for inference in generalized linear models. The R Journal, 2(1):39–43, June 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2010-1/RJournal_
2010-1_Werft+Benner.pdf.

Wangler:2006:PRP

[WBS06] M. Wangler, J. Beyersmann, and M. Schumacher. changeLOS: An R-package for change in length of hospital stay based on the Aalen– Johansen estimator. R News: the Newsletter of the R Project, 6(2):31–35, May 2006. CO-DEN ???? ISSN 1609-3631. URL http://

CRAN.R-project.org/doc/Rnews/.

Wang:2017:CRP

[WC17] Chao Wang and Kung-Sik Chan. carx: an R package to estimate censored autoregressive time series with exogenous covariates. The R Journal, 9(2):213–231, December 2017.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2017/RJ-2017-064.

Winslow:2016:SPC

[WCAR16] Luke A. Winslow, Scott Chamberlain, Alison P. Appling, and Jordan S. Read. sbtools: a package connecting R to cloud-based data for collaborative online research. The R Journal, 8(1):387–398, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-029.

Chang:2020:CMS

[wCTKL20] Hsin wen Chang, Pei-Yuan Tsai, Jen-Tse Kao, and Guo-You Lan. Comparing multiple survival functions with crossing hazards in R. The R Journal, 12(2):20–42, December 2020.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2021/RJ-2021-002.

Wilhelm:2013:ESP

[WdM13] Stefan Wilhelm and Miguel Godinho de Matos. Estimating spatial probit models in R. The R Journal, 5(1):130–144, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_wilhelm-matos.pdf.

Wang:2015:MGP

[WFRM15] Juhui Wang, Robert Faivre, Hervé Richard, and Hervé Monod. mtk: a general-purpose and extensible R environment for uncertainty and sensitivity analyses of numerical experiments. The R Journal, 7(2):206–226, December 2015.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2015/RJ-2015-031.

Weidmann:2010:MMC

[WG10] Nils B. Weidmann and Kristian Skrede Gleditsch. Mapping and measuring country shapes. The R Journal, 2(1):18– 24, June 2010. CODEN ???? ISSN

                                        2073-4859.  URL http://journal.r-
                                        project.org/archive/2010-1/RJournal_
                                        2010-1_Weidmann+Skrede~Gleditsch.pdf.

Wickham:2010:PMC

[Wic10] Hadley Wickham. stringr: Modern, consistent string processing. The R Journal, 2(2):38–40, December 2010. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2010-2/RJournal_
                                        2010-2_Wickham.pdf.

Wickham:2011:PGS

[Wic11] Hadley Wickham. testthat: Get started with testing. The R Journal, 3(1):5–10, June 2011. CODEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2011-1/
                                        RJournal_2011-1_Wickham.pdf.

Wickham:2013:Ea

[Wic13a] Hadley Wickham. Editorial. The R Journal, 5(1):4–6, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_editorial.pdf.

Wickham:2013:Eb

[Wic13b] Hadley Wickham. Editorial. The R Journal, 5(2):3–4, December 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-2/RJournal_
2013-2_editorial.pdf.

Wieland:2017:MAA

[Wie17] Thomas Wieland. Market area analysis for retail and service locations with MCI. The R Journal, 9(1):298–323, June 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-020.

Williams:2009:PDM

[Wil09] Graham J. Williams. Rattle: A data mining GUI for R. The R Journal, 1(2):45–55, December 2009. CODEN ???? ISSN 2073-4859. URL

http://journal.r-project.org/archive/
2009-2/RJournal_2009-2_Williams.pdf.

Winter:2017:RRP

[Win17] David J. Winter. rentrez: an R package for the NCBI eUtils API. The R Journal, 9(2):520–526, December 2017. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2017/RJ-2017-058.

Wickham:2008:IP

[WLLS08] Hadley Wickham, Michael Lawrence, Duncan Temple Lang, and Deborah F. Swayne. An introduction to rggobi. R News: the Newsletter of the R Project, 8(2):3–7, October 2008. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Wilhelm:2010:PPT

[WM10] Stefan Wilhelm and B. G. Manjunath. tmvtnorm: A package for the truncated multivariate normal distribution. The R Journal, 2(1):25–29, June 2010. CODEN ????

ISSN 2073-4859. URL http://journal.

r-project.org/archive/2010-1/RJournal_
2010-1_Wilhelm+Manjunath.pdf.

Wood:2001:PGG

[Woo01] Simon N. Wood. mgcv: GAMs and generalized ridge regression for R. R News: the Newsletter of the R Project, 1(2):20–25, June 2001. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Wurm:2018:SGL

[WR18] Michael J. Wurm and Paul J. Rathouz. Semiparametric generalized linear models with the gldrm package. The R Journal, 10(1):288–307, July 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-027.

Wright:2016:UDV

[Wri16] Erik S. Wright. Using DECIPHER v2.0 to analyze big biological sequence data in R. The R Journal, 8(1):352–359, August 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-025.

Weiss:2018:FCR

[WRR18] Christoph E. Weiss, Eran Raviv, and Gernot Roetzer. Forecast combinations in R using the ForecastComb package. The R Journal, 10(2):262–281, December 2018. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2018/RJ-2018-052.

Wang:2011:PDO

[WS11] Haizhou Wang and Mingzhou Song. Ckmeans.1d.dp: Optimal k-means clustering in one dimension by dynamic programming. The R Journal, 3(2):29–33, December 2011. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2011-2/RJournal_
2011-2_Wang+Song.pdf.

Wang:2013:PRP

[WS13] Weizhen Wang and Guogen Shan. ExactCIdiff: An R package for computing exact confidence intervals for the difference of two proportions. The R Journal, 5(2):62–71, December 2013. CODEN ???? ISSN 2073-4859. URL http:/

/journal.r-project.org/archive/2013-2/
RJournal_2013-2_wang-shan.pdf.

Wansouwe:2016:ARP

[WSK16] Wanbitching E. Wansouwé, Sobom M. Somé, and Célestin C. Kokonendji. Ake: an R package for discrete and continuous associated kernel estimations. The R Journal, 8 (2):258–276, December 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-045.

Watkins:2006:NLR

[WV06] Peter Watkins and Bill Venables. Non-linear regression for optimising the separation of carboxylic acids. R News: the Newsletter of the R Project, 6(3):2–7, August 2006. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Wabersich:2014:PPR

[WV14] Dominik Wabersich and Joachim Vandekerckhove. The RWiener package: An R package providing distribution functions for the Wiener diffusion model. The R Journal, 6(1):49–56, June 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-1/RJournal_
2014-1_vandekerckhove-wabersich.pdf.

Wehrens:2006:MDX

[WW06] Ron Wehrens and Egon Willighagen. Mapping databases of X-ray powder patterns. R News: the Newsletter of the R Project, 6(3): 24–28, August 2006. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Weisberg:2007:NH

[WW07] Sanford Weisberg and Hadley Wickham. Need a hint? R News: the Newsletter of the R Project, 7(3):36–38, December 2007. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Xie:2008:PPS

[XC08] Yihui Xie and Xiaoyue Cheng. animation: A package for statistical animations. R News: the Newsletter of the R Project, 8(2):23–27, October 2008. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-project.org/
                                        doc/Rnews/.

Xiang:2013:GSA

[XGSH13] Yang Xiang, Sylvain Gubian, Brian Suomela, and Julia Hoeng. Generalized simulated annealing for global optimization: the GenSA package for R. The R Journal, 5(1):13–29, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

                                        r-project.org/archive/2013-1/RJournal_
                                        2013-1_xiang-gubian-suomela-etal.pdf.

Xie:2005:RLR

[Xie05] Xianhong Xie. Running long R jobs with Condor DAG. R News: the Newsletter of the R Project, 5(2):13–15, November 2005. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

                                        project.org/doc/Rnews/.

Xiong:2019:SRP

[XY19] Juan Xiong and Grace Y. Yi. swgee: an R package for analyzing longitudinal data with response missingness and covariate measurement error. The R Journal, 11(1):416–426, June 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-031.

Yan:2002:PYA

[Yan02] Jun Yan. geepack: Yet another package for generalized estimating equations. R News: the Newsletter of the R Project, 2(3):12–14, December 2002. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Yan:2004:FRB

[Yan04] Jun Yan. Fusing R and BUGS through Wine. R News: the Newsletter of the R Project, 4(2): 19–21, September 2004. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Yee:2008:PP

[Yee08] Thomas W. Yee. The VGAM package. R News: the Newsletter of the R Project, 8(2):28–39, October 2008. CODEN ???? ISSN 1609-

3631. URL http://CRAN.R-project.org/
doc/Rnews/.

Yengo:2016:VCH

[YJBC16] Loïc Yengo, Julien Jacques, Christophe Biernacki, and Mickael Canouil. Variable clustering in high-dimensional linear regression: The R package clere. The R Journal, 8(1):92–106, August 2016. ISSN 2073-

4859. URL https://journal.r-project.
org/archive/2016/RJ-2016-006.

Yeomans:2018:PPD

[YKT18] Michael Yeomans, Alejandro Kantor, and Dustin Tingley. The politeness package: Detecting politeness in natural language. The R Journal, 10(2):489–502, December 2018.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2018/RJ-2018-079.

Yu:2019:CPC

[YLGT19] Hengshi Yu, Fan Li, John A. Gallis, and Elizabeth L. Turner. cvcrand: a package for covariate-constrained randomization and the clustered permutation test for cluster randomized trials. The R Journal, 11 (2):191–204, December 2019. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2019/RJ-2019-027.

Yoo:2018:BAS

[Yoo18] Jae Keun Yoo. Basis-adaptive selection algorithm in dr-package. The R Journal, 10 (2):124–132, December 2018. ISSN 2073-

                                        4859. URL https://journal.r-project.
                                        org/archive/2018/RJ-2018-045.

Young:2016:NTI

[You16] Derek S. Young. Normal tolerance interval procedures in the tolerance package. The R Journal, 8(2):200–212, December 2016.

                                        ISSN 2073-4859. URL https://journal.r-
                                        project.org/archive/2016/RJ-2016-041.

Yan:2003:BMW

[YR03] Jun Yan and A. J. Rossini. Building Microsoft Windows versions of R and R packages under Intel Linux. R News: the Newsletter of the R Project, 3(1):15–17, June 2003. CODEN ???? ISSN 1609-3631. URL http://CRAN.

                                        R-project.org/doc/Rnews/.

Yu:2002:PPS

[Yu02] Hao Yu. Rmpi: Parallel statistical computing in R. R News: the Newsletter of the R Project, 2(2):10–14, June 2002. CODEN ???? ISSN

                                        1609-3631. URL http://CRAN.R-project.
                                        org/doc/Rnews/.

Zabala:2014:PPE

[Zab14] Aiora Zabala. qmethod: A package to explore human perspectives using Q methodology. The R Journal, 6(2):163–173, December 2014. CO-DEN ???? ISSN 2073-4859. URL http:/

                                        /journal.r-project.org/archive/2014-2/
                                        RJournal_2014-2_zabala.pdf.

Zagaglia:2013:PMA

[Zag13] Paolo Zagaglia. PIN: Measuring asymmetric information in financial markets with R. The R Journal, 5(1):80–87, June 2013. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2013-1/RJournal_
2013-1_zagaglia.pdf.

Zeng:2020:BPM

[ZB20] Yaohui Zeng and Patrick Breheny. The biglasso package: a memory- and computationefficient solver for lasso model fitting with big data in R. The R Journal, 12(2):6–19, December 2020. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2021/RJ-
2021-001.

Zhu:2016:MRP

[ZC16] Xiang-Wei Zhu and Jian-Yi Chen. mixtox: an R package for mixture toxicity assessment. The R Journal, 8(2):421–433, December 2016.

ISSN 2073-4859. URL https://journal.r-
project.org/archive/2016/RJ-2016-056.

Zeileis:2001:PTS

[Zei01] Achim Zeileis. strucchange: Testing for structural change in linear regression relationships. R News: the Newsletter of the R Project, 1(3):8–11, September 2001. CODEN ???? ISSN 1609-3631. URL http://CRAN.R-

project.org/doc/Rnews/.

Zeileis:2005:CTV

[Zei05] Achim Zeileis. CRAN task views. R News: the Newsletter of the R Project, 5(1):39–40, May 2005. CODEN ???? ISSN 1609-3631. URL

http://CRAN.R-project.org/doc/Rnews/.

Zhang:2004:TIE

[ZG04] Jianhua Zhang and Robert Gentleman. Tools for interactively exploring R packages. R News: the Newsletter of the R Project, 4(1):

20–25, June 2004. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Zeileis:2002:DCR

[ZH02] Achim Zeileis and Torsten Hothorn. Diagnostic checking in regression relationships. R News: the Newsletter of the R Project, 2(3): 7–10, December 2002. CODEN ???? ISSN

1609-3631. URL http://CRAN.R-project.
org/doc/Rnews/.

Zhao:2006:DPD

[Zha06] Jing Hua Zhao. Drawing pedigree diagrams with R and graphviz. R News: the Newsletter of the R Project, 6(2):38–41, May 2006. CODEN ???? ISSN 1609-3631. URL http:/

/CRAN.R-project.org/doc/Rnews/.

Zhang:2014:P

[ZHC^{+}14] Jonathan Zhang, Nancy Heckman, Davor Cubranic, Joel G. Kingsolver, Travis Gaydos, and J. S. Marron. Prinsimp. The R Journal, 6(2):27–42, December 2014. CODEN ???? ISSN 2073-4859. URL http://journal.

r-project.org/archive/2014-2/RJournal_
2014-2_zhang-etal.pdf.

Zhu:2017:MRP

[ZM17] Xuwen Zhu and Volodymyr Melnykov. ManlyMix: an R package for manly mixture modeling. The R Journal, 9(2):176–197, December 2017. ISSN 2073-4859. URL https://

journal.r-project.org/archive/2017/RJ-
2017-060.

Zhu:2019:OSD

[ZZZ^{+}19] Ruoqing Zhu, Jiyang Zhang, Ruilin Zhao, Peng Xu, Wenzhuo Zhou, and Xin Zhang. orthoDr: Semiparametric dimension reduction via orthogonality constrained optimization. The R Journal, 11(2):24–37, December 2019. ISSN 2073-4859. URL https:// REFERENCES 98

journal.r-project.org/archive/2019/RJ-
2019-006.